### v2

# CoBRA: A Universal Strategyproof Confirmation Protocol for Quorum-based Proof-of-Stake Blockchains

Abstract
We present a formal analysis of quorum-based State Machine Replication (SMR) protocols in Proof-of-Stake (PoS) payment systems under a hybrid threat model comprising honest, Byzantine, and rational validators. Our analysis of traditional quorum-based protocols establishes two fundamental impossibility results: (1) in partially synchronous networks, no quorum-based protocol can achieve SMR when rational and Byzantine validators comprise more than $1/3$ of participants, and (2) even in synchronous networks, SMR remains impossible when rational and Byzantine validators comprise $2/3$ or more of participants.
Given these limitations, we propose two complementary solutions that achieve SMR in the presence of less than $2/3$ Byzantine and rational validators. To achieve this, we first introduce a protocol that enforces an upper bound on the total volume of transactions finalized within any time window $\Delta$ and prove that this bound is necessary for secure SMR protocols in our model. Second, we present the strongest chain rule, which enables efficient finalization of transactions when the majority of honest participants provably support the SMR execution. Through empirical analysis of Ethereum and Cosmos networks, we demonstrate that validator participation for over $99\%$ of blocks exceeds the required $5/6$ threshold, establishing the practical feasibility of our solution in production PoS systems.
1 Introduction
State Machine Replication (SMR) protocols are the backbone of fault-tolerant distributed systems, ensuring agreement on an ordered sequence of operations among participants. Traditionally, these protocols classify participants as either honest (strictly following the protocol), or Byzantine (deviating arbitrarily). However, this binary model fails to capture the behavior of rational participants, who strategically deviate only when it maximizes their utility.
This distinction is particularly critical in Proof-of-Stake (PoS) blockchains, which collectively secure $0.5$ trillion dollars. Participation in these systems is open and financially motivated, with validators staking assets to partake in consensus and earn rewards. While this incentive structure is designed to encourage honest behavior, it also introduces new attack vectors: profit-maximizing validators may deviate to increase their gains. Worse, external financial markets, such as derivative markets, may create additional incentives for validators to undermine the system. Capturing this multi-dimensional behavior requires a more nuanced model that accounts for honest, Byzantine, and rational validators. Ensuring the economic security of quorum-based SMR protocols (e.g., PBFT [1], HotStuff [2], Tendermint [3]) within this extended model is essential to safeguard existing PoS blockchains.
Several recent works examine the economic security of quorum-based PoS blockchains, e.g., [4, 5, 6, 7, 8, 9]. However, they either, like [6] adopt a different extended model and rely on a random beacon, which provides only probabilistic security. Or, in contrast, similar to [4] consider only Byzantine and honest validators but solely focus on slashing misbehaving validators while protecting honest stake, without accounting for liquid token transfers, which adversaries can exploit. Or, like [7], primarily focus on achieving economic safety after recovering from attacks. Yet, none of these works address the following fundamental question: Can existing PoS quorum-based SMR protocols be adapted to provably withstand both Byzantine and rational validators?
1.1 Contributions
We answer this question in the affirmative and introduce the first transformation that enables optimal rational resilience in PoS blockchains. We make the following contributions:
  1. 1.
    Impossibility Results: We prove that no quorum-based SMR protocol can be resilient when Byzantine and rational validators together exceed critical thresholds:
    • Partial Synchrony: Quorum-based SMR is impossible when Byzantine and rational validators collectively control $1/3$ or more of the validator set.
    • Synchrony: Quorum-based SMR is impossible when Byzantine and rational validators collectively control more than $2/3$ of the validator set.
  2. 2.
    Protocol Transformation: We introduce CoBRA (Correct over Byzantine Rational Adversary), a transformation that enhances existing quorum-based PoS protocols to tolerate rational behavior . Our approach comprises:
    • A finality gadget that detects equivocations and prevents rational validators from benefiting from forking. Our finality gadget safeguards honest validators as well as clients that do not actively monitor the SMR protocol execution.
    • A stake-bounded finalization rule that enables finalizing a total transaction volume equal to the stake validators lock in the system before a synchronous bound elapses.
    • A strongest-chain finalization rule that enables unbounded transaction throughput when at least $5/6$ of the validators actively follow the protocol.
    • An incentive mechanism consisting of (i) slashing conditions to disincentivize rational validators to equivocate in collusion with the Byzantine ones; (ii) a novel reward mechanism to incentivize rational validators to actively participate in the SMR protocol, and not engage in censoring or silencing honest proposers.
  3. 3.
    Incentive-based Security Analysis: We prove that CoBRA can securely transform the class of quorum-based SMR protocols to withstand rational participants, effectively reducing the required fraction of honest validators while maintaining security guarantees.
  4. 4.
    Empirical Validation: We evaluate validator participation rates in Ethereum and Cosmos, revealing they exceed the strongest-chain threshold for over 99% of the blocks, highlighting CoBRA is well-suited for deployment in existing systems.
2 Preliminaries
2.1 Model and Assumptions
In this section, we outline the model and assumptions used throughout the paper.
Proof-of-Stake-based blockchain
A PoS blockchain implements a distributed ledger protocol [10] (i.e., a state machine replication protocol), where participants are selected to create and validate new blocks based on the amount of coins (cryptocurrency) they hold and lock up as stake. The blockchain serves as a payment system in which participants hold and transfer fungible assets, referred to as coins. We also discuss how to extend this model to general-purpose smart contract platforms in Section 7. Each party is associated with a private-public key pair which forms its client identity in the system. The parties that actively participate in maintaining the state of the system (i.e., run the SMR) are known as validators. To do so, they lock coins (i.e., stake) and process transactions, providing clients an ordered sequence of confirmed transactions, referred to as the transaction ledger.
We assume the existence of a genesis block $G$, which contains the initial stake distribution and the identities of the validators participating in the SMR protocol. For the remainder of this paper, we assume a static validator setting, where the stake distribution is known in advance to all participants. In Section 7 we discuss how we can extend our system to a dynamic validator setting.
Validators
For simplicity, we assume a uniform model where all validators lock the same stake equal to $D$ coins and have thus the same “voting power”. If an entity locks some multiple $\alpha\cdot D$, they may control up to $\alpha$ individual validators. Thus, the SMR protocol involves a fixed set of validators, $N$, which are known to all participants. Let $n=|N|$ denote the total number of validators. We assume a public key infrastructure (PKI), where each validator is equipped with a public-private key pair for signing messages, and their public keys are available to all parties.
Clients
Unlike validators, the number of active clients is not known to all parties. Clients have public keys, which they can use to issue authenticated transactions. We assume clients are silent [11], meaning they neither actively interact with nor observe the execution of the SMR protocol among validators. Instead, clients query a subset of validators to obtain confirmation of any transaction (i.e., the transaction ledger) from their responses.
Cryptography
We assume that all participants are computationally bounded, i.e., they are modeled as probabilistic polynomial-time (PPT) interactive Turing machines. We further assume the existence of cryptographically secure communication channels, hash functions, and digital signatures.
Time
Time progresses in discrete logical units indexed as $u=0,1,2,\dots$. We assume that the maximum clock offset between any two parties is bounded, as any such offset can be absorbed into the network delay bound.
Network
We assume that all participants are connected through perfect point-to-point communication channels. We consider two communication models: partial synchrony and synchrony. In the former, we show impossibility results; in the latter, we present both impossibility and our main results.
(a) Partial synchronous model: There is an unknown time called Global Stabilization Time ($GST$), such that after $GST$ messages arrive within a known time bound $\Delta$. Therefore, every message sent at time $t$ eventually arrives at time $t^{\prime}\leq max\{GST,t\}+\Delta$.
(b) Synchronous model: Any message sent is guaranteed to arrive within a known time bound $\Delta^{*}$. Thus, a message sent at time $t$ is guaranteed to arrive by time $t+\Delta^{*}$.
When applying our transformation protocol to partially synchronous SMR protocols, we provide an alternative interpretation of our synchronous network model in Section 7, following prior work [12, 8].
Adversarial model
We consider a mixed model where validators are honest, Byzantine, or rational. Honest validators always follow the protocol. Byzantine validators deviate arbitrarily from the protocol. We assume a Byzantine adversary controls at most $f=(n-1)/3$ validators, and does not have any liquid coins, i.e., all their coins are staked to be able to maximize their influence.
Rational validators, on the other hand, will deviate from the protocol specification if and only if doing so maximizes their utility. We define a party’s utility as its monetary profit, calculated as the rewards from participation minus any potential penalties (e.g., slashing). For simplicity, we assume no loss in opportunity cost. We denote the number of rational validators in the system by $k$.
We make the following assumptions on rational validators: (a) The cost of operating a validator is assumed to be negligible. Consequently, any positive reward provides sufficient incentive for rational validators to participate. This assumption simplifies the analysis by canceling out the constant operational costs of running a PoS validator (e.g., bandwidth, CPU, storage) with a corresponding minimum constant reward. (b) Second, the cost of voting for a block proposed by another validator (e.g., adding a signature) is negligible and does not decrease their utility. (c) Third, a rational validator deviates from the protocol only if such deviation results in strictly greater utility, as devising a strategy to deviate may incur additional costs to running the default client. For example, a rational validator will censor or double spend only if they can secure a bribe larger than their potential loss due to slashing. This stands in contrast to Byzantine validators, who may deviate arbitrarily, capturing scenarios that fall outside the scope of our utility-based PoS model. (d) A rational validator considers a strategy only if the resulting utility is guaranteed within the transaction ledger, as we do not account for external trustless mechanisms  [^1] Any deviation from this assumption, such as validators influenced by external trustless mechanisms is modeled as Byzantine behavior.. Promises or unverifiable commitments from the adversary regarding future actions are not considered credible and, thus, will not influence the validator’s strategy. Given the appropriate incentives, however, rational validators may collude with other rational and Byzantine validators.
2.2 Definitions and Properties
At a high level, in an SMR protocol, validators must perform two primary tasks: order a set of transactions [^2] Validators execute the transactions in the agreed order to acquire the state of the state machine. and reply to clients. Traditional SMR protocols such as PBFT [1] and HotStuff [2] assume an honest supermajority of validators – specifically $12f+$ 1 out of the total $13f+$ 1 validators must be honest. Under this assumption, clients can extract the confirmation of transactions simply by waiting for responses from $f+1$ validators. In contrast, our setting introduces rational validators and assumes only an honest minority. For this reason, we separate these two responsibilities: (i) ordering transactions and (ii) replying to clients.
SMR problem
Given as input transactions, the validators execute an SMR protocol $\Pi$ and output an ordered sequence of confirmed transactions, known as the transaction ledger. We say that $\Pi$ solves the SMR problem when: i) correct validators agree in a total order [13], ii) the transaction ledger is certifiable.
Definition 1.
We say that correct validators agree in a total order, when the transaction ledger output by $\Pi$ satisfies:
  • Safety. The transaction ledgers of any two correct validators are prefixes of one another,
  • Liveness. Any valid transaction witnessed by all correct validators, will eventually be included in the transaction ledger of every correct node.
Definition 2.
We say that $\Pi$ satisfies certifiability if the following holds,
  • Client-safety: Clients accept the confirmation of a transaction only if it is included to the transaction ledger of at least one correct validator,
  • Client-liveness: Clients accept confirmation for any transaction included to the transaction ledger of all correct validators.
Resiliency of SMR protocols
Traditional SMR protocols [14, 1, 15] are secure when considering $f$ Byzantine validators and $n-f$ correct validators. We say that a protocol satisfying this condition is $(n,f)$-resilient.
Definition 3.
A protocol $\Pi$ executed by $n$ validators is $(n,f)$-resilient if it remains secure in the presence of up to $f$ Byzantine validators, assuming that the remaining $n-f$ validators follow the protocol correctly.
However, in a traditional $(n,f)$-resilient SMR protocol, one rational validator is enough to violate safety and liveness as Byzantine validators can bribe the rational validator to collude with them. We outline such attacks in Appendix A. To capture both byzantine and rational behavior in our protocol, we introduce the notion of $(n,k,f)$-resiliency.
Definition 4.
A protocol $\Pi$ executed by $n$ validators is $(n,k,f)$-resilient if it remains secure in the presence of $f$ Byzantine and $k$ rational validators, assuming that the remaining $n-f-k$ validators follow the protocol correctly.
$q$-commitable SMR protocols
We focus on a family of SMR protocols that progress by collecting certificates of $q$ votes across multiple phases, similar to several BFT SMR protocols (e.g. [1, 14]). We define a quorum certificate (QC) as a certificate consisting of $q$ votes. A protocol is q-commitable if a QC on a specific message $m$ (e.g., $m=\texttt{commit}$ in PBFT [1]) is both a necessary and sufficient condition for validators to commit a transaction (or block) $b$ to their local ledger.
Definition 5.
A protocol $\Pi$ is a $q$-commitable SMR protocol when every correct validator includes a block $b$ in its local transaction ledger if and only if it has witnessed a QC on a specific message $m$, denoted $QC_{m}(b)$.
For simplicity, we may often omit the specific message $m$ and refer to the certificate that finalizes $b$ as $QC(b)$. We stress that $QC(b)$ may not be sufficient for certifiability from the clients’ perspective, as they do not actively monitor the SMR execution (i.e., they are silent). However, for clients that monitor the SMR execution (e.g., the validators), $QC(b)$ is sufficient for certifiability.
We study SMR protocols in the context of proof-of-stake (PoS) blockchains, where validators batch transactions into blocks, as in HotStuff [2]. Our objective is to determine when an $(n,f)$-resilient SMR protocol can be transformed into an $(n,k,f)$-resilient PoS blockchain. Specifically, we present a transformation that converts any $(n,f)$-resilient $q$-certifiable PoS blockchain [1, 14, 15] into an $(n,k,f)$-resilient PoS blockchain. At the core of this transformation, we introduce a new certification mechanism for finalizing transactions, ensuring security under our hybrid model.
Finality latency
Finality latency refers to the time that elapses between a transaction being submitted by a client and its irreversible commitment to the transaction ledger.
2.3 SMR game
In this section, we relate game-theoretic rational behavior to protocol-level resilience guarantees. While $(n,k,f)$-resilience specifies that a protocol remains safe and live when at most $f$ validators are Byzantine and $k$ are rational, it does not explain why rational validators would not prefer to violate the protocol properties. To address this, we model the protocol execution as a finite game and introduce the notion of a rationally robust protocol. This notion ensures that rational validators have no incentive to deviate in ways that compromise safety or liveness, regardless of the strategies thereby ensuring the protocol’s resilience.
Definition 6 (SMR Game).
The SMR game is a tuple:
$G=(n,(S_{i})_{i\in[n]},(u_{i})_{i\in[n]},H)$
where:
  • $n\in\mathbb{N}$ is the number of players. Each player is either correct, Byzantine, or rational.
  • $H$ is the global history of the protocol execution. Each player $i\in[n]$ observes a local view $H_{i}\subseteq H$.
  • $S_{i}$ is the set of all possible strategies available to player $i$, where a strategy is a function mapping local histories $H_{i}$ to actions.
  • A global strategy profile is $s=(s_{1},s_{2},\dots,s_{n})\in S=\prod_{i\in[n]}S_{i}$.
  • The utility function $u_{i}:S\to\mathbb{R}$ gives the total payoff to player $i$ over the entire execution.
We denote by $s_{-i}$ the strategy profile excluding player $i$, and by $S_{-i}$ the set of all such profiles.
We partition the space of strategy profiles into two sets: $S_{\text{bad}}$, consisting of strategy profiles resulting in safety violations (i.e., conflicting blocks are finalized with valid certificates) or liveness violation (i.e., permanent censorship of a transaction), and $S_{good}$ which preserve both properties.
For each validator $p$, we define $S_{\mathsf{bad}}^{p}$ as the set of individual strategies $s_{p}$ for which there exists a strategy profile $s=(s_{p},s_{-p})\in S_{\mathsf{bad}}$ where $p$ actively contributes to the violation. Conversely, let $S_{\mathsf{good}}$ be the set of strategy profiles that preserve both safety and liveness. For each validator $p$, denote by $S_{\mathsf{good}}^{p}$ the set of individual strategies $s_{p}$ such that there exists a strategy profile $s=(s_{p},s_{-p})\in S_{\mathsf{good}}$.
We say that a protocol is rationally robust if, for every rational validator $p$, any strategy $s_{p}\in S_{\mathsf{bad}}^{p}$ that contributes to a violation is weakly dominated by some strategy $s_{p}^{\prime}\in S_{\mathsf{good}}^{p}$. In other words, rational validators always have an equally good or better option than causing violations.
Definition 7.
A strategy $s_{i}\in S_{i}$ weakly dominates a strategy $s_{i}^{\prime}\in S_{i}$ if, for every possible strategy profile $s_{-i}\in S_{-i}$, the following holds:
$u_{i}(s_{i},s_{-i})\geq u_{i}(s^{\prime}_{i},s_{-i}),$
and there exists at least one $s_{-i}\in S_{-i}$ such that:
$u_{i}(s_{i},s_{-i})>u_{i}(s^{\prime}_{i},s_{-i}).$
Definition 8.
We say that an SMR game is rationally robust if for every rational validator $p$, and every strategy $s_{p}\in S_{\mathsf{bad}}^{p}$, there exists a strategy $s_{p}^{\prime}\in S_{\mathsf{good}}^{p}$ that weakly dominates $s_{p}$.
Rational robustness relaxes traditional incentive compatibility, which requires that following the protocol is always the utility-maximizing strategy. Such a requirement is too strong in adversarial settings, where Byzantine validators may accumulate participation rewards and use them to bribe rational ones. Instead, rational robustness focuses only on the minimal conditions to guarantee safety and liveness.
3 Impossibility results
In the following, we identify settings in which a $q$-committable $(n,f)$–resilient PoS blockchain cannot be transformed into an $(n,k,f)$–resilient PoS blockchain. Additionally, we determine when $q$-committable SMR protocols fail to guarantee finality before the synchrony bound $\Delta^{*}$ elapses. We provide a summary of our results along with proof sketches and defer the complete proofs to Appendix C.
Partial synchrony
In Theorem 3, we show that designing an $(n,k,f)$-resilient $q$-commitable SMR protocol is impossible for $f\geq\lceil\frac{h}{2}\rceil$, where $h$ is the number of correct validators. Notice that this result depends only on the proportion of Byzantine and correct validators. This is because to guarantee liveness, the size of the quorum $q$ is at most $n-f$. For $f\geq\lceil\frac{h}{2}\rceil$, we cannot guarantee that two quorums intersect in at least one correct validator. Rational validators, in turn, will collude with Byzantine validators to form quorums of conflicting blocks to maximize their monetary gains.
To establish this result, we extend the proof of Budish et al. [12, Theorem 4.2] within our model. The core issue is that rational validators can collude with Byzantine to generate valid quorum certificates while avoiding penalties. This is possible because, for any predefined withdrawal period that allows validators to reclaim their stake, correct validators cannot reliably distinguish between an honest execution with network-induced delays and malicious behavior. Hence, rational validators may engage in equivocation without facing slashing, undermining the protocol’s security.
{restatable}
theoremPartialSynchrony Assuming silent clients, a static setting, and a partially synchronous network, no $q$-commitable SMR protocol can be $(n,k,f)$–resilient for any $q$ when $f\geq\lceil\frac{h}{2}\rceil$, where $h=n-k-f$.
Proof sketch of Theorem 3
Proof.
The adversary can partition correct validators in two distinct subsets $H_{1}$, $H_{2}$ and perform the double-spending attack described in Lemma 8.This attack can be beneficial for rational validators since for any value of stake $D$ they lock before the run, there is an input of conflicting transactions $T_{1}$ and $T_{2}$ where rational validators transfer all their liquid coins to distinct subset of clients $C_{1}$ and $C_{2}$ respectively. Notably, misbehaving validators can withdraw their stake without even being slashed.
Let us fix the stake $D$ and any withdrawal period $\Delta_{w}$. Consider as input a set of conflicting transactions $T_{1}$ and $T_{2}$ constructed as explained before. To guarantee liveness, correct validators must finalize any input of valid transactions. Therefore, validators in $H_{1}$ must finalize transactions in $T_{1}$, and validators in $H_{2}$ must finalize transactions in $T_{2}$.
However, since $GST$ occurs at an unknown time, it is impossible for correct validators in different subsets to detect the conflicting blocks including transactions in $T_{1},T_{2}$ before finalizing the transactions. Similarly, it impossible for the correct validators to slash the misbehaving ones before they withdraw their stake. ∎
It follows directly from Theorem 3 that no $q$-commitable protocol can be $(n,f)$–resilient when $f=\lceil h/2\rceil$, implying resilience is impossible if Byzantine validators exceed one-third of the network ($f\geq n/3$). Furthermore, even if a $q$-commitable protocol is resilient for $f=\lfloor h/2\rfloor$, replacing just one correct validator with a rational one breaks resilience if the total number of Byzantine and rational validators reaches at least one third the network ($f+k\geq n/3$), as shown in Corollary 3.1. This result highlights the fragility of quorum-based protocols under rational behavior, as even a small shift from correct to rational participants significantly impacts security.
Corollary 3.1.
Consider any $q$-commitable protocol $\Pi$ that is $(n,f)$–resilient for $f=\lfloor\frac{h}{2}\rfloor$ where $h=n-f$ is the number of correct validators. For any $k\geq 1$, $\Pi$ is not $(n,k,f)$–resilient in our model in the partial synchronous setting.
Corollary 3.1 implies that no quorum-based SMR protocol $\Pi$ can achieve $(3f+1,k,f)$–resilience for $k\geq 1$ in the partial synchronous model. That means that quorum-based SMR protocols are not resilient for $f+k\geq n/3$ under the model described in Theorem 3.
Synchrony
To circumvent the impossibility result of Theorem 3, we assume synchrony with a known time delay $\Delta^{*}$. However, we now show that even in synchronous settings, designing resilient SMR protocols remains impossible for $f\geq h$. Here, a client cannot distinguish whether correct validators participated in a quorum certificate. Consequently, Byzantine and rational validators can generate certificates that the clients will accept (to ensure client-liveness) without any participation from correct validators.
{restatable}
theoremSynchrony Assuming silent clients, a static setting, and a synchronous network, no $q$-commitable SMR protocol can be $(n,k,f)$-resilient, when $f\geq h$, where $h=n-k-f$.
Proof sketch of Theorem 3
Proof.
When $f+k\geq q$, Byzantine and rational validators can successfully construct valid certificates for an arbitrary number of conflicting blocks and provide them to clients. For any stake $D$ that rational validators have locked before the run, there is a configuration in which each rational validator can double-spend more than $D$ coins. That is because the number of clients is arbitrary and unknown before the run. Even if the conflicting blocks are detected and the misbehaving validators get slashed, the attack is still beneficial for rational validators. ∎
By setting $k=0$ in Theorem 3, we conclude that no $q$-commitable protocol can be $(n,f)$–resilient when $f=h$. Furthermore, Theorem 3 establishes that even if a $q$-commitable protocol remains $(n,f)$–resilient for $f=h-1$, we cannot replace a single correct validator to a rational one, implying that resilience breaks down when $f+k\geq 2n/3$. This result is formalized in Corollary 3.2.
Corollary 3.2.
Consider any $q$-commitable protocol $\Pi$ that is $(n,f)$–resilient for $f=h-1$ where $h=n-f$ is the number of correct validators. For any $k\geq 1$, $\Pi$ is not $(n,k,f)$–resilient under our model in the synchronous setting.
In other words, no quorum-based SMR protocol can achieve $(n,f+1,f)$–resilience, even in the synchronous model. Therefore, quorum-based SMR protocols are not resilient for $f+k\geq 2n/3$.
Finality Latency
Theorem 3 establishes that if participation rewards are used to incentivize rational validators, no protocol can ensure finality during a $\Delta^{*}$-bounded window when $f\geq\lceil\frac{h}{2}\rceil$. That is because Byzantine and rational validators can still produce valid quorum certificates for conflicting blocks within $\Delta^{*}$, with each rational validator double-spending an amount exceeding their stake. This attack remains profitable for rational validators, even if they get slashed and forfeit their stake.
{restatable}
theoremresponsiveness Let us assume participation rewards are distributed to validators, silent clients, a static PoS setting, and a synchronous network. No SMR protocol can be both (i) $(n,f)$–resilient $q$-commitable with finality latency $t$, and (ii) $(n,k,f)$–resilient with finality latency $t^{\prime}<t+\Delta^{*}$, when $f\geq\lceil\frac{h}{2}\rceil$, for $h=n-k-f$.
Proof sketch of Theorem 3
Proof.
The adversary can perform the double-spending attack illustrated in Lemma 8, within the last $\Delta^{*}$. This attack can be beneficial for rational validators since for any value of stake $D$ they lock before the run, there is an input of conflicting transactions $T_{1}$ and $T_{2}$ where rational validators transfer $D+\epsilon,\epsilon>0$ coins to clients in $C_{1}$ and $C_{2}$ respectively, and therefore they can double-spend more than $D$ coins. Even if they are slashed and forfeit their stake, the attack is still beneficial.
Now, let us fix any value of stake $D$. Similar to Theorem 3, a safety violation remains possible within $\Delta^{*}$. That is because both cases assume the same resilience against Byzantine adversaries and operate under similar network conditions during a network partition in $\Delta^{*}$.
We only need to argue that we can construct a configuration where validators can spend more than $D$ coins in $T_{1}$ and $T_{2}$. To show that, we argue that validators can spend their funds multiple times. This is feasible because they can reclaim their initial funds through participation rewards, allowing them to spend them again, and repeat that sufficiently many times during the period $\Delta^{*}$.
According to Theorem 3, if we run an $(n,f)$–resilient quorum-based SMR protocol $\Pi$ with $f\geq\lceil\frac{h}{2}\rceil$ validators, it is impossible to finalize every transaction before $\Delta^{*}$ elapses.
4 CoBRA
In this section, we outline how to transform any $(n,f)$-resilient $(2f+1)$-commitable SMR protocol $\Pi$ into a $(n,k,f^{*})$-resilient proof-of-stake SMR protocol for $n=3f+1$. First, we consider $f^{*}\leq f$ Byzantine validators, preserving the security guarantees of the underlying SMR protocol while allowing the model to accommodate more rational nodes. To accommodate at least one rational validator ($k\geq 1$), we assume a synchronous network model with message delays bounded by $\Delta^{*}$, as required by Theorem 3. Under these conditions, we provide an $(n,k,f^{*})$-resilient protocol with $k+f^{*}\leq 2f$ and $f^{*}\leq f$, which is optimal as shown in Theorem 3.
4.1 CoBRA Overview
Our protocol builds on an underlying SMR protocol $\Pi$ and ensures that validators finalize transactions securely. The key challenge is that rational validators may attempt to maximize their gains by colluding with Byzantine validators to double-spend or censor transactions. Furthermore, as proven in Theorem 3, no protocol can ensure finality before the maximum message delay elapses. To address these challenges, our protocol introduces the following components:
  • Safety. A finality gadget that defines confirmation rules for validators and clients, and detects equivocations; together with a slashing mechanism that penalizes validators upon evidence of equivocation. These components ensure that rational validators have no incentive to fork, as such behavior yields no reward.
  • Liveness. A reward scheme that incentivizes the participation of rational validators and mitigates censorship attacks.
  • Efficiency. A strong chain mechanism that enables higher-volume finalization during periods of high participation.
4.2 SMR protocol
We model the SMR protocol $\Pi$ as a black-box abstraction that proceeds in rounds. In some rounds, one or more validators may propose blocks. Validators can vote on proposed blocks either explicitly, by casting votes, or implicitly, e.g., by extending a block with a new proposal. A block is considered confirmed in $\Pi$ once a validator obtains a valid quorum certificate for it. Whenever a validator $p$ obtains a valid certificate $QC(b)$ for block $b$ at height $h$, the protocol outputs to $p$: deliver$(QC(b),b,h)$.
This abstraction captures a broad class of $q$-committable blockchain protocols, including leader-based, chained, DAG-based, and leaderless designs. For instance, in leader-based protocols, designated leaders propose blocks in certain rounds, followed by one or more voting rounds leading to block finalization.
We classify the action space of validators during the execution of $\Pi$ into two categories: safety-preserving and liveness-preserving actions.
Safety-Preserving Actions
Rational validators may collude with byzantine validators to fork the system. More specifically, although the SMR protocol $\Pi$ may output a single block $b$ with a valid certificate $QC(b)$ to $p_{i}$, before the message delay elapses, rational and Byzantine validators can produce a conflicting block $b^{\prime}$ with its own valid certificate $QC(b^{\prime})$ by obtaining signatures from correct validators that have not yet witnessed $QC(b)$.
We assume a broadcast-based SMR protocol $\Pi$ [^3] Otherwise, validators in our construction can broadcast the valid certificates to propagate the required information., ensuring that all conflicting blocks with valid certificates are detected by every correct validator within $2\Delta^{*}$  [^4] Clients only accept blocks with $12f+$ 1 finality votes according to the client confirmation rule. Validators commit finality votes only for blocks with valid certificates that meet conditions specified by the protocol. Hence, rational validators cannot benefit from hiding votes from correct validators. Therefore, we consider only certificates witnessed by at least one correct validator.. We further analyze this attack in Appendix A.
Liveness-Preserving Actions
Since $\Pi$ is $(n,f)$-resilient, it guarantees liveness as long as there are at most $f$ Byzantine validators and at least $n-f$ correct validators. Given that the set of correct and rational validators is $f+1+k\geq n-f$, the protocol remains live if the $k$ rational validators participate in liveness-preserving actions. In the setting of $q$–committable blockchains, however, a valid certificate is sufficient to commit a block. Therefore, it is enough to guarantee that there exist specific intervals during which rational validators participate in the liveness-preserving steps.
Liveness-preserving actions include (i) proposing a valid block when it is their turn according to $\Pi$, (ii) voting for valid blocks proposed by others, and (iii) engaging in additional actions such as view-change protocols specified by $\Pi$. Assuming synchronous periods are sufficiently long, these conditions guarantee that at least one correct validator will eventually propose and finalize a valid block. As a result, any valid transaction observed by all correct validators will eventually be included in the ledger.
In contrast, rational validators may choose to abstain from liveness-preserving actions if offered a sufficiently high bribe by the adversary to censor specific transactions. We further elaborate on liveness attacks in Appendix A and provide examples of bribing mechanisms in Section 7.
4.3 Finality Gadget
The finality gadget ensures that validators finalize only transactions that cannot be reversed and enables them to provide inclusion proofs to clients.
We present the finality gadget in Algorithm 1. Moreover, below, we describe the validator and client protocols for accepting a transaction. Since clients do not monitor the protocol execution, their confirmation rule differs from that of validators. As a result, validators may finalize transactions based on protocol guarantees, whereas clients rely on additional proofs to verify finality.
Validator confirmation
When a validator $p_{i}$ witnesses a valid certificate $QC(b)$ for a block $b$, it stores the certificate (line 7) along with the block and the current local timestamp (line 8). Then $p$ initiates the block finalization function.
Since conflicting blocks may exist but remain unwitnessed within the last $2\Delta^{*}$, validators employ the Stake-bounded finalization rule. More specifically, each validator $p$ traverses its local ledger and using the timestamp of the blocks and its local clock, determines which blocks are $2\Delta^{*}$–old (output $2\Delta^{*}$ ago), or $2\Delta^{*}$–recent (output within the last $\Delta^{*}$).  [^5] To measure elapsed time, $p_{i}$ relies on its local clock rather than requiring synchronized clocks (lines 8, 18). Then, $p$ finalizes blocks as follows.
$2\Delta^{*}$–old Blocks
If no conflicting block has been detected, validators finalize any block $b^{\prime}$ that was output by $\Pi$ at least $2\Delta^{*}$ ago (lines 19 and 20). If a conflicting block existed, it would have been observed by this time.
$2\Delta^{*}$–recent Blocks
For blocks output within the last $2\Delta^{*}$, validators cannot guarantee the absence of conflicting chains. Instead, each validator finalizes only a subset of the most recent blocks, ensuring that the total finalized transaction volume does not exceed $C=D$ coins (lines 22 and 23). This guarantees that any rational validator engaging in an attack can gain at most $D$ coins. For intuition, we provide a tight example in Figure 1, where the adversary partitions all correct validators to create $f+1$ conflicting ledgers before the bound $\Delta^{*}$ elapses. Assuming Byzantine validators claim no rewards and instead use them to bribe the $f$ rational validators they need to successfully double-spend, the $f$ rational validators share a total of $f\times D$ double-spent coins, resulting in a per-validator gain of $D$.
Figure 1: Forking attack.
At round $t$, a subset $\mathcal{F}_{R}$ of $2f$ misbehaving validators (Byzantine and rational) fork the system. The adversary partitions correct validators into $f+1$ distinct sets and collects their signatures to create conflicting blocks with valid certificates. Each correct validator finalizes blocks with transaction volume up to $C=D$ within the last $\Delta^{*}$. The misbehaving rational validators share the rewards and, consequently, each rational validator is able to double-spend $D$ coins.
Client Confirmation
After finalizing a block $b$, validators include finality votes for $b$ in subsequent blocks. A block $b$ is considered finalized for clients once at least $12f+$ 1 finality votes for $b$ or a block extending $b$ have been recorded.
To confirm the inclusion of a transaction $tx$, validators provide clients with a proof consisting of:
  • An inclusion proof (e.g., a Merkle proof) for $tx$ in block $b$.
  • A finality proof containing at least $12f+$ 1 finality votes for $b$ or a descendant block.
This ensures that clients can securely verify the transaction’s finalization, ensuring that the SMR protocol is certifiable.
4.4 Rewards and Penalties
Slashing Mechanism
To ensure misbehaving validators are penalized, we implement a two-stage withdrawal process [16]:
  • Validators must stake $D$ coins to participate in $\Pi$.
  • Withdrawals are only allowed after a waiting period $\Delta_{W}>2\Delta^{*}$, during which validators cannot participate in the protocol.
  • If a validator is detected signing conflicting blocks, it is blacklisted (line 13), preventing permanently its withdrawal.
All correct validators will detect the equivocations within $2\Delta^{*}$. Thus, misbehaving validators cannot reclaim their stake. This disincentivizes rational validators from forking.
Recovery
If a fork is detected, honest validators halt participation in the underlying SMR protocol $\Pi$. As a result, the system stalls due to an insufficient number of active validators to ensure liveness. However, as we will prove, forking is not a profitable strategy for rational validators and thus does not occur in equilibrium. We briefly discuss recovery and reconfiguration procedures following misbehavior, such as rational validators misestimating their utility, in Appendix 7.
Algorithm 1 Finality Gadget
[…]
5on SMR.deliver($QC(b),b,h$) do
6    if $quorumstore[h]=\bot$ then
7       $quorumstore[h]\leftarrow QC(b)$
8       $ledger\leftarrow ledger\cup(b,localtime)$
9       finalize-blocks()
10    else
11       $(block,\_\;)\leftarrow ledger[h]$
12       if $block\neq b$ then
13          slash($quorumstore[h]\cap QC(b)$)
14          $quorumstore[h].add(QC(b))$
15
16procedure finalize-blocks
17    $suffixvalue=0$
18    for $(block,blocktime)\in ledger$ from suffixindex to $|ledger|$ do
19       if $blocktime\leq localtime-2\Delta^{*}\land|blacklist|=\emptyset$ then .
20          $block.finalize$
21          $suffixindex++$
22       else if $suffixvalue+block.value\leq C\land|blacklist|=\emptyset$ then
23          $block.finalize$
24          $suffixvalue\leftarrow suffixvalue+block.value$
25       else
26          return
27
28procedure slash($validators$)
29    $blacklist\leftarrow blacklist\cup validators$
30
Rewards
Participation rewards are sourced from transaction fees and distributed to validators based on their contributions to block finalization. Rewards are proportional to the fraction of blocks with valid certificates proposed by each validator and are distibuted every $n$ rounds. Specifically, we define the interval of rounds $[kn+1,(k+1)n]$ for $k\in\mathbb{Z}$ as the leader window $W_{k}$. After the end of the window $W_{k}$ the total transaction fees from finalized transactions $F_{W_{k}}$, are allocated based on the proportion of unique proposers $P_{W_{k}}$ relative to $n$: $R_{W_{k}}=F_{{W_{k}}}\cdot\dfrac{P_{W_{k}}}{n}.$ To ensure that correct validators successfully propose blocks, rewards are distributed only if $P_{W_{k}}\geq n-f$, i.e., at least $n-f$ unique validators submit proposals within $n$ rounds. This guarantees censorship resilience, meaning that rational validators have no incentive to accept bribes from Byzantine validators to censor specific transactions or remain silent when a correct block proposer is present.
4.5 Extending Finalization with Strong Chains
Recall that validators finalize transaction volumes up to $D$ coins within a time window $2\Delta^{*}$. This is the lower bound in our threat model. When the system is not under attack (i.e., fewer faulty nodes), blocks often receive more than $12f+$ 1 signatures. We leverage the certificates with additional signatures to further limit forks and finalize larger transaction volumes without waiting for $2\Delta^{*}$.
To achieve this, we introduce the concept of $i$-strong chains [17]. A ledger is $i$-strong if it contains valid certificates signed by $2f+1+i$ unique validators within $2\Delta^{*}$. With at least $1+i$ correct validators, the maximum number of forks decreases to $f-i$, allowing secure finalization of higher transaction volumes. In particular, when a certificate has $12.5f+$ 1 signatures, we can define a strongest chain that can finalize infinite transaction volume (regardless of $D$).
Definition 9 ($i$-strong chain).
Consider a validator $p$ and its local data structure $\text{ledger}^{p}$ (Algorithm 1, line 1). Let $S$ represent the set of distinct signatures from validators included in the valid certificates of blocks within the set of blocks in $\text{ledger}^{p}$ that were output by the SMR protocol $\Pi$ within the last $2\Delta^{*}$. We say that $\text{ledger}^{p}$ is an $i$-strong chain, where $i=|S|-2f-1$.
Strong Chain finalization rule
Validators that monitor an $i$-strong chain for $i>f/4$, can finalize transaction volume up to $C=\frac{f}{f-i}D$. Otherwise, validators finalize $C=D$ coins. Notice that for $i=f$, $C\to\infty$ because all validators participate in that ledger, allowing the finalization of any block with a valid certificate.
$C=\begin{cases}\frac{f}{f-i}D,&\text{if }i>\frac{f}{4},\\ D,&\text{else.}\end{cases}$
Strongest Chain finalization rule
We additionally define the strongest chain, as an i-strong chain with $i>\frac{f+1}{2}$. More than half of the correct validators participate in the strongest chain. This guarantees that only one strongest chain can exist within $2\Delta^{*}$, and all correct validators will shift on it after detecting a fork, under this rule:
$C=\begin{cases}\infty,&\text{if }i>\frac{f+1}{2},\\ \frac{f}{f-i}D,&\text{if }\frac{f}{4}<i\leq\frac{f+1}{2},\\ D,&\text{otherwise.}\end{cases}$
5 Analysis
We will prove that CoBRA is both a safety-preserving and liveness-preserving transformation. To this end, let $S_{f}\subseteq S_{\mathsf{bad}}$ denote the set of strategy profiles that cause a safety violation (possibly alongside a liveness violation), and define $S_{l}=S_{\mathsf{bad}}\setminus S_{f}$ as the set of strategy profiles that lead solely to liveness violations.
For each validator $p$, let $S_{f}^{p}$ (or $S_{l}^{p}$) be the set of individual strategies $s_{p}$ such that there is a strategy profile $s=(s_{p},s_{-p})\in S_{f}\textit{ (or $S_{l}$)}$ where $p$ actively contributes to the safety (or liveness) violation, and let $S_{f}^{-p}$ (or $S_{l}^{-p}$) denote the set of strategies $s_{-p}$ of all other validators in such profiles.
Safety-Preserving Transformation
First, we will show that regardless of the actions of the rest validators it is not beneficial for rational validators to engage in a forking attack. In that way, we demonstrate that for every validator $p$, and any strategy $s_{p}\in S^{p}_{f}$, there exists an alternative strategy $s_{p}^{\prime}\in S^{p}_{good}$ that weakly dominates $s_{p}$.
Liveness-preserving transformation
To establish that CoBRA is a liveness-preserving transformation, we show that at any round there will be future “good” leader windows during which at least one correct validator successfully proposes a block. Since this prevents liveness violations, we conclude that for every validator $p$ and any strategy $s_{p}\in S^{p}_{l}$, there exists an alternative strategy $s_{p}^{\prime}\in S^{p}_{\mathsf{good}}$ that weakly dominates $s_{p}$.
Combining both guarantees demonstrates that the composition of $\Pi$ and CoBRA forms a rationally robust protocol. To formalize this, we define the utility function of the rational validators.
Utility of rational validators
To model dynamic behavior over intervals of rounds, we define the global history up to round $r$ as $H^{r}$, and the local view of player $i$ at that round as $H_{i}^{r}\subseteq H^{r}$. Let $S_{i}^{I}$ be the set of strategies available to player $i$ over an interval $I$, where each strategy is a function of their local view $H_{i}^{r}$ at each round $r\in I$. The overall strategy space for the interval is $S^{I}=\prod_{i\in[n]}S_{i}^{I}$, and the utility function of player $i$ over $I$ is given by $u_{i}^{I}:S^{I}\to\mathbb{R}$.
We now consider the set of strategies available to a rational validator $p$ over some interval $I$. First, $p$ may choose to participate in a coalition that attempts to fork the protocol. [^6] We assume a worst-case scenario: if $p$ chooses to join a coalition $\mathcal{F}$, all other validators in $\mathcal{F}$ are colluding. Otherwise, if $p$ does not engage in a forking attack, it may either follow actions that preserve liveness or abstain from doing so.
Forking. To perform this attack, a rational validator $p$ can join a coalition $\mathcal{F}\in\mathcal{F}_{R}$, where $\mathcal{F}_{R}$ consists of all the susbets of Byzantine and rational validators sufficienlty large to create a fork. We denote the partipication of $p$ in a coalition $\mathcal{F}$ during the interval $I$ with the indicator variable $b^{I}_{\mathcal{F},p}$. The function $\mathcal{M}$ gives an upper bound of the coins that can be double-spent during the attack:
$\mathcal{M}(\mathcal{F},I)=\frac{\gamma\cdot C(I)}{R(\mathcal{F})},$
where $\gamma$ is the maximum number of forks that can be created by the coalition, $C(I)$ is an upper bound of coins finalized per fork during $I$ as specified by the protocol, and $R(\mathcal{F})$ returns the number of rational validators in the coalition [^7] To provide stronger guarantees, we assume that all the rewards from double-spending are distributed exclusively and equally to the rational validators..
Moreover, we denote the participation of $p$ in at least one coalition $\mathcal{F}\in\mathcal{F}_{R}$ by the indicator variable $b^{I}_{p}$. Let $\mathbf{b}^{I}_{p}$ be the vector whose elements are all indicator variables $b^{I}_{\mathcal{F},p}$ for every $\mathcal{F}\in\mathcal{F}_{R}$ and let $P$ be the slashing penalty imposed on $p$ (which is equal to the stake $D$).
Liveness-preserving actions. If validator $p$ does not join any coalition, it may instead follow the protocol’s liveness-preserving actions—such as proposing valid blocks when selected as leader, voting for valid blocks, participating in view-change procedures, and issuing finality votes for blocks committed to its local ledger.
We denote by $\sigma^{I}_{p}$ the strategy of validator $p$ over the interval $I$ with respect to these liveness-preserving actions. Let $R(\sigma^{I}_{p})$ denote the total participation rewards that $p$ earns by following $\sigma^{I}_{p}$, and let $B(\sigma^{I}_{p})$ denote any bribes received from an adversary to refrain from signing specific blocks.
Finally, the utility of $p$ in the interval $I$ while following the strategy $s_{p}=(\mathbf{b}^{I}_{p},\sigma^{I}_{p})$ is expressed as:
$u_{p}^{I}(s_{p},s_{-p})=\begin{cases}\sum_{\mathcal{F}\in\mathcal{F}_{R}}b^{I}_{\mathcal{F},p}\cdot\mathcal{M}(\mathcal{F,I})-P&\text{if $b^{I}_{p}=1$},\\ R(\sigma^{I}_{p})+B(\sigma^{I}_{p})&\text{otherwise}.\end{cases}$
5.1 Lemmas for proving safety
In Lemma 1, we prove that once a correct validator finalizes a $2\Delta^{*}$-old block $b$, no conflicting block can exist. Thus, misbehaving validators can only double-spend in $2\Delta^{*}$-recent blocks.
Lemma 1.
Assume a correct validator finalizes a block $b_{1}$ that is $2\Delta^{*}$-old. Then, no correct validator will ever finalize a block $b_{2}$ conflicting to $b_{1}$.
Proof.
Assume, for the sake of contradiction, that the following scenario occurs. A correct validator $p_{i}$ includes block $b_{1}$ in its local ledger at time $localTime_{b_{1}}$ and triggers $b_{1}.\mathsf{finalize}$ at time $localTime$, where $localTime\geq localTime_{b_{1}}+2\Delta^{*}$. Furthermore, another correct validator $p_{j}$ triggers $b_{2}.\mathsf{finalize}$ for a block $b_{2}$ that conflicts with $b_{1}$.
By assumption, at $localTime_{b_{1}}$, $p_{i}$ broadcasts the valid certificate $QC(b_{1})$ for block $b_{1}$. Thus, $p_{j}$ receives $QC(b_{1})$ at some time $t^{*}\in\{localTime_{b_{1}},\dots,localTime_{b_{1}}+\Delta^{*}\}$. We distinguish two cases:
Case 1: $p_{j}$ triggers $b_{2}.\mathsf{finalize}$ at time $t<t^{*}$. Then $p_{j}$ must have received $QC(b_{2})$ at some time $t^{\prime}\leq t<t^{*}$. Since $p_{j}$ broadcasts $QC(b_{2})$, validator $p_{i}$ receives it by some time $localTime_{b_{2}}\leq t+\Delta^{*}<localTime_{b_{1}}+2\Delta^{*}$. At time $localTime_{b_{2}}$, $p_{i}$ detects the equivocation and slashes the misbehaving validators (line 13), resulting in a non-empty blacklist: $QC(b_{1})\cap QC(b_{2})\neq\emptyset$. Therefore, at time $localTime$, the finalization condition (line 19) is not satisfied, and $p_{i}$ will not finalize $b_{1}$ (line 20), contradicting the assumption.
Case 2: $p_{j}$ triggers $b_{2}.\mathsf{finalize}$ at time $t\geq t^{*}$. At time $t$ $p_{j}$ has received both $QC(b_{1})$ and $QC(b_{2})$. Let $t_{b_{2}}$ and $t_{b_{1}}$ be the times at which $p_{j}$ received $QC(b_{2})$ and $QC(b_{1})$, respectively, and assume without loss of generality that $t_{b_{1}}\geq t_{b_{2}}$. At time $t_{b_{1}}$, $p_{j}$ detects the equivocation and slashes the misbehaving validators (line 13), resulting in $QC(b_{1})\cap QC(b_{2})\neq\emptyset$. Therefore, at time $t$, the finalization condition (line 19) is not satisfied, and $p_{j}$ will not finalize $b_{2}$ (line 20), leading to a contradiction.
In Lemma 2, we show that when validators employ the stake-bounded finalization rule, any rational validator engaging in a forking attack can gain at most $D$ coins.
Lemma 2.
Assume that validators employ the stake-bounded finalization rule and that at least one coalition of validators forks the system in some round.
  1. 1.
    Upper bound of double-spending: Any rational validator participating in the attack can double-spend at most an amount of coins equal to the stake $D$.
  2. 2.
    Tight example: There is a configuration where a rational validator in the coalition double-spends exactly $D$ coins.
Proof.
(1) Upper bound of double-spending. Misbehaving validators can double-spend only coins in conflicting blocks that are accepted by clients. Any inclusion proof $\pi_{tx}$ for a transaction $tx$ included in a block $b$ is accepted by some client $c$ only after witnessing at least $12f+$ 1 finality votes for $b$. Therefore at least one correct validator $p_{i}$ has finalized block $b$. Since correct validators do not trigger $finalize$ for conflicting blocks, and at least one correct validator is required to finalize a block, there can be at most $f+1$ conflicting ledgers. Moreover, according to Lemma 1, there can only be $2\Delta^{*}$-recent conflicting blocks.
Assume the worst case, where there are $f+1$ conflicting ledgers and let us fix any rational validator $p$ participating in $f+1-m$ conflicting ledgers $L_{i\in[f+1-m]}$ with $m\in[f+1]$. The coins that misbehaving validators double-spend on these conflicting ledgers is upper bounded by the total coins spent in conflicting blocks finalized by at least one correct validator. To bound this, we fix $L_{i}$ for some $i\in[f+1-m]$, and upper bound the spent coins in $2\Delta^{*}$-recent blocks that are finalized by at least one correct validator.
Let $Suffix^{t^{\prime}}_{L_{i}}$ be the set of all $2\Delta^{*}$-recent blocks with valid certificates in $L_{i}$ until time $t^{\prime}$. For every block $b\in Suffix^{t^{\prime}}_{L_{i}}$, $p_{i}$ updates the variable $suffixvalue$ with the total output sum of transactions in the block $b$ and only finalizes $b$ if $suffixvalue$ is at most $C=D$ (line 22). Therefore, the amount of coins spent in $L_{i}$ is at most $D$.
Now, we will upper bound the total amount of coins validator $p$ can double-spend in all the $f+1-m$ conflicting ledgers. First, at most $m+1$ correct validators participate in each ledger. We prove that by reaching a contradiction. Assume there is a ledger with at least $m+2$ correct validators. Then, in the remaining $f-m$ ledgers, there must be at least one correct validator. Therefore, there are, in total, at least $m+2+f-m=f+2$ correct validators in the system. We reached a contradiction. Since there are at most $f$ byzantine and $m+1$ correct validators per ledger, at least $f-m$ rational validators participate in each ledger. Therefore, at least $f-m$ rational validators can double-spend at most $C=D$ coins in $f-m$ ledgers, and since they distribute the benefit equally, each validator double-spends $max_{m\in[f+1]}\frac{f-m}{f-m}D=D$ coins.
(2) Tight example. Consider a run with $f$ byzantine, $f$ rational, $f+1$ correct validators. Moreover, consider as input a set of transactions $Tx=\cup_{i=1}^{f+1}Tx_{i}$ constructed as follows. There is a partition of clients in distinct sets $C=\cup_{i=1}^{f+1}C_{i}$ s.t. for each $i$, in $Tx_{i}$, each of the $f$ rational validators transfers $D$ coins to a subset of clients in $C_{i}$.
At some time, a malicious (Byzantine or rational) validator $p$ is the block proposer. Within the last $\Delta^{*}$, the adversary partitions the $f+1$ correct validators in the distinct sets $H_{i}$ for $i\in[f+1]$. Now, $p$ constructs the blocks $b_{i\in[f+1]}$, where each $b_{i}$ includes the transactions in $Tx_{i}$. Validator $p$ sends the block $b_{i}$, only to the correct node in $H_{i}$ and successfully generates a valid certificate for $b_{i}$ along with the respective finality votes which will present to the clients in $C_{i}$. Therefore, each rational validator spends the $D$ coins in the $f+1$ conflicting ledgers. In total, each rational validator double-spends $\frac{f*D}{f}=D$ coins.
Now, we argue that any misbehaving validator will be detected and forfeit their stake.
Lemma 3.
Assume a coalition of validators $\mathcal{F}$ forks the system. No validator in $\mathcal{F}$ can withdraw its stake.
Proof.
Assume that the validators in $\mathcal{F}$ fork the system at time $t$ [^8] We refer to the first round when distinct correct validators observe conflicting blocks with valid certificates. As before, we use the local clock of any validator to refer to time elapsed between events. resulting in $j\in[f+1]$ conflicting ledgers $L_{i\in\{1,...,j\}}$. Let us fix any validator $p\in\mathcal{F}$. We will show that $p$ cannot withdraw its stake in any of the conflicting ledgers.
Consider the first ledger $L_{i}$ in which validator $p$’s withdrawal request $R_{p}$ is included in a block $b$ at time $t^{\prime}$. We argue that $t\leq t^{\prime}+\Delta^{*}$. Suppose, for contradiction, that $t>t^{\prime}+\Delta^{*}$. Since block $b$ has a valid certificate, it must have been witnessed by at least one correct validator $p_{j}$, who then broadcasts $R_{p}$ at time $t^{\prime}$. Therefore all correct validators receive $R_{p}$ by time $t^{\prime}+\Delta^{*}$ and ignore the participation of $p$ at time $t$. This contradicts the assumption that $p$ participated in the forking attack at time $t$.
To complete the withdrawal, a second transaction $R_{p,s}$ must be committed to $L_{i}$ at a time $>t^{\prime}+2\Delta^{*}$. Since $L_{i}$ is the first ledger to include $R_{p}$, the same condition applies to any other conflicting ledger, namely the withdrawal can only be finalized after $>t^{\prime}+2\Delta^{*}$. However, since correct validators broadcast valid certificates, every correct validator will witness all conflicting ledgers at some time $t^{*}\leq t+\Delta^{*}\leq t^{\prime}+2\Delta^{*}$. As a result, they will detect equivocation and blacklist $p$, preventing the finalization of $R_{p,s}$. Thus, $p$’s withdrawal cannot succeed in any of the conflicting ledgers.
Strong(-est) Chain Finalization Rule
We provide only sketch proofs for the Strong and Strongest finalization rules (see Section 4.5), and defer the full analysis to Appendix B.
We leverage additional signatures to constrain the number of possible forks and, consequently, the total amount of coins that can be spent in conflicting blocks. Specifically, we bound the coins validators can finalize within $\Delta^{*}$ based on the number of signatures in their local chain. In the most optimistic scenario, with $65n/$ 6 signatures, we show that, regardless of other forks, only one strongest chain can exist (Lemma 10), to which all validators will transition within $\Delta^{*}$. In both cases, we prove that rational validators attempting to equivocate across the other forks can gain at most $D$ coins from conflicting blocks (Lemma 9, Lemma 11). Given our slashing mechanism (Lemma 3), rational validators cannot increase their profit through equivocation, ensuring that the strongest chain finalization rule preserves safety.
Finally, in Lemmas 45, we conclude that rational validators have no incentive to engage in forking attacks, as any potential gain does not exceed their staked amount, which is subject to slashing.
Lemma 4.
When correct validators use the i) Stake-bounded finalization rule, or ii) the Strong Chain finalization rule, or iii) the Strongest Chain finalization rule, for any rational validator, following the protocol weakly dominates participating in a coalition to fork the system.
Proof.
Let us fix any rational validator $p$ and compute its utility to join coalitions forking the system. In Lemmas 2911, we showed that by performing the attack, $p$ can double-spend at most $D$ coins when validators employ i) finalization rule presented in 4, ii) strong ledger finalization rule, iii) strongest finalization rule respectively. According to Lemma 3, $p_{j}$ cannot withdraw its stake $D$ before the conflicting ledgers are detected. Moreover, upon detecting equivocation the system halts and thus misbehaving validators cannot utilize further bribes. Therefore, for any strategy $s_{p}\in S^{p}_{\mathsf{f}}$ and $s_{-p}\in S^{-p}_{\mathsf{f}}$, the utility of $p$ is $u_{p}(s_{p},s_{-p})\leq 0$, while following the protocol yields non-negative utility.
To show that following the protocol weakly dominates the strategy of forking the system, we will construct a configuration where $p$ chooses a strategy $s_{p}\in S^{p}_{\mathsf{f}}$ and fix the rest strategies $s_{-p}\in S^{-p}_{\mathsf{f}}$ such that $u_{p}(s_{p},s_{-p})<0$ (this is sufficient since following the protocol has non-negative utility). To this end, suppose that $p$ participates in a coalition $\mathcal{F}$ consisting of $f$ Byzantine and $\lceil\frac{f+1}{2}\rceil$ rational validators. Moreover, consider a subset $H_{1}$ of $\lceil\frac{f+1}{2}\rceil$ correct validators, and a subset $H_{2}$ distinct to $H_{1}$ that consists of $\lfloor\frac{f+1}{2}\rfloor$ correct validators. Validators in $\mathcal{F}$ construct to conflicting ledgers $T_{1}$ and $T_{2}$ as described in Lemma 8. The utility of $p$ in that case is $u_{p}(s_{p},s_{-p})\leq\frac{D}{\lceil\frac{f+1}{2}\rceil}-D<0$.
Lemma 5.
A fork of the system will never occur.
Proof.
For the sake of contradiction assume that two distinct correct validators $p_{1}$, $p_{2}$ finalize the conflicting blocks $b_{1}$ and $b_{2}$ respectively. Validator $p_{1}$ finalized $b_{1}$ after witnessing the respective valid certificate $QC(b_{1})$, and $p_{2}$ finalized $b_{2}$ after witnessing $QC(b_{2})$. By quorum intersection there exists a validator $p$ that is not Byzantine and has voted in both $QC(b_{1})$ and $QC(b_{2})$. Since correct validators do not equivocate, $p$ must be rational. However, according to Lemma 4, for rational validators, following the protocol weakly dominates participating in a coalition that forks the system. So, $p$ will not sign for conflicting blocks and thus we reached a contradiction.
5.2 Lemmas for proving liveness
Since $\Pi$ is $(n,f)$-resilient and $f+1+k\geq n-f$, liveness is ensured as long as rational validators participate in liveness-preserving actions during some intervals. In Lemma 6, we show that there exist windows during which it is a dominant strategy for rational validators to follow the protocol specifications. Building on this, Lemma 7 concludes that some correct validator will successfully propose a block that gathers a valid certificate and finality votes.
Lemma 6.
For any round $r$, there is a leader window $W_{k}$ starting at some round $r^{\prime}\geq r$, such that following the protocol in every round of $W_{k}$ is a weakly dominant strategy for any rational validator.
Proof.
Assume for contradiction that in every window $W_{m}$ after round $r$, there exists at least one rational validator who prefers a strategy that deviates from the protocol. Since following the protocol yields a non-negative utility, any such deviating strategy must yield strictly positive utility during window $W_{m}$. Rational validators can deviate as follows:
Forking. From Lemma 4, any strategy in the strategy profile $S^{p}_{f}$ (i.e., where $p$ actively contributes to a fork), is weakly dominated by following the protocol. Therefore, no rational validator will engage in forking attacks.
Only abstaining from liveness actions. A rational validator will only choose to abstain if the adversary offers a bribe using their participation rewards from previous rounds that outweighs the utility from honest participation.
Suppose the adversary is able to bribe at least one rational validator in every window $W_{m}$ following round $r$. In such windows, at most $2f$ validators can successfully propose valid blocks. Since participation rewards are distributed only when at least $12f{+}$ 1 distinct validators succeed, no rewards are paid out in these bribed windows. Because the adversary’s budget is finite, it is eventually depleted and the adversary cannot bribe any rational validator anymore. Therefore, there exists a future window $W_{k}$ (starting at some round $r^{\prime}\geq r$) such that, for any validator $p$, any strategy $s_{p}\in S_{p}$, and any $s_{-p}\in S_{-p}$ under which $p$ abstains from liveness actions during $W_{k}$, we have $u_{p}(s_{p},s_{-p})\leq 0$.
To show that following the protocol during $W_{k}$ is a weakly dominant strategy, fix any rational validator $p$, and consider two strategies: $s_{1}$, where $p$ follows the protocol during $W_{k}$, and $s_{2}$, where $p$ abstains from liveness actions during $W_{k}$ (we have already established the result for forking in 4). Let $s_{-p}$ be a strategy profile in which at least $2f$ other validators follow the protocol during $W_{k}$. In the combined strategy profile $s_{a}=(s_{1},s_{-p})$, at least $12f+$ 1 validators follow the protocol, satisfying the reward condition, and $p$ receives the corresponding participation rewards. In contrast, under $s_{b}=(s_{2},s_{-p})$, only $2f$ validators follow the protocol, so no rewards are distributed. Hence, $u^{W_{k}}_{p}(s_{a})>u^{W_{k}}_{p}(s_{b})$.
Lemma 7.
For any round $r$, there exists a round $r^{*}\geq r$ at which a correct validator proposes a block $b$ such that:
  1. 1.
    $b$ is finalized by every correct validator, and
  2. 2.
    $12f+$ 1 finality votes for $b$ or a subsequent block are included in blocks with valid certificates.
Proof.
According to Lemma 6, for any round $r$, there exists a round $r^{\prime}\geq r$ such that a time window $W_{k}$ begins, during which it is a weakly dominant strategy for rational validators to follow the protocol. Consequently, some block $b$ proposed by an honest validator during $W_{k}$ will obtain a valid quorum certificate $QC(b)$ and receive at least $12f+$ 1 finality votes (either for $b$ or for one of its descendants).
5.3 Final Theorems
Theorem 5.1.
Validators running the composition of CoBRA and $\Pi$ agree in a total order.
Proof.
Safety. In Lemma 5, we show that validators will never finalize conflicting blocks.
Liveness. Consider any transaction $tx$ witnessed by all correct validators at some round $r$. According to Lemma 7, there is a round $r^{*}\geq r$ where a correct validator $p$ will successfully propose a valid block $b$. If $tx$ has not already been included in a valid block proposed before round $r^{*}$, then, $p$ will include $tx$ at block $b$.
Theorem 5.2.
The composition of CoBRA and $\Pi$ satisfies certifiability.
Proof.
Client-safety. Consider any client $c$. We will show that any transaction $tx$ for which $c$ has accepted an inclusion proof is included in the transaction ledger of at least one correct validator.
To this end, consider the inclusion proof $\pi_{tx}$ for the transaction $tx$. Client $c$ has accepted $\pi_{tx}$ only after receiving a proof showing the inclusion of $tx$ in a block $b$, where $b$ has $12f+$ 1 finality votes. At least one correct validator $p$ has committed a finality vote for $b$. Validator $p$ commits finality votes only for blocks included in its transaction ledger.
Client-liveness. For every transaction $tx$ committed to a block $b$ accepted by all correct validators, any correct validator $p_{i}$ can construct a proof $\pi$ as follows. Validator $p_{i}$ sends an inclusion proof $\pi_{tx}$ of $tx$ in $b$ to a client $c$, along with a proof $\pi_{f}(b)$ consisting of $12f+$ 1 finality votes for $b$ or subsequent block, which is feasible according to Lemma 7. Any client $c$ will accept the proof $\pi$ by definition of the client confirmation rule.
Theorem 5.3.
The composition of CoBRA and $\Pi$ is an $(n,k,f^{*})$-resilient SMR protocol.
Proof.
Validators running the composition of CoBRA and $\Pi$ agree in a total order according to Theorem 5.1. Moreover, in Theorem 5.2 we show that the composition of CoBRA and $\Pi$ satisfies certifiability.
6 Case-Study
In this section, we present an empirical case study to evaluate the practical feasibility of CoBRA. To this end, we analyzed one million consecutive blocks from the Cosmos blockchain. Based on our analysis, over 99% of Cosmos blocks exceeded the $\frac{5}{6}$ participation threshold required by our strongest-chain finalization rule, allowing CoBRA to instantly finalize the vast majority of blocks.
The remaining blocks were produced during periods of lower participation, lasting up to 147 blocks. Given Cosmos’s $\sim$6-second block time, this corresponds to about 15 minutes. With a median daily transaction volume of \$17 million, this 15-minute window corresponds to an average transaction volume of approximately \$180,000. As such, given the minimum slashable stake of active validators of approximately \$470,000, CoBRA can compensate for low-participation periods by instantly finalizing blocks using our stake-based finalization rule.
These high participation rates are not unique to Cosmos. On Ethereum, the average daily participation has also consistently remained above 99% [18]. Thus, we conclude that CoBRA could be seamlessly integrated into production systems like Cosmos or Ethereum, enhancing system resilience without impacting block confirmation times.
7 Limitations and Extensions
Dynamic Validator Setting
We outline how our approach extends to dynamic settings where new validators may join the system:
  • New validator: A validator that wishes to join the system in round $r$ broadcasts a join request and waits until round $r+2\Delta^{*}$ before fully participating. This waiting period allows the validator to detect any past consistency violations.
  • Existing validator: Fix any existing validator $p$ that receives the join request at some round $r^{\prime}\in\{r,r+\Delta^{*}\}$. It then broadcasts any consistency violations it has observed up to round $r^{\prime}$. From that point onward, it treats the new validator as part of the system and interacts with it according to the protocol.
By round $r^{\prime}+\Delta^{*}$, the new validator will have received all consistency violation reports known to validator $p$ for rounds $\leq r^{\prime}$. For future rounds, the new validator learns of inconsistencies through normal protocol participation. Since this argument holds for any existing validator, it is guaranteed that if an honest validator observes a consistency violation, the joining validator will be aware of the violation.
Recovery Protocols
Our solution ensures that rational validators have no positive utility in misbehaving, thus eliminating the need for a recovery protocol, as consistency violations will not occur in our model. However, in practice, it would be beneficial if the system could recover from failures in the event a rational validator incorrectly estimates its utility. For this, we refer to [4], which –assuming the same threshold of correct validators – describes a protocol that merges conflicting branches, compensates affected users from the stake of misbehaving validators, and reconfigures the validator set accordingly. This is possible as CoBRA guarantees that the stake of the rational validators always exceeds the potential utility (i.e. also economic damage) they can extract. Specifically, in the presence of $i\leq f$ conflicting ledgers, misbehaving validators can double-spend in total up to $(i-1)D$ coins. However, upon equivocation, validators in CoBRA detect the misbehavior of at least $f+1$ validators and slash their stake which amounts to $(f+1)D$.
Assessing Transaction Value
An important part of our protocol is assessing the value of transactions in a given block to weigh it against the locked stake. While this is simple in a payment system environment, it is not straightforward in a smart contract environment. While transactions transferring native on-chain currency can easily be evaluated, it is challenging to evaluate the value of ERC-20 tokens, NFTs, or even layer-2 rollup checkpoints.
To address this, clients can include an estimated transaction value in their submissions. If the receiver agrees with the estimate, they accept the stake-based finality rule; otherwise, they wait for the full finality threshold $2\Delta^{*}$ (or a strong chain). However, this mechanism introduces a vulnerability: a malicious client can delay the finality of other transactions by regularly issuing, or falsely declaring, high-value transactions. To mitigate this, we introduce a two-tier system in which high-value transactions always finalize at $\Delta^{*}$ (or via strong chains), while only lower-value ones qualify for faster stake-based finality. This makes such attacks costly, as the adversary must submit many high-value transactions to be effective.
Bribing Mechanism
Trustless bribing mechanisms enable cooperation between Byzantine and rational nodes. A Byzantine node can bribe a rational validator $p$ to sign two conflicting blocks $b$ and $b^{\prime}$ (e.g., by sending pre-signed transactions). Similarly, censorship of a target client $p^{\prime}$ can be incentivized, effectively rewarding exclusion of $p^{\prime}$’s transactions. A malicious client can deploy a smart contract, funded every second block, that pays out in the next block if $p^{\prime}$’s wallet balance remains unchanged, allowing rational validators to claim the bribe only if $p^{\prime}$’s wallet is not updated.
Responsive Synchronous Model
[…]
When applied to partially synchronous protocols, our transformation assumes a model similar to prior work [12, 8]. In addition to the usual partial synchrony assumptions, where message delays are bounded by $\Delta$ after GST, we introduce a larger synchronous bound $\Delta^{*}$ that applies even before GST. This captures the presence of slower, potentially out-of-band communication channels that remain functional during network disruptions or adversarial conditions, with $\Delta^{*}$ possibly orders of magnitude greater than $\Delta$.
8 Related Work
Blockchains usually operate in financial contexts, and thus, their validators which are responsible for their maintenance and security, are typically financially motivated. Consequently, designing incentive-compatible consensus protocols is crucial, particularly given the permissionless nature of blockchains. While production blockchains such as Bitcoin and Ethereum incorporate incentive mechanisms, they are not incentive-compatible [19, 20, 21, 22, 23].
In the PoW setting, the state-of-the-art incentive-compatible blockchain protocol is FruitChains [24], which remains secure as long as rational miners do not form a coalition that exceeds $1/2$ of the total mining power. Several subsequent works have proposed diverse incentives mechanisms for PoW blockchains to optimize specific metrics, e.g., latency [25].
While PoW blockchains continue to evolve, environmental concerns and scalability limitations have driven the blockchain community towards Proof of Stake (PoS) blockchains. Recent research has particularly focused on accountable PoS systems [26, 27, 28, 29, 30], where misbehaving participants can be identified and penalized.
Closely related works that adopt a mixed Byzantine-rational model similar to ours include BAR [31] and Trap [5]. Trap [5], introduces a baiting game to hold misbehaving nodes accountable, solving one-shot consensus in the partial synchronous model for $n>\max\left(\frac{3}{2}k+3f,2(k+f)\right)$. In contrast, our work focuses on multi-shot consensus to realize an SMR protocol. Our solution achieves better resiliency while tolerating up to $f+k<\frac{2}{3}$, compared to Trap’s limit of $f+k<\frac{1}{2}$. While our protocol relies on synchrony, we prove that achieving SMR in a partially synchronous network is impossible under our model. On the other hand, BAR is designed for general-purpose state machine replication and does not address the economic incentive structures or adversarial dynamics specific to payment systems. As such, it is not directly applicable to our work, which explicitly tackles rational deviations in the context of decentralized payment protocols, providing guarantees to validators and clients.
The most closely related works are BLR24 [4] and ZLB [6], both of which investigate economic safety in payment systems. BLR24 examines economic safety in PoS blockchains with the goal to determine when slashing mechanisms can effectively deter misbehaving validators from launching equivocation attacks while ensuring honest validators retain their stake. They prove that slashing is impossible when more than $2f$ out of $13f+$ 1 validators misbehave, and propose a slashing mechanism when at most $2f$ do. Although our bounds align, we also provide resilience up to $2f$ Byzantine and rational nodes, our focus is fundamentally different. BLR24 [4] does not account for liquid coin transfers during SMR execution or explicitly model profit generation beyond slashing deterrents.
ZLB provides probabilistic economic safety in the presence of up to $\frac{5}{9}$ adversarial nodes, where less than $\frac{2}{3}$ can be Byzantine while the rest follow the a-b-c model [32]. Although ZLB supports the partially synchronous model, it does so probabilistically by leveraging a random beacon to construct random committees. In contrast, our approach provides stronger, deterministic guarantees for transaction finality, ensuring that any transaction accepted by clients is never reverted.
Unlike BLR24 [4] and ZLB [6], we introduce a finer classification of adversaries: rational validators, who optimize profit strictly within the PoS framework (e.g., via double-spending), and Byzantine validators, who may have external incentives. Under this model, we prove that finalized transactions will never be reverted, ensuring that rational validators find equivocation unprofitable and that Byzantine adversaries alone cannot cause equivocation. To achieve this stronger finality guarantee, we design novel confirmation rules for both validators and clients, explicitly considering liquid coin transfers and block signature counts within a bounded time window.
Several recent works explore recovery protocols for when the $\frac{1}{3}$ fault threshold is exceeded, as well as the provision of cryptoeconomic safety guarantees for clients [7, 33, 9]. These findings are complementary to ours. Our solution ensures that rational validators have no incentive to misbehave, thereby preventing consistency violations and eliminating the need for recovery protocols within the scope of our model. Nevertheless, in practice, incorporating recovery mechanisms could be beneficial to handle failures caused by unforeseen validator behavior that falls outside the assumptions of our rational model.
Finally, our impossibility results under partial synchrony align with [34, 4], which show that Byzantine validators can cause safety violations and avoid slashing before cashing out once they exceed the security threshold. We extend these findings by showing that, even when the Byzantine threshold is not exceeded, the attack remains feasible and profitable if just one honest validator is replaced by a rational one. Moreover, our finality impossibility (Theorem 3) demonstrates that such attacks remain feasible and profitable for protocols finalizing blocks before a large synchronous bound elapses.
Appendix A Rational attacks
In this section, we illustrate the challenges of extending $(n,f)$–resilient SMR protocols towards $(n,k,f)$–resiliency for $k\geq 1$.
Safety attacks
Conflicting ledgers and double-spending. Two blocks $b_{1}\neq b_{2}$ are conflicting if they both extend the same transaction ledger $T$. Moreover, we say that two ledgers $T_{1}$, $T_{2}$ are conflicting with each other, if there are two conflicting blocks $b_{1}\in T_{1}$, $b_{2}\in T_{2}$. Last, we say that there is a fork in the system if there are at least two conflicting ledgers $T_{1}$ and $T_{2}$.
The attack. Rational validators can collude with adversarial validators to create two conflicting blocks $b_{1}$ and $b_{2}$ with valid inclusion proofs, where they spend $\alpha$ coins to two distinct subsets of clients $C_{1}$ and $C_{2}$. For clients in $C_{1}$ (or $C_{2}$), the $\alpha$ coins have not been spent yet, and therefore, rational validators can spend them again. In that case, we say that rational validators double-spend $\alpha$ coins. We present this attack in Lemma 8, showing that $(n,f)$-resilient blockchains are vulnerable to double-spending.
Lemma 8 (Double-Spending Attack).
Consider a system of $n=f+k+h$ validators, where $f$ are adversarial, $k$ are rational, and $h$ are correct validators. Let $\Delta^{*}$ denote the maximum message delay (either known or not before execution). For any $q\leq f+k+\lfloor h/2\rfloor$ and $f+k>1$ and any q–commitable protocol $\Pi$ that is $(n,f)$-resilient:
  1. 1.
    $\Pi$ is vulnerable to double-spending attacks,
  2. 2.
    if $\Pi$ is broadcast based, any equivocation is witnessed by all validators within $2\Delta^{*}$.
Proof.
(1): Assume the leader $L_{r}$ proposing the next block at some round $r$ (w.l.o.g.) is malicious. If conflicting blocks already exist, the claim holds. Otherwise, all correct validators initially agree on the same ledger $T$. A malicious leader partitions the correct validators into two subsets, $H_{1}$ and $H_{2}$, where $|H_{1}|=\lfloor h/2\rfloor$ and $|H_{2}|=\lceil h/2\rceil$. Before the bound $\Delta^{*}$ elapses, the leader sends a conflicting block $b_{1}$ to $H_{1}$ and $b_{2}$ to $H_{2}$. Rational validators sign both $b_{1}$ and $b_{2}$ because they can present inclusion proofs for $b_{1}$ and $b_{2}$ to disjoint client sets $C_{1}$ and $C_{2}$ in exchange for off-chain assets. Each conflicting block gathers $q$ signatures since $f+k+\lfloor h/2\rfloor\geq q$ and $f+k+\lceil h/2\rceil\geq q$. Therefore, there is a valid certificate for the conflicting blocks $b_{1}$ and $b_{2}$.
(2): Let us fix any honest validator $p_{1}\in H_{1}$ that witnesses the valid certificate $QC(b_{1})$ for block $b_{1}$ at time $t_{1}$ [^9] We use $p_{i}$’s local time as a reference, as each validator measures elapsed time using its own clock.. Similarly, let $p_{2}\in H_{2}$ be an honest validator that witnesses the valid certificate $QC(b_{2})$ for a conflicting block $b_{2}$ at time $t_{2}$, and without loss of generality, assume $t_{1}\leq t_{2}$.
First, it must hold that $t_{2}-t_{1}\leq\Delta^{*}$. That is because every validator in $H_{2}$ will have receive $QC(b_{1})$ at time $t_{1}+\Delta^{*}$ and would not vote for the conflicting $QC(b_{2})$. Moreover, each validator will have received and witnessed both certificates by time $t_{2}+\Delta^{*}\leq t_{1}+2\Delta^{*}$. Thus, any two conflicting blocks with valid certificates will be observed by all honest validators within $2\Delta^{*}$.
Liveness Attacks
To guarantee liveness, q–commitable protocols typically assume at least $h\geq q$ correct validators. When this is not the case, rational validators must be incentivized to follow the protocol, introducing new challenges in aligning incentives with correct execution.
Censoring transactions. In traditional Proof-of-Stake blockchains, the majority of transaction fees are awarded to the validator that successfully proposes a block, incentivizing the inclusion of as many transactions as possible. However, if a proposer fails to produce a valid block, the pool of unprocessed transactions grows, potentially increasing the rewards available to subsequent proposers. This dynamic creates an incentive for rational validators to launch censorship attacks [35] against competing proposers to maximize their own rewards.
Bribery attacks. Byzantine validators can use their share of the rewards to bribe rational validators and censor all correct validators. As an example, we consider a $12f+$ 1–commitable SMR protocol with $f$ Byzantine, $f$ rational, and $f+1$ correct validators. To successfully launch the attack, Byzantine validators only need to bribe a single rational validator $p$, by distributing all their rewards to $p$. In that scenario, all the rest block proposers gather at most $2f$ signatures for their proposals. Consequently, only the byzantine validators along with $p$ would successfully propose a block. As a result, the adversary would manage to censor any transaction held by correct validators.
Appendix B Omitted proofs
Lemma 9.
Assume that correct validators employ the Strong Chain finalization rule and at least one coalition of validators forks the system. Any rational validator participating in the attack can double spend at most $D$ coins.
Proof.
Similar to Lemma 2, we will upper bound the total amount of coins in conflicting $2\Delta^{*}$-recent blocks finalized by at least one validator.
Case 1: Every correct validator finalizes up to $D$. The proof is similar to Lemma 2.
Case 2: At least one correct validator $p_{i}$ monitors an $i$-strong chain $L_{p_{i}}$ with $i>f/4$. Since $1+i$ correct validators monitor $L_{p_{i}}$, there can be at most $f-i$ additional conflicting ledgers. We assume the worst case, which is that $f-i$ conflicting ledgers exist. We fix any rational validator $p$ engaging in the attack, participating in $f-m-i$ conflicting ledgers $L_{i\in[f-m-i]}$, with $m\in[f-i]$.
For each $i\in[f-m-i]$, at most $m+1$ correct validators and therefore at least $f-m$ rational validators monitor ledger $L_{i}$. We prove that by contradiction. Assume there is a ledger $L_{i}$ with $m+2$ correct validators. At least $i+1$ correct validators monitor $L_{p_{i}}$, at least $m+2$ correct validators monitor $L_{i}$ and at least one correct validator exists in the rest $f-m-i-1$ conflicting ledgers $L_{j\neq i}$. In total, this accounts for $f+2$ correct validators, leading to a contradiction. Furthermore, since there are at most $m+1$ correct validators per ledger, the total amount of coins finalized by correct validators on any ledger is at most $\frac{f}{f-m}D$.
Finally, there are $f-m-i$ conflicting ledgers with at most $f-m$ rational validators spending up to $\frac{f}{f-m}D$ per ledger. So, the amount of double-spent coins is bounded by $max_{\{m\in[f-i]\}}\frac{(f-m-i)f}{(f-m)^{2}}D$ which is bounded by $D$ for $i\in\{\frac{f}{4},...,f\}$. We can show that, by taking the continuous extension of the function $F(m)=\frac{(f-m-i)f}{(f-m)^{2}}D$ for $m\in[0,f-i]$ and some constant $i\in\{\lfloor\frac{f}{4}+1\rfloor,...,f\}$, which takes its maximum value for $m=f-2i$. Since the function takes maximum value at an integer value, the maximum will be the same for the discrete function and thus $\frac{f}{4i}D\leq D$.
Lemma 10.
At any round, there can only exist a unique strongest chain.
Proof.
For the sake of contradiction, assume that at some round $r$, there exist two conflicting chains $T_{1}$, $T_{2}$ that both form a strongest chain. That means that $T_{1}$ is an $i_{1}$-strong chain for some $i_{1}>\frac{f-1}{2}$ and $T_{2}$ is an $i_{2}$-strong chain for some $i_{2}>\frac{f-1}{2}$. Therefore, there exist two distinct subsets $H_{1}$ and $H_{2}$ that consist of $1+i_{1}$ correct validators participating in $T_{1}$ and $1+i_{2}$ correct validators participating in $T_{2}$ respectively. However, $11+i_{1}+1+i_{2}>f+$ 1. Since there are $f+1$ only correct validators in the system, we reached a contradiction. ∎
Lemma 11.
Assume that correct validators employ the Strongest Chain finalization rule and at least one coalition of validators forks the system. Any rational validator participating in the attack can double-spend at most $D$ coins.
Proof.
We will prove that when validators employ the strongest chain finalization rule, rational validators participating in coalitions can double-spend up to $D$ coins. Similar to Lemma 2, we will upper bound the total amount of coins spent in conflicting $2\Delta^{*}$-recent finalized blocks.
Case 1: There is no correct validator $p_{i}$ monitoring an $i_{1}$-strong chain with $i_{1}>(f+1)/2$. If every correct validator finalizes up to $D$, the proof follows from Lemma 2. If at least one correct validator monitors an $i_{1}$-strong chain with $i_{1}>f/4$, the proof follows from Lemma 9.
Case 2: At least one correct validator $p_{i}$ monitors an $i_{1}$-strong chain with $i_{1}>(f+1)/2$. By upper bounding the total coin value in the last $2\Delta^{*}$ blocks of conflicting ledgers finalized by at least one validator, we also bound the double-spent coins per misbehaving validator.
Case 2a: Every correct validator in a conflicting ledger, finalizes up to $D$. A number of $1+i_{1}$ correct validators monitor $L_{p_{i}}$, and thus there are up to $f-i_{1}$ extra conflicting ledgers. Assume that there are $f-i_{1}$ extra conflicting ledgers and let us any rational validator participating in $f-m-i_{1}$ of the conflicting ledgers for $m\in[f-i]$. In each fork, there are at most $m+1$ correct validators and at least $f-m$ rational validators (with the same analysis done in Lemma 2). Therefore, rational misbehaving validators can double-spend up to $max_{\{m\in[f-i_{1}]\}}\frac{(f-m-i_{1})}{(f-m)}D\leq{m\in[f-i_{1}]\}}\frac{(f-m)}{(f-m)}D=D$.
Case 2b: At least one correct validator $p_{j}$ monitors an $i_{2}$-strong chain $L_{p_{j}}$ with $i_{2}>f/4$.. Let $i_{1}^{*}=i_{1}+1$ and $i_{2}^{*}=i_{2}+1$. In that scenario, at least $i_{1}^{*}>\lceil\frac{f+3}{2}\rceil$ correct validators monitor $L_{p_{i}}$ and at least $i_{2}^{*}$ correct validators, where $\lceil\frac{f+4}{4}\rceil<i_{2}^{*}<\lceil\frac{f-1}{2}\rceil$, monitor ledger $L_{p_{j}}$. Therefore, there are at $i_{3}=f+1-i_{1}^{*}-i_{2}^{*}<\frac{f-6}{4}<\frac{f}{4}$ remaining correct validators to participate in the rest conflicting ledgers. Thus, the there are up to $i_{3}$ conflicting ledgers, where correct validators finalize up to $D$ coins (since there is a participation of less than $f/4$ correct validators). Assume the worst case, where there are $i_{3}$ such conflicting ledgers $L_{i\in[i_{3}]}$.
Now, we fix any rational validator $p$ participating in $L_{p_{i}}$ and in $i_{3}-m$ of the conflicting ledgers $L_{i\in[i_{3}-m]}$ for $m\in[i_{3}]$. We will upper bound the coins that $p$ can spend in these conflicting blocks. First, in $L_{p_{j}}$, there are finalized up to $\frac{f}{(f-m)(f-i_{2})}D<A(m)=\frac{f}{(f-m)(f-\frac{f-1}{2})}D$. For the conflicting ledgers $L_{i\in[i_{3}-m]}$, we can show that there are at most $m+1$ correct validators and at least $f-m$ rational validators per fork, with the same analysis done in 2. Thus, each validator can double-spend up to $\frac{f+1-i_{1}^{*}-i_{2}^{*}-m}{f-m}D=(1+\frac{1-i_{1}^{*}-i_{2}^{*}}{f-m})D\leq B(m)=1-\frac{3(f+2)}{4(f-m)})D$. Now we need to upper bound $C(m)=A(m)+B(m)=(1-\frac{3f^{2}+f+6}{4(f-m)(f+1)})D<D$ since $m\in[f+1-i_{1}^{*}-i_{2}^{*}]$ and $f>0$.
Appendix C Impossibility results
Proof of Theorem 3
Proof.
For the sake of contradiction assume that there is a q–commitable SMR protocol $\Pi$ that is $(n,k,f)$-resilient for $f\geq\lceil h/2\rceil$, where $h=n-k-f$ is the number of correct validators. Assume that each validator has deposited $D\in\mathbf{N}$ stake to participate in $\Pi$ and assume any period $\Delta_{W}$ where a validator can withdraw their stake.
We will construct a run where colluding with Byzantine validators to create valid certificates for conflicting blocks is a dominant strategy for rational validators. This will lead to a safety violation.
Double spending attack. Since $f\geq\lceil h/2\rceil$, it holds that $f+k+\lfloor h/2\rfloor\geq q$. In Lemma 8, we showed that a malicious validator can partition correct validators in two distinct subsets $H_{1}$ and $H_{2}$ monitoring two conflicting ledgers $T_{c}\cup T_{1}$ and $T_{c}\cup T_{2}$ respectively, where $T_{c}$ is the common part of the ledgers. Assume the attack takes place during an interval  [^10] We can use a reference point the local clock of any correct validator. $[t,t^{\prime}]$ where $t^{\prime}>t+\Delta_{W}$ and let $\mathcal{F}$ be the set of the $f+k$ misbehaving validators.
We will construct $T_{1}$ and $T_{2}$ as follows. There is a partition of the clients in two sets $C_{1},C_{2}$, and a set of transactions $Tx=Tx_{1}\cup Tx_{2}$ input to the SMR protocol $\Pi$, where a set $\mathcal{R}$ of at least $q-f-min(|H_{1}|,|H_{2}|)$  [^11] If $q-f-\min(|H_{1}|,|H_{2}|)=0$, meaning there are enough Byzantine validators to carry out the attack, the analysis follows trivially. rational validators spend all their liquid coins $\epsilon>0$ in $Tx_{1}$ to a subset of clients in $C_{1}$ and the same $\epsilon>0$ coins in $Tx_{2}$ to a subset of clients $C_{2}$. Moreover, in $Tx_{1}$ and $Tx_{2}$, there are included all necessary transactions so that validators can withdraw their stake.
We will show that such an attack is feasible. That is, if rational nodes in $\mathcal{R}$ perform the attack, it leads to a safety violation. Consider a correct validator $p_{1}\in H_{1}$ and any correct validator $p_{2}\in H_{2}$. Since $\Pi$ solves SMR, it guarantees liveness, and therefore, every transaction in $Tx$ is committed to the ledger of $p_{1}$ and $p_{2}$  [^12] We assume that correct validators gossip valid transactions, so the transaction will eventually be witnessed by every correct validators. . We prove that $p_{1}$ and $p_{2}$ will finalize the blocks in the conflicting ledgers $T_{1}$ and $T_{2}$ respectively, using an indistinguishability argument.
World 1. (GST = 0.) Validators in $H_{1}$ and $F$ are correct. Validators in $H_{2}$ are Byzantine and do not participate in the protocol. The input to $\Pi$ are the transactions in $Tx_{1}$. Validator $p_{1}\in H_{1}$ witnesses blocks in $T_{1}$  [^13] We mean that $p_{1}$ sequentially finalizes blocks in $T_{1}$. with quorum certificates of $n-f$ votes.
World 2. (GST = 0) This is symmetrical to World 1. Now, validators in $H_{2}$ and $F$ are correct. Validators in $H_{1}$ are Byzantine and do not participate in the protocol. The input to $\Pi$ are the transactions in $Tx_{2}$. Validator $p_{2}\in H_{2}$ witnesses blocks in $T_{2}$ along with the quorum certificates of $n-f$ votes.
World 3. ($GST=t^{\prime}$). Validators in $F$ perform the Double Spending Attack, as explained before, until $GST$. First, the adversary partitions the network between validators in $H_{1}$ and $H_{2}$, such that messages sent by validators in $H_{1}$ are received from validators in $H_{2}$ only after $GST$ happens, and vice versa. The input to $p_{1}\in H_{1}$ are the transactions in $Tx_{1}$ and the input to $p_{2}\in H_{2}$ are the transactions in $Tx_{2}$. Then, every validator in $F$ simulates two correct validators. The first one simulates the execution of World 1 to any validator $p_{1}\in H_{1}$ and the second one simulates the execution of World 2 to any validator $p_{2}\in H_{2}$.
In the view of $p_{1}$ World 1 and World 3 are indistinguishable. Thus, $p_{1}$ finalizes blocks in $T_{1}$ after $T_{c}$. In the view of $p_{2}$ World 2 and World 3 are indistinguishable. Therefore, $p_{2}$ finalizes blocks in $T_{2}$ after $T_{c}$ leading to a safety violation.
Then, to argue that the attack is profitable for rational nodes we will show that clients will accept the inclusion proofs for conflicting blocks. Since $\Pi$ satisfies client-liveness, there must be a rule so that clients in $C_{1}$ (or $C_{2}$) can receive inclusion proofs for blocks in $T_{c}$. We will show that any rule $\rho$ that satisfies client-liveness, leads to clients accepting the inclusion proofs for the conflicting blocks in the double spending attack.
In q–commitable protocols, the rule $\rho$ can either depend on some time bound $\Delta$ or on a certificate of votes. However, we show that rule $\rho$ cannot depend on a time bound in the partial synchronous model using an indistinguishability argument. Consider the time bound $\Delta$ after GST. Moreover, consider any rule $\rho_{1}$ according which clients wait for the message delay $k\Delta$ for any $k\in\mathbb{Z}$ to receive a proof of equivocation and, w.l.o.g., let us fix some client $c_{1}\in C_{1}$.
World A. (GST=0). The next blocks with valid certificates extending $T_{c}$ are the blocks in $T_{1}$. Validator, $p_{1}$ commits blocks in $T_{1}$ in its local ledger. Client $c_{1}$ receives a valid certificate for any block $b\in T_{1}$ from $p_{i}$ and waits for $k\Delta$. Since there are not conflicting blocks in the system, client $c_{1}$ does not receive a proof of equivocation.
World B. ($GST>t^{\prime}+k\Delta)$. Validators in $\mathcal{F}$ perform the double spending attack at time $t$. Validator $p_{1}\in H_{1}$ finalizes blocks in $T_{1}$ to extend $T_{c}$ and validator $p_{2}\in H_{2}$ finalizes blocks in $T_{2}$ to extend $T_{c}$. Validator $p_{1}$ sends the inclusion proof for blocks in $T_{1}$ to $c_{1}$. Client $c_{1}$ waits for $k\Delta$. At $t^{\prime}+k\Delta<GST$, $c_{1}$ does not receive any message from validators in $H_{2}$ and does not witness conflicting blocks.
Client $c_{1}$ cannot distinguish between World A and World B and therefore, the rule $\rho_{1}$ is not sufficient. Next, consides certificates of stages of at least $t$ votes. since the $f$ Byzantine validators might never participate, $t\leq n-f$. Now, consider any rules $\rho_{a}$ and $\rho_{b}$, which are receiving stages of at least $t_{1}=n-f$ votes and $t_{2}$ votes for some $t_{2}\in[1,n-f)$ respectively. Any certificate validating $\rho_{1}$ also validates $\rho_{2}$. Therefore, an inclusion proof consisting of stages of $t_{1}=n-f\geq q$ votes is the strongest certificate that clients can receive, and will therefore accept the inclusion proofs for the blocks in the conflicting ledgers $T_{1},T_{2}$.  [^14] A different approach is to notice that $c_{1}$ cannot distinguish between World 1 and World 2 of the Double Spending Attack.
Proof of Theorem 3
Proof.
For the sake of contradiction assume that there is a q–commitable SMR protocol $\Pi$ that is $(n,k,f)$-resilient for $f\geq h$, where $h=n-k-f$ is the number of correct validators. Assume that each validator has deposited $D\in\mathbf{N}$ stake to participate in $\Pi$.
We will construct a run where it is a dominant strategy for rational validators to collude with Byzantine validators and create valid certificates for blocks that are not witnessed, and therefore not finalized, by any correct validator. Clients will accept the certificates, leading to a client-safety violation. Towards this direction, consider a run of $\Pi$ and let us denote by $C_{t}$ the number of clients in the system at time $t$. Assume that all correct validators agree in the transaction ledger $T_{c}$ at time $t-\epsilon$ for some $\epsilon>0$.
Double spending attack. The input of $\Pi$ at time $t$ is a set of transactions $Tx=\cup_{i=1}^{C_{t}}Tx_{i}$ constructed as follows. There is a set $\mathcal{R}$ of at least $r=max(q-f,0)$ rational validators such that every validator in $\mathcal{R}$ transfers $c>0$ coins to the client $c_{i\in\{1,\dots,C_{t}\}}$ in $Tx_{i\in\{1,\dots,C_{t}\}}$. Moreover, we construct the run such that $(C_{t}-1)C>D$. Since the deposit $D$ is decided before the run of the protocol and $C_{t}$ is arbitrary, it is feasible to construct such a run.
Validators in $\mathcal{R}$, along with the $f$ Byzantine validators, create the blocks $b_{i\in\{1,\dots,C_{t}\}}$, where for each $i,b_{i}$ includes the transactions $Tx_{i}$. Since $r+f\geq q$, they can construct $s$ phases (for any $s$ specified by $\Pi$) of $q$ votes for every blocks $b_{i\in\{1,\dots,C_{t}\}}$. Note that none of the correct validators has witnessed the valid certificate for any of those blocks.
We will show that clients $c_{i\in\{1,\dots,C_{t}\}}$ will accept the inclusion proofs for blocks $b_{i\in\{1,\dots,C_{t}\}}$. In that scenario, any rational validator $p\in\mathcal{R}$ double spends $(C_{t}-1)C>D$ coins. Even if validator $p$ is slashed in $\Pi$, the attack is still beneficial. This will lead to violation of client-safety.
Clients must accept the inclusion proofs. Since $\Pi$ ensures client-liveness, there is a rule $\rho$ that allows clients to accept inclusion proofs for every block committed to the ledger $T_{c}$. We will show that for any such rule $\rho$, all clients $c_{i\in\{1,\dots,C_{t}\}}$ will accept inclusion proofs for blocks $b_{i\in\{1,\dots,C_{t}\}}$.
First, consider a rule $\rho_{1}$ where clients wait for a fixed time bound $k\Delta$, where $\Delta$ is the synchrony bound and $k>0$. We fix a client $c_{i\in\{1,\dots,C_{t}\}}$. We prove that this rule is insufficient using an indistinguishability argument.
World 1. The next block extending $T_{c}$ with a valid certificate is block $b_{i}$. At least one correct validator $p_{i}$, commits $b_{i}$ in its local ledger. There are not conflicting blocks in the system. Client $c_{i}$ receives a valid certificate for $b_{i}$ from $p_{i}$ and waits for $k\Delta$. After the period passes $c_{i}$ detects no conflict.
World 2. Rational nodes collude with Byzantine nodes to perform the double spending attack described above. A rational node $p\in\mathcal{R}$ sends the inclusion proof of $b_{i}$ to $c_{i}$. Client $c_{i}$ waits for $k\Delta$. Since no correct validator has witnessed the certificates for any conflicting block $b_{j\in\{1,\dots,C_{t}\},j\neq i}$, client $c_{i}$ detects no conflict.
Client $c_{i}$ cannot distinguish between World 1 and World 2 and therefore, the rule $\rho_{1}$ is not sufficient. Now, consider any rule $\rho_{1}$ which requires stages of at least $t$ votes. Similar to the proof of Theorem 3, an inclusion proof of stages of $t=n-f$ votes is the strongest certificate that clients can receive. Therefore, each client $c_{i\in\{1,\dots,C_{t}\}}$ clients will accept the inclusion proofs for the respective block $b_{i}$.
Proof of Theorem 3
Assume, for contradiction, that there exists a $q$-participation SMR protocol $\Pi$ that is $(n,k,f)$-resilient for $f\geq\lceil h/2\rceil$, where $h=n-k-f$ represents the number of correct validators, and $\Pi$ does not incur a finality delay of at least $\Delta^{*}$. Moreover, suppose each validator has staked $D\in\mathbf{N}$ to take part in $\Pi$.
We will construct an attack where it is dominant strategy for rational validators to collude with Byzantine validators and produce valid certificates for conflicting blocks, leading to safety violation. Similar to the proof of Theorem 3, a malicious validator can partition correct validators in two distinct subsets $H_{1}$ and $H_{2}$ which monitor the conflicting ledgers $T_{c}\cup T_{1}$ and $T_{c}\cup T_{2}$ respectively. Additionally, there is a partition of the clients in two sets $C_{1},C_{2}$, and the input to the blockchain protocol is $Tx=Tx_{1}\cup Tx_{2}$, where a set $\mathcal{R}$ of at least $p-f-min(|H_{1}|,|H_{2}|)$ rational validators spend $D+\epsilon_{1}$ coins in $Tx_{1}$ to a subset of clients in $C_{1}$ and $D+\epsilon_{2}$ coins in $Tx_{2}$ to a subset of clients $C_{2}$ where $\epsilon_{1},\epsilon_{2}>0$. Then, every transaction in $Tx_{1}$ must be committed to $T_{1}$, and every transaction in $Tx_{2}$ must be committed to $T_{2}$ within $\Delta^{*}$ by assumption.
We will now show that such a run is feasible. Specifically, we will construct a run where, for any $D$, every rational validator spend $D+\epsilon_{1}$ coins in $T_{1}$ (same for $T_{2}$). To this end, cosnider the following notation. $G^{p}_{0}$ denotes the coins of validator $p$ in $T_{c}$. Then, we denote by $S^{p}_{i}$ the total coins paid to clients in $C_{1}$ by validator $p$ and by $R^{p}_{i}$ the participation rewards $p$ has received until the $i$-th block of $T_{1}$ (excluding $T_{c}$). We need to show that for every rational validator $p$ there is a $k$ s.t. $S^{p}_{k}>D$, or $G^{p}_{0}+R^{p}_{k}>D$.
We will show that the participation rewards $p$ receives cannot be bounded. Toward this direction, consider the first $i\in\mathcal{N}$ such that $R^{p}_{i}=\mathcal{C}^{p}_{0}$ as follows. Every validator in $\mathcal{R}$ spends all their coins, and therefore all coins belong to the hands of all the other users. When enough blocks has been processed in the system, any rational validator $p$ actively participating (by proposing or voting blocks depending on the reward scheme) will receive at least $\mathcal{C}^{p}_{0}$ in participation rewards. Specifically, w.l.o.g., if the participation rewards per block are $r$,$n$ blocks have been processed in the system and $p$ has participated in a percentage of $\alpha$ of them, then there is a run where $\alpha\cdot n\cdot r\geq\mathcal{C}^{p}_{0}$, which holds since it can be $n\to\infty$. Then, rational validators will spend all of their money to pay clients in $C_{1}$ and the above scenario is repeated for $j$ times such that there is $i\in\mathcal{N}$, $R^{p}_{i}=j\mathcal{C}^{p}_{0}$ and $(j+1)\mathcal{C}^{p}_{0}>D$. Since it can be that $j\to\infty$, this is a feasible scenario.
Theorems 3 and 3 share the same Byzantine resilience and assume similar network conditions during a partition lasting up to $\Delta^{*}$. Thus, the proof of Theorem 3 extends to show that: (1) safety is violated, with correct validators $p_{1}\in H_{1}$ and $p_{2}\in H_{2}$ finalizing conflicting blocks in ledgers $T_{1}$ and $T_{2}$, and (2) clients accept inclusion proofs backed by valid certificates. ∎