### v3

# CoBRA: A Universal Strategyproof Confirmation Protocol for Quorum-based Proof-of-Stake Blockchains

Abstract.
The integrity of many Proof-of-Stake (PoS) payment systems relies on quorum-based State Machine Replication (SMR) protocols mitigating diverse adversarial attacks. While traditional analyses assume a purely Byzantine threat model, practical security depends on resilience against both arbitrary faults and strategic, profit-driven actors. In this work, we study this challenge by analyzing SMR protocols under a hybrid threat model comprising honest, Byzantine, and rational validators. We first establish the fundamental limitations of traditional consensus mechanisms, proving two impossibility results: (1) in partially synchronous networks, no quorum-based protocol can achieve SMR when rational and Byzantine validators collectively exceed $1/3$ of the participants; and (2) even under synchronous network assumptions, SMR remains unattainable if this coalition comprises more than $2/3$ of the validator set.
Motivated by these boundaries, we develop a protocol that achieves SMR in the presence of up to $1/3$ Byzantine and $1/3$ rational validators. Our approach relies on two complementary mechanisms to circumvent our impossibility results. First, we introduce a protocol constraint that enforces an upper bound on the transaction volume finalized within any time window $\Delta$, and we prove that such a bound is a necessary condition for security. Second, we define the strongest chain rule, a finality gadget that enables instant transaction confirmation when a supermajority of participants provably supports the execution. We validate the feasibility of this design through an empirical analysis of the Ethereum and Cosmos networks, demonstrating that validator participation exceeds the required $5/6$ threshold in over $99\%$ of observed blocks.
Finally, we explore adversarial settings with $\geq 1/3$ Byzantine validators. There we provide a recovery mechanism capable of restoring safety and liveness following consistency violations, even when Byzantine validators control up to $5/9$ of the stake and rational actors control $1/9$. This mechanism ensures security by guaranteeing that any clients incurring provable economic losses due to safety faults are fully reimbursed.
Consensus, Blockchain, Proof-of-Stake, Accountability, Recovery
conference: ; $c$s: Theory of computation Algorithmic game theory and mechanism designccs: Computer systems organization Dependable and fault-tolerant systems and networksccs: Security and privacy Distributed systems security
1. Introduction
State Machine Replication (SMR) protocols are the backbone of secure distributed ledgers, ensuring agreement on an ordered sequence of transactions among distributed participants. Traditional SMR protocols model participants as either honest, strictly following the protocol, or Byzantine, deviating arbitrarily. Quorum-based Byzantine Fault Tolerant (BFT) protocols, such as PBFT (Castro and Liskov, 2002), HotStuff (Yin et al., 2018), and Tendermint (Buchman et al., 2018), operate under partial synchrony and guarantee safety and liveness provided that less than $1/3$ of the participants are Byzantine. Once this threshold is exceeded, safety violations are unavoidable.
While the Byzantine model is deliberately pessimistic, the complementary assumption that a $2/3$ supermajority of participants behaves honestly at all times is particularly strong. This is evident in modern Proof-of-Stake (PoS) blockchains, where participation is economically motivated. Validators stake assets to earn rewards from block production, transaction fees, and external financial positions, collectively securing hundreds of billions of dollars in value. In this setting, participants are not altruistic entities but profit-maximizing agents. Deviations from the protocol may therefore be strategic rather than arbitrary, arising whenever doing so increases expected utility. As a result, violations of the honest supermajority assumption may stem not only from malicious behavior, but also from economically rational behavior.
This observation raises a fundamental question for the design of secure distributed ledgers: What assurances can we provide when fewer than $2/3$ of the participants behave honestly? In particular, can we design protocols that both (i) incentivize rational, profit-maximizing validators to follow the protocol, and (ii) recover from safety violations without imposing irreversible economic losses on clients, even when more than $1/3$ of the participants behave Byzantine?
1.1. Related Work
A substantial body of work has studied consensus under violations of the classic $2/3$ honest supermajority assumption. Broadly, this literature falls into two main categories.
Rational and hybrid fault models.
A first line of work studies consensus under hybrid fault models distinguishing honest, rational, and Byzantine participants. BAR (Aiyer et al., 2005) formalizes this setting, and subsequent works extend it to incentive-compatible and equilibrium-based consensus, ensuring honest behavior is a Nash equilibrium under appropriate rewards and penalties (Amoussou-Guenou et al., 2019; McMenamin et al., 2021; Lev-Ari et al., 2019). These approaches typically assume rational participants act independently, ignoring collusion. This assumption is problematic in open, financially motivated environments such as Proof-of-Stake blockchains, validators often coordinate in pools, making deviations like double-spending profitable when executed jointly.
To capture collusions, Abraham et al. (Abraham et al., 2006, 2008) formalize solution concepts for strategic interaction between rational and Byzantine participants in secret sharing and multi-party computation. Building on this, TRAP (Ranchal-Pedrosa and Gramoli, 2022) introduces a baiting game to hold misbehaving nodes accountable, solving one-shot consensus in a partially synchronous model for $n>\max\left(\frac{3}{2}k+3f,2(k+f)\right)$ where $k$ in the number of rational and $f$ the number of Byzantine participants. While one-shot consensus allows bounding the maximum double-spend, this does not extend to multi-shot SMR, where repeated strategic behavior can accumulate unbounded economic damage.
Accountability and recovery.
A second line of work studies consensus protocols that tolerate violations of the classic $n/3$ Byzantine threshold by weakening safety guarantees and providing accountability (Sheng et al., 2021; Civit et al., 2021, 2023) and recovery mechanisms (Lewis-Pye and Roughgarden, 2025; Ranchal-Pedrosa and Gramoli, 2024; Gong et al., 2025).
Accountable protocols focus on detecting equivocation and identifying misbehaving validators, but do not address recovery. As such, accountability alone does not restore safety or protect clients from economic losses incurred during safety violations.
Recoverable consensus protocols address this limitation by incorporating a mechanism that enables the system to internally restore a consistent state after safety violations, even when more than $1/3$ of the participants behave maliciously (Gong et al., 2025; Lewis-Pye and Roughgarden, 2025; Ranchal-Pedrosa and Gramoli, 2024; Sridhar et al., 2023b). However, these protocols either assume communication between clients to detect conflicting certificates (Sridhar et al., 2023b) or provide no economic guarantees to clients (Gong et al., 2025; Lewis-Pye and Roughgarden, 2025; Ranchal-Pedrosa and Gramoli, 2024): transactions accepted during safety violations may be reverted without compensation.
As a result, existing work does not resolve our fundamental question. There is no SMR protocol that (i) incentivizes profit-maximizing participants to follow the protocol and (ii) provides economic security to clients when the classic $1/3$ Byzantine threshold is exceeded.
1.2. Our Contributions
This paper presents CoBRA, which closes this gap. It is the first generic transformation for quorum-based SMR that provides optimal rational resilience when the Byzantine parties are less than 1/3 and economic protection for clients when they are more than 1/3 and less than 5/9. Table 1 compares CoBRA to prior work, highlighting the trade-offs in fault tolerance, recovery guarantees, and client-level economic protection.
Impossibility results.
We establish fundamental limits for quorum-based SMR under a hybrid fault model with honest, Byzantine, and rational validators, and non-interactive clients:
  • Under partial synchrony, we show that quorum-based SMR is impossible once Byzantine and rational validators together control at least $1/3$ of the validator set. This bound is tight and generalizes the classic Byzantine threshold to economically motivated deviations.
  • Under synchrony, we further prove that SMR remains impossible when this coalition exceeds $2/3$.
Protocol transformation.
We introduce CoBRA, a generic transformation applicable to existing quorum-based PoS protocols, yielding three operational regimes:
  • Standard regime: In partial synchrony, CoBRA preserves the safety and liveness guarantees of the underlying SMR protocol against up to $1/3$ Byzantine validators.
  • Rational-resilient regime: Assuming a (possibly large) known synchronous bound $\Delta^{*}$ on message delays, CoBRA ensures safety and liveness even when, in addition to $1/3$ Byzantine validators, up to another $1/3$ of the participants are rational and may engage in strategic collusion. This is achieved by modifying only the finalization rule of the underlying protocol.
  • Recoverable economic-security regime: Assuming a synchronous bound $\Delta^{*}$ when Byzantine behavior exceeds the standard $1/3$ threshold, CoBRA enables recovery from consistency violations and guarantees full client reimbursement. CoBRA provides these guarantees without requiring minting new coins or client participation in the recovery procedure.
2. Preliminaries
2.1. Model and Assumptions
In this section, we outline the model and assumptions used throughout the paper.
Proof-of-stake blockchain
A PoS blockchain implements a distributed ledger protocol (Garay et al., 2024) (i.e., a state machine replication protocol), where participants are selected to create and validate new blocks based on the amount of coins (cryptocurrency) they hold and lock up as stake. The blockchain serves as a payment system in which participants hold and transfer fungible assets, referred to as coins. We also discuss how to extend this model to general-purpose smart contract platforms in Section 8. Each party is associated with a private-public key pair which forms its client identity in the system. The parties that actively participate in maintaining the state of the system (i.e., run the SMR) are known as validators. To do so, they lock coins (i.e., stake) and process transactions, providing clients an ordered sequence of confirmed transactions, referred to as the transaction ledger.
We assume the existence of a genesis block $G$, which contains the initial stake distribution and the identities of the validators participating in the SMR protocol. For the remainder of this paper, we assume a static validator setting, where the stake distribution is known in advance to all participants. In Section 8, we discuss how we can extend our system to a dynamic validator setting.
Validators
For simplicity, we assume a uniform model where all validators lock the same stake equal to $D$ coins and have thus the same “voting power”. If an entity locks some multiple $\alpha\cdot D$, they may control up to $\alpha$ individual validators. Thus, the SMR protocol involves a fixed set of validators, $N$, which are known to all participants. Let $n=|N|$ denote the total number of validators. We assume a public key infrastructure (PKI), where each validator is equipped with a public-private key pair for signing messages, and their public keys are available to all parties.
Clients
Unlike validators, the number of active clients is not known to all parties. Clients have public keys, which they can use to issue authenticated transactions. We assume clients are silent (Sridhar et al., 2024), meaning they neither actively interact with nor observe the execution of the SMR protocol among validators. Instead, clients query a subset of validators to obtain confirmation of any transaction (i.e., the transaction ledger) from their responses.
Cryptography
We assume that all participants are computationally bounded, i.e., they are modeled as probabilistic polynomial-time (PPT) interactive Turing machines. We further assume the existence of cryptographically secure communication channels, hash functions, and digital signatures.
Time
Time progresses in discrete logical units indexed as $u=0,1,2,\dots$. Each participant has their own local clock. We assume that the maximum clock offset between any two parties is bounded, as any such offset can be absorbed into the network delay bound.
Network
We assume that all participants are connected through perfect point-to-point communication channels. We consider two communication models: partial synchrony and synchrony. In the former, we show impossibility results; in the latter, we present both an impossibility and our main results.
(a) Partial synchronous model: There is an unknown time called Global Stabilization Time ($GST$), such that after $GST$ messages arrive within a known time bound $\Delta$. Therefore, every message sent at time $t$ eventually arrives at time $t^{\prime}\leq max\{GST,t\}+\Delta$.
(b) Synchronous model: Any message sent is guaranteed to arrive within a known time bound $\Delta^{*}$. Thus, a message sent at time $t$ is guaranteed to arrive by time $t+\Delta^{*}$.
Adversarial model
We consider a mixed model where validators are honest, Byzantine, or rational. Honest (or correct) validators always follow the protocol. Byzantine validators may deviate arbitrarily from the protocol. We assume a Byzantine adversary does not have any liquid coins, i.e., all their coins are staked to be able to maximize their influence. Rational validators, on the other hand, will deviate from the protocol specification if and only if doing so maximizes their utility. We define a party’s utility as its monetary profit, calculated as the rewards from participation minus potential penalties (e.g., slashing).
We denote the number of Byzantine and rational validators by $f^{*}$ and $k$, respectively. In Section 3, we derive the feasible bounds for $f^{*}$ and $k$. In our solution, we consider $n=3f+1$, where $f$ represents the maximum number of Byzantine validators tolerated by standard BFT protocols, e.g., (Abraham et al., 2018; Buchman et al., 2018; Castro and Liskov, 2002). In Section 4, we specify that $f^{*}\leq f$ and $k\leq 2f-f^{*}$. Then, in Section 5, we extend our solution to Byzantine failures beyond the classic threshold, in particular, $f+1\leq f^{*}<5n/9$ and $k<2n/3-f^{*}$.
We make the following assumptions on rational validators: (a) We assume that participation rewards exceed both the opportunity cost and the cost of operating a validator. Thus, these costs are treated as negligible, simplifying the analysis by offsetting the constant operational expenses (e.g., bandwidth, CPU, storage) or voting for a block proposed by another validator (e.g., adding a signature) and the constant opportunity cost with a minimum constant reward. In practice, rational validators account for these costs beforehand and conclude that participation is beneficial. (b) A rational validator considers a strategy only if the resulting utility is guaranteed within the transaction ledger, as we do not account for external trustless mechanisms. Promises or unverifiable commitments from the adversary regarding future actions are not considered credible and, thus, will not influence the validator’s strategy. Given the appropriate incentives, however, rational validators may collude with other rational and Byzantine validators. (c) A rational validator deviates from the protocol only if such deviation results in strictly greater utility, as devising a strategy to deviate may incur additional costs to running the default client. For example, a rational validator will censor or double spend only if they can secure a bribe larger than their potential loss due to slashing.
2.2. Definitions and Properties
At a high level, in an SMR protocol, validators must perform two primary tasks: order a set of transactions which they execute to extract the state, and reply to clients. Traditional SMR protocols such as PBFT (Castro and Liskov, 2002) and HotStuff (Yin et al., 2018) assume an honest supermajority of validators – specifically $2f+1$ out of the total $3f+1$ validators must be honest. Under this assumption, clients can extract the confirmation of transactions simply by waiting for responses from $f+1$ validators. In contrast, our setting introduces rational validators and assumes only an honest minority. For this reason, we separate these two responsibilities: (i) ordering transactions, and (ii) replying to clients.
SMR problem
Given as input transactions submitted by clients, the validators execute an SMR protocol $\Pi$ and output an ordered sequence of confirmed transactions, known as the transaction ledger. We say that $\Pi$ solves the SMR problem when: (i) correct validators agree in a total order (Défago et al., 2004), (ii) the transaction ledger is certifiable to clients.
Definition 0.
We say that correct validators agree in a total order, when the transaction ledger output by $\Pi$ satisfies:
  • Safety. The transaction ledgers of any two correct validators are prefixes of one another,
  • Liveness. Any valid transaction witnessed by all correct validators, will eventually be included in the transaction ledger of every correct validator.
Definition 0.
We say that $\Pi$ satisfies certifiability if the following holds,
  • Client-safety: Clients accept the confirmation of a transaction only if it is included to the transaction ledger of at least one correct validator,
  • Client-liveness: Clients accept confirmation for any transaction included to the transaction ledger of all correct validators.
We emphasize that total order and certifiability are distinct guarantees. Some protocols achieve total order but not certifiability (e.g., Dolev–Strong (Dolev and Strong, 1983) for $f>n/2$ Byzantine faults). To satisfy both properties, validators and clients may adopt different confirmation rules, in line with the flexible BFT paradigm (Malkhi et al., 2019).
Resiliency of SMR protocols
Traditional SMR protocols (Abraham et al., 2018; Castro and Liskov, 2002; Gelashvili et al., 2022) are secure when considering $f$ Byzantine validators and nfn-f correct validators. We say that a protocol satisfying this condition is $(n,f)$-resilient.
Definition 0.
A protocol $\Pi$ executed by $n$ validators is $(n,f)$-resilient if it remains secure in the presence of up to $f$ Byzantine validators, assuming that the remaining nfn-f validators follow the protocol correctly.
However, in a traditional $(n,f)$-resilient SMR protocol, one rational validator is enough to violate safety and liveness as Byzantine validators can bribe the rational validator to collude with them. We outline such attacks in Appendix A. To capture both byzantine and rational behavior in our protocol, we introduce the notion of $(n,k,f)$-resiliency.
Definition 0.
A protocol $\Pi$ executed by $n$ validators is $(n,k,f)$-resilient if it remains secure in the presence of $f$ Byzantine and $k$ rational validators, assuming that the remaining nfkn-f-k validators follow the protocol correctly.
$q$-commitable SMR protocols
We focus on a family of SMR protocols that progress by collecting certificates of $q$ votes across multiple phases, similar to several BFT SMR protocols (e.g. (Castro and Liskov, 2002; Abraham et al., 2018)). We define a quorum certificate (QC) as a certificate consisting of $q$ votes. A protocol is q-commitable if a QC on a specific message $m$ (e.g., $m=\texttt{commit}$ in PBFT (Castro and Liskov, 2002)) is both a necessary and sufficient condition for validators to commit a transaction (or block) $b$ to their local ledger.
Definition 0.
A protocol $\Pi$ is a $q$-commitable SMR protocol when every correct validator includes a block $b$ in its local transaction ledger if and only if it has witnessed a QC on a specific message $m$, denoted $QC_{m}(b)$.
For simplicity, we may often omit the specific message $m$ and refer to the certificate that finalizes $b$ as $QC(b)$. We stress that $QC(b)$ may not be sufficient for certifiability from the clients’ perspective, as they do not actively monitor the SMR execution (i.e., they are silent). However, for clients that monitor the SMR execution (e.g., the validators), $QC(b)$ is sufficient for certifiability.
We study SMR protocols in the context of proof-of-stake (PoS) blockchains, where validators batch transactions into blocks, as in HotStuff (Yin et al., 2018). Our objective is to determine when an $(n,f)$-resilient SMR protocol can be transformed into an $(n,k,f)$-resilient PoS blockchain. Specifically, we present a transformation that converts any $(n,f)$-resilient $q$-certifiable PoS blockchain (Castro and Liskov, 2002; Abraham et al., 2018; Gelashvili et al., 2022) into an $(n,k,f)$-resilient PoS blockchain. At the core of this transformation, we introduce a new certification mechanism for finalizing transactions, ensuring security under our hybrid model.
Finality latency
Finality latency refers to the time that elapses between a transaction being submitted by a client and its irreversible commitment to the transaction ledger.
Recovery and economic restitution
A protocol that is $(n,k,f)$-resilient constitutes a complete solution to SMR: when at most $f$ validators are Byzantine and at most $k$ are rational, the protocol preserves both safety and liveness. We now consider executions in which the actual number of Byzantine validators $f^{*}$ exceeds the resilience threshold $f$ of an $(n,f)$-resilient system. In such executions, the protocol may suffer both safety and liveness violations.
Liveness attacks cannot, in general, be deterred, since Byzantine validators may simply refuse to participate. For this reason, the literature on recoverable consensus typically assumes an adversary that is willing to participate in the protocol (e.g., to earn participation rewards) but may opportunistically violate safety; this type of faults are called a-b-c Byzantine (Malkhi et al., 2019; Ranchal-Pedrosa and Gramoli, 2024; Gong et al., 2025).
In the same spirit, we focus only on safety attacks with economic impact, namely conflicts among blocks that are accepted by clients [^1] Recall that the certificate which clients accept might differ with the certificate validators use to finalize.. Our goal is to ensure that, even if a client-side safety violation occurs, the system can recover and resume execution from a consistent state in which all honest clients harmed by the violation are economically reimbursed, without minting new coins [^2] Minting new coins inflates the currency and essentially penalizes honest validators..
Definition 0.
An SMR protocol regains economic restitution after safety violations with recovery parameter $\Delta_{R}$ if, after any safety violation, (i) all honest validators eventually recover to a common state $S$ for which there exists a verifiable state certificate that every honest validator witnesses at most $\Delta_{R}$ time after the equivocation, (ii) in state $S$, each client receives the full monetary value of every transaction it has accepted; (iii) validators responsible for the equivocation are punished, (iv) new coins are not minted.
2.3. SMR Game and Coalition-Compliance
To connect game-theoretic rational behavior to protocol-level resilience guarantees, we introduce the solution concept of $(n,k,f)$-coalition-compliance [^3] Coalition-compliance extends the notion of compliance originally introduced in (Karakostas et al., 2022).. Informally, a protocol is $(n,k,f)$-coalition-compliant if no rational validator prefers to follow a strategy that violates the SMR protocol, regardless of the actions of Byzantine validators. This property holds even if all rational and Byzantine validators collude.
By transforming an $(n,f)$-resilient protocol into an $(n,k,f)$-coalition-compliant protocol, we ensure that rational validators have no incentive to violate the SMR security properties. As a result, the transformed protocol effectively achieves $(n,k,f)$-resilience, even when rational and Byzantine parties collude.
Definition 0 (SMR Game).
The SMR game is a tuple:
$G=(n,(S_{i})_{i\in[n]},(u_{i})_{i\in[n]},H)$
where:
  • $n\in\mathbb{N}$ is the number of players. Each player is either correct, Byzantine, or rational.
  • $H$ is the global history of the protocol execution. Each player $i\in[n]$ observes a local view $H_{i}\subseteq H$.
  • $S_{i}$ is the set of all possible strategies available to player $i$, where a strategy is a function mapping local histories $H_{i}$ to actions.
  • A global strategy profile is $s=(s_{1},s_{2},\dots,s_{n})\in S=\prod_{i\in[n]}S_{i}$.
  • The utility function $u_{i}:S\to\mathbb{R}$ gives the total payoff to player $i$ over the entire execution.
We denote by $s_{-i}$ the strategy profile excluding player $i$, and by $S_{-i}$ the set of all such profiles.
We partition the space of strategy profiles into two sets: $S_{\text{bad}}$, consisting of strategy profiles resulting in safety violations (i.e., conflicting blocks are finalized with valid certificates) or liveness violations (i.e., permanent censorship of a transaction), and $S_{\text{good}}$ which preserve both properties.
For each validator $p$, we define $S_{\mathsf{bad}}^{p}$ as the set of individual strategies $s_{p}$ for which there exists a strategy profile $s=(s_{p},s_{-p})\in S_{\mathsf{bad}}$ where $p$ actively contributes to the violation. Conversely, for each validator $p$, we denote by $S_{\mathsf{good}}^{p}$ the set of individual strategies $s_{p}$ such that there exists a strategy profile $s=(s_{p},s_{-p})\in S_{\mathsf{good}}$.
We say that a protocol is coalition-compliant if no rational validator can profit from contributing to safety or liveness violations, even under full collusion with Byzantine validators. In other words, rational validators always have a protocol-compliant strategy that is equally good or better than any deviation causing violations.
Definition 0.
A strategy $s_{i}\in S_{i}$ weakly dominates a strategy $s_{i}^{\prime}\in S_{i}$ if, for every possible strategy profile $s_{-i}\in S_{-i}$, it holds that $u_{i}(s_{i},s_{-i})\geq u_{i}(s^{\prime}_{i},s_{-i}),$ and there exists at least one $s_{-i}\in S_{-i}$ such that: $u_{i}(s_{i},s_{-i})>u_{i}(s^{\prime}_{i},s_{-i}).$
Definition 0.
We say that an SMR game is $(n,k,f)$-coalition-compliance if for any rational validator $p$, and every strategy $s_{p}\in S_{\mathsf{bad}}^{p}$, there exists a strategy $s_{p}^{\prime}\in S_{\mathsf{good}}^{p}$ that weakly dominates $s_{p}$.
Coalition-compliance is a stronger notion than equilibrium, as it accounts for arbitrary collusion among subsets of rational and Byzantine validators. However, coalition-compliance relaxes traditional incentive compatibility, which requires that following the protocol is always the utility-maximizing strategy. Such a requirement is too strong in adversarial settings, where Byzantine validators may accumulate participation rewards and use them to bribe rational ones. Instead, rational coalition-compliance focuses only on the minimal conditions to guarantee safety and liveness.
3. Impossibility results
In the following, we identify settings in which a $q$-committable (n,f)(n,f)–resilient PoS blockchain cannot be transformed into an $(n,k,f^{*})$–resilient PoS blockchain for $f^{*}\leq f$. Additionally, we determine when $q$-committable SMR protocols fail to guarantee finality before the synchrony bound $\Delta^{*}$ elapses.
We provide a summary of our results and defer the proof sketches along with the complete proofs to Appendix C. We also illustrate the main results in Figure 1. We emphasize that our impossibility results apply to the setting of silent clients, where clients decide whether to accept a block based solely on an accompanying certificate, without engaging in additional communication rounds with other participants. Finally, note that (n,k,f)(n,k,f)–coalition compliance is a deterministic guarantee which must hold in every execution of the protocol; thus, a single attack suffices to prove impossibility. We discuss probabilistic guarantees in Section 8.
Partial synchrony
Impossibilities. (A): for $f^{*}>\lceil h/2\rceil$ (Theorem 3), (B): for $f^{*}+k>f$ (Theorem 3).
(C) Any $(n,f)$-resilient SMR protocol is also $(n,f^{*},k)$-resilient for any $f^{*}+k\leq f$.
(D) Enlarging the quorum to $q=n-f^{*}$ allows resilience for all $k\leq n-1$.
Synchrony
(A) Impossibility when $f^{*}+k\geq 2f+1$ (Theorem 3.1).
(B) Possibility of SMR only by limiting transaction volume (CoBRA, Theorem 3).
Possibilities. (C)–(D) As in the partial synchrony setting.
Figure 1. Design space of $2f+1$-committable SMR protocols.
Partial synchrony
In Theorem 3, we show that designing an $(n,k,f)$-resilient $q$-commitable SMR protocol is impossible for $f\geq\lceil\frac{h}{2}\rceil$, where $h$ is the number of correct validators. Notice that this result depends only on the proportion of Byzantine and correct validators. This is because to guarantee liveness, the size of the quorum $q$ is at most nfn-f. For $f\geq\lceil\frac{h}{2}\rceil$, we cannot guarantee that two quorums intersect in at least one correct validator. Rational validators, in turn, will collude with Byzantine validators to form quorums of conflicting blocks to maximize their monetary gains.
To establish this result, we extend the proof of Budish et al. (Budish et al., 2024a, Theorem 4.2) within our model. The core issue is that rational validators can collude with Byzantine to generate valid quorum certificates while avoiding penalties. This is possible because, for any predefined withdrawal period that allows validators to reclaim their stake, correct validators cannot reliably distinguish between an honest execution with network-induced delays and malicious behavior. Hence, rational validators may engage in equivocation without facing slashing, undermining the protocol’s security.
{restatable}
theoremPartialSynchrony Assuming silent clients, a static setting, and a partially synchronous network, no $q$-commitable SMR protocol can be (n,k,f)(n,k,f)–resilient for any $q$ when $f\geq\lceil\frac{h}{2}\rceil$, where h=nkfh=n-k-f.
It follows directly from Theorem 3 that no $q$-commitable protocol can be (n,f)(n,f)–resilient when $f=\lceil h/2\rceil$, which agrees with the well-known result that resilience is impossible once Byzantine validators exceed one third of the network ($f\geq n/3$). Furthermore, the additional insight provided by our result is that even if a $q$-commitable protocol is resilient for $f=\lfloor h/2\rfloor$, replacing just one correct validator with a rational one breaks resilience if the total number of Byzantine and rational validators reaches at least one third the network ($f+k\geq n/3$), as shown in Corollary 3.1. This result highlights the fragility of quorum-based protocols under rational behavior, as even a small shift from correct to rational participants significantly impacts security.
Corollary 3.1.
Consider any $q$-commitable protocol $\Pi$ that is (n,f)(n,f)–resilient for $f=\lfloor\frac{h}{2}\rfloor$ where h=nfh=n-f is the number of correct validators. For any $k\geq 1$, $\Pi$ is not (n,k,f)(n,k,f)–resilient in our model in the partial synchronous setting.
Corollary 3.1 implies that no quorum-based SMR protocol $\Pi$ can achieve (3f+1,k,f)(3f+1,k,f)–resilience for $k\geq 1$ in the partial synchronous model. That means that quorum-based SMR protocols are not resilient for $f+k\geq n/3$ under the model described in Theorem 3.
Fewer Byzantine.
Theorem 3 characterizes the maximum Byzantine resilience of $(n,k,f)$-resilient SMR protocols, independently of the quorum size $q$. A natural question is what happens below this maximum Byzantine threshold. Can an SMR protocol remain secure when $f^{*}<\lfloor h/2\rfloor$ and the combined number of Byzantine and rational validators exceeds one third of the network, i.e., $f^{*}+k>n/3$?
When $f=\lfloor h/2\rfloor$, even a single rational validator can collude with Byzantine validators to equivocate without being detected by other rational validators. However, for $f^{*}<\lfloor h/2\rfloor$, more than one rational validators must collude to equivocate. This opens a potential design space in which rational validators can be incentivized to expose misbehavior. One such approach is a baiting strategy, leveraged by TRAP (Ranchal-Pedrosa and Gramoli, 2022) to achieve consensus on a single value, where rational validators are rewarded for reporting equivocation. Theorem 3 characterizes the conditions under which, even in the presence of baiting strategies, solving SMR remains impossible.
{restatable}
theorempartialSync2 Let us assume participation rewards are distributed to validators, silent clients, a static PoS setting, and a partially synchronous network. Without any bounds i) on the liquid funds held by rational validators or malicious clients, ii) or on the participation rewards, the following holds: no $q$-commitable SMR protocol can be (n,k,f)(n,k,f)–resilient for any $q$ when $q\leq f+k+\lfloor\frac{h}{2}\rfloor$, where h=nkfh=n-k-f.
Theorem 3 implies that no quorum-based SMR protocol can achieve $(3f+1,k,f^{*})$–resiliency with $k=f-f^{*}+1$ in partial synchrony without enlarging the quorum size, as formalized below.
Corollary 3.2.
Consider any $q$-commitable protocol $\Pi$ that is (n,f)(n,f)–resilient for $f=\lfloor\tfrac{h}{2}\rfloor$, where h=nfh=n-f is the number of correct validators. For any $f^{*}<f$ and $k=f-f^{*}+1$, the protocol $\Pi$ is not $(n,k,f^{*})$–resilient in our model under partial synchrony, without enlarging the quorum size.
Thus, 2f+12f+1–commitable quorum-based SMR protocols with a quorum of size $q=2f+1$ are not resilient for any $f^{*}+k\geq n/3$ under the assumptions of Theorem 3.
Closing the gaps
Corollary 3.2 establishes that, in partial synchrony, reducing the actual number of Byzantine validators does not improve resilience unless we increase the quorum size. Following the flexible BFT approach (Malkhi et al., 2019), in this setting, we can enlarge the quorum size used by clients from the classical q=nfq=n-f, to a higher quorum $q^{\prime}=n-f^{*}$, where $f^{*}$ is the actual number of Byzantine validators present. This choice ensures that $q^{\prime}>f^{*}+k+\lfloor h/2\rfloor$, so that any two quorums intersect in at least one honest validator and equivocation becomes impossible. Liveness is guaranteed by the incentives of rational validators, who are motivated to follow the protocol in order to obtain participation rewards.
Synchrony
To circumvent the impossibility result of Theorem 3, we assume synchrony with a known time delay $\Delta^{*}$. However, we now show that even in synchronous settings, designing resilient SMR protocols remains impossible for $f\geq h$. Here, a client cannot distinguish whether correct validators participated in a quorum certificate. Consequently, Byzantine and rational validators can generate certificates that the clients will accept (to ensure client-liveness) without any participation from correct validators.
{restatable}
theoremSynchrony Assuming silent clients, a static setting, and a synchronous network, no $q$-commitable SMR protocol can be $(n,k,f)$-resilient, when $f\geq h$, where h=nkfh=n-k-f.
By setting $k=0$ in Theorem 3, we conclude that no $q$-commitable protocol can be (n,f)(n,f)–resilient when $f=h$. Furthermore, Theorem 3 establishes that even if a $q$-commitable protocol remains (n,f)(n,f)–resilient for f=h1f=h-1, we cannot replace a single correct validator to a rational one, implying that resilience breaks down when $f+k\geq 2n/3$. This result is formalized in Corollary 3.3.
Corollary 3.3.
Consider any $q$-commitable protocol $\Pi$ that is (n,f)(n,f)–resilient for f=h1f=h-1 where h=nfh=n-f is the number of correct validators. For any $k\geq 1$, $\Pi$ is not (n,k,f)(n,k,f)–resilient under our model in the synchronous setting.
In other words, no quorum-based SMR protocol can achieve (n,f+1,f)(n,f+1,f)–resilience, even in the synchronous model. Therefore, quorum-based SMR protocols are not resilient for $f+k\geq 2n/3$.
Beyond $n/3$ Byzantine.
Some SMR protocols in synchrony can tolerate $n/3\leq f^{*}<n/2$ Byzantine validators by assuming a known bound on message delays (Abraham et al., 2020). However, the guarantees of these protocols critically depend on the network respecting this synchronous bound. In contrast, $q$-committable protocols are designed to provide security under partial synchrony, independently of any synchrony bound.
Our focus is to maintain $(n,f)$-resilience even when the network is only partially synchronous. If the synchronous bound holds, the protocol can provide additional guarantees, such as increased fault tolerance, faster finality, or improved recovery guarantees. For this reason, we do not depict this family of solutions in Figure 1.
Finality latency
Theorem 3 establishes that without bounds on the liquid coins or the participation rewards, no protocol can ensure finality during a $\Delta^{*}$-bounded window when $f\geq\lceil\frac{h}{2}\rceil$ or or $q\leq f+k+\lceil\frac{h}{2}\rceil$. That is because Byzantine and rational validators can still produce valid quorum certificates for conflicting blocks within $\Delta^{*}$, with each rational validator double-spending an amount exceeding their stake. This attack remains profitable for rational validators, even if they get slashed and forfeit their stake.
{restatable}
theoremresponsiveness Let us assume participation rewards are distributed to validators, silent clients, a static PoS setting, and a synchronous network. Without any bounds i) on the liquid funds held by rational validators or malicious clients, or ii) on the participation rewards, the following holds: No SMR protocol can be both (i) (n,f)(n,f)–resilient $q$-commitable with finality latency $t$, and (ii) (n,k,f)(n,k,f)–resilient with finality latency $t^{\prime}<t+\Delta^{*}$, when $f\geq\lceil\frac{h}{2}\rceil$ or $q\leq f+k+\lfloor\frac{h}{2}\rfloor$, for h=nkfh=n-k-f, for any execution.
According to Theorem 3, if we run an (n,f)(n,f)–resilient quorum-based SMR protocol $\Pi$ with $f\geq\lceil\frac{h}{2}\rceil$ validators or $q\leq f+k+\lfloor\frac{h}{2}\rfloor$, it is impossible to finalize every transaction before $\Delta^{*}$ elapses. That means that we must limit the transaction volume finalized per any $\Delta^{*}$ time window.
4. CoBRA: Finalization Rule
In this section, we present the CoBRA finalization protocol, a transformation that converts any $(n,f)$-resilient $(2f+1)$-committable SMR protocol $\Pi$ into an $(n,k,f^{*})$-resilient proof-of-stake SMR protocol for $n=3f+1$.
Throughout this section, we assume a synchronous network with message delay bounded by $\Delta^{*}$, as required by Theorem 3. We consider executions in which the number of Byzantine and rational validators satisfy $f^{*}\leq f$ and $k\leq 2f-f^{*}$, respectively. Our protocol achieves the optimal resilience bounds established in Theorem 3. Recovery from executions exceeding this Byzantine threshold is deferred to Section 5.
Overview.
As established in Section 3, in the presence of rational validators and silent clients, quorum-based SMR protocols are vulnerable to two major attack vectors: (i) safety attacks, where rational validators collude with Byzantine ones to produce conflicting blocks with valid quorum certificates within the synchrony window $\Delta^{*}$, enabling double-spending attacks against clients; and (ii) censorship attacks, where the adversary uses accumulated participation rewards to bribe rational validators into withholding votes for blocks containing targeted transactions, thereby preventing their finalization.
To address these challenges, our protocol introduces the following components:
  • A finality gadget that ensures safety by enforcing stake-bounded finalization: within any $\Delta^{*}$-bounded window, the maximum transaction volume that can be finalized is limited by the total stake at risk; making equivocations unprofitable once slashing penalties are applied. This ensures that rational validators have no incentive to fork.
  • A reward scheme that incentivizes the participation of rational validators and mitigates censorship attacks (liveness). Transaction fees are distributed to participating validators only in execution windows in which at least nfn-f distinct validators successfully contribute to block production.
  • A strong chain mechanism that enables higher-volume finalization during periods of high participation, improving efficiency.
4.1. Finality Gadget
The finality gadget ensures that validators finalize only transactions that cannot be reversed and enables them to provide inclusion proofs to clients. Below, we describe the validator and client protocols for accepting transactions. Since clients do not monitor the protocol execution, their confirmation rule differs from that of validators. As a result, validators may finalize transactions based on protocol guarantees, whereas clients rely on additional proofs to verify finality.
Validator confirmation
We assume a $q$-committable SMR protocol $\Pi$ that is broadcast-based [^4] Otherwise, validators in our construction can broadcast the valid certificates to propagate the required information. to ensure that safety violations are detected by every correct validator within $\Delta^{*}$. We treat $\Pi$ as a black-box abstraction: whenever a validator obtains a valid quorum certificate $QC(b)$ for a block $b$ at height $h$, the protocol outputs $\mathsf{deliver}(QC(b),b,h)$ to that validator (line 10). Then, the validator stores the certificate (line 12) along with the block and the current local timestamp (line 13), and initiates the block finalization function.
Since conflicting blocks may exist but remain unwitnessed within the last $\Delta^{*}$, validators employ the stake-bounded finalization rule. More specifically, each validator $p$ traverses its local ledger and using the timestamp of the blocks and its local clock, determines which blocks are $\Delta^{*}$–old (output $\Delta^{*}$ ago or more), or $\Delta^{*}$–recent (output within the last $\Delta^{*}$) [^5] To measure elapsed time, the validator relies on its local clock rather than requiring synchronized clocks (lines 13, 36).. Then, $p$ finalizes blocks as follows.
$\Delta^{*}$–old Blocks
If no conflicting block has been detected, validators finalize any block $b^{\prime}$ that was output by $\Pi$ at least $\Delta^{*}$ ago (lines 37 and 38).
$\Delta^{*}$–recent Blocks
For blocks output within the last $\Delta^{*}$, each validator finalizes only a subset of the most recent blocks, ensuring that the total finalized transaction volume does not exceed $C=D$ coins (lines 41 and 42). This guarantees that any rational validator engaging in an attack can gain at most $D$ coins. For intuition, we provide a tight example in Figure 2 for $f^{*}=f$ Byzantine and $k=f$ rational validators.
Figure 2. Forking attack.
A subset $\mathcal{F}_{R}$ of $2f$ misbehaving validators (Byzantine and rational) fork the system. The adversary partitions all correct validators into $f+1$ distinct sets and collects their signatures to create conflicting blocks with valid certificates before $\Delta^{*}$ elapses. Each correct validator finalizes blocks with transaction volume up to $C=D$ within $\Delta^{*}$. Assuming Byzantine validators claim no rewards and instead use them to bribe the $f$ rational validators they need to successfully double-spend, then the $f$ rational validators share a total of $f\times D$ double-spent coins, resulting in a per-validator gain of $D$.
Finality votes
To enable inclusion proofs and ensure accountability, validators include finality votes for a block $b$ in subsequent blocks once $b$ is finalized. We say that $b$ has a finality certificate, denoted $\mathcal{F}(b)$, if at least $2f+1$ finality votes for $b$ (or for any block extending $b$) are recorded on-chain, with each vote included in a block carrying a valid quorum certificate.
Conflicting certificates.
Correct validators never commit finality votes for conflicting blocks. If a validator observes conflicting blocks with valid certificates or finality certificates (lines 17, 29), it immediately stops finalizing new blocks (lines 37, 41). By limiting the transaction volume so that the total potential gain for any rational validator does not exceed their stake, we can ensure that, once they are slashed, rational validators have no incentive to engage in equivocations.
Client confirmation
Clients, who do not participate directly in the consensus protocol, cannot observe when validators’ internal confirmation rules are met. The finality certificate thus serves as a verifiable proof of finality, allowing clients to confirm that a block has been finalized according to the validators’ rules.
Specifically, to confirm the inclusion of a transaction $tx$, validators provide clients with a proof consisting of:
  • An inclusion proof (e.g., a Merkle proof) for $tx$ in block $b$.
  • A finality certificate $\mathcal{F}(b)$ for the block $b$, including the finality votes and the valid quorum certificates of the blocks including these votes.
In this way, clients can securely verify transaction finalization, ensuring that the SMR protocol is certifiable.
4.2. Rewards and Penalties
Slashing mechanism
To ensure misbehaving validators are penalized, we implement a two-phase withdrawal process (Org, 2024):
  • Validators must stake $D$ coins to participate in $\Pi$.
  • Withdrawals are only allowed after a period $\Delta_{W}>2\Delta^{*}$, during which validators cannot participate in the protocol.
  • If a validator is detected signing finality votes for conflicting blocks, it is blacklisted (line 30), preventing permanently its withdrawal.
Note that the withdrawal period is set to be larger than $2\Delta^{*}$ (instead of $\Delta^{*}$). Although equivocations are observed within time $\Delta^{*}$—which is sufficient to limit the transaction volume—only after $2\Delta^{*}$ can we be certain that no conflicting block exists. For example, consider a block finalized at time tt [^6] We can fix the local clock of any validator, as only the elapsed time matters. by $2f$ rational and Byzantine validators along with one honest validator $p_{i}$. Just before time $t+\Delta^{*}$, the $2f$ malicious validators may finalize a conflicting block with the help of another honest validator $p_{j}$. In the worst case, $p_{i}$ will become aware of the conflict only at time $t+2\Delta^{*}$. Setting the extended withdrawal period larger than $2\Delta^{*}$ is sufficient to prevent malicious validators from withdrawing their funds prematurely.
Rewards
Participation rewards are sourced from transaction fees and distributed to validators based on their contributions to block finalization. Rewards are proportional to the fraction of blocks with valid certificates proposed by each validator and are distibuted every $n$ rounds. Specifically, we define the interval of rounds $[kn+1,(k+1)n]$ for $k\in\mathbb{Z}$ as the leader window $W_{k}$. After the end of the window $W_{k}$ the total transaction fees from finalized transactions $F_{W_{k}}$ are allocated based on the proportion of unique proposers $P_{W_{k}}$ relative to $n$: $R_{W_{k}}=F_{{W_{k}}}\cdot\dfrac{P_{W_{k}}}{n}.$ To ensure that correct validators successfully propose blocks, rewards are distributed only if $P_{W_{k}}\geq n-f$, i.e., at least nfn-f unique validators submit proposals within $n$ rounds.
This guarantees censorship resilience, as Byzantine validators can only earn rewards in windows that include at least one honest proposer. Even if a Byzantine validator attempts to participate normally to accumulate rewards in order to bribe rational validators later, they would eventually deplete their coins. With no coins left to bribe, rational validators prefer to simply follow the protocol and collect the participation rewards.
4.3. Extending Finalization with Strong Chains
Recall that validators finalize transaction volumes up to $D$ coins within a time window $\Delta^{*}$. This is the lower bound in our threat model. When the system is not under attack (i.e., fewer faulty nodes), blocks often receive more than $2f+1$ signatures. We leverage the certificates with additional signatures to further limit forks and finalize larger transaction volumes without waiting for $\Delta^{*}$.
To achieve this, we extend the concept of $i$-strong chains (Xiang et al., 2021) in our model. A ledger is $i$-strong if it contains valid certificates signed by $2f+1+i$ unique validators within $\Delta^{*}$. With at least $1+i$ correct validators, the maximum number of forks decreases to fif-i, allowing secure finalization of higher transaction volumes. In particular, when a certificate has $2.5f+1$ signatures, we can define a strongest chain that can finalize infinite transaction volume (regardless of $D$).
Definition 0 ($i$-strong chain).
Consider a validator $p$ and its local data structure $\text{ledger}^{p}$ (Algorithm 1, line 1). Let $S$ represent the set of distinct signatures from validators included in the valid certificates of blocks within the set of blocks in $\text{ledger}^{p}$ that were output by the SMR protocol $\Pi$ within the last $\Delta^{*}$. We say that $\text{ledger}^{p}$ is an $i$-strong chain, where i=|S|2f1i=|S|-2f-1.
Strong chain finalization rule
Validators that monitor an $i$-strong chain for $i>f/4$, can finalize transaction volume up to $C=\frac{f}{f-i}D$. Otherwise, validators finalize $C=D$ coins. Notice that for $i=f$, $C\to\infty$ because all validators participate in that ledger, allowing the finalization of any block with a valid certificate.
$C=\begin{cases}\frac{f}{f-i}D,&\text{if }i>\frac{f}{4},\\ D,&\text{else.}\end{cases}$
Strongest chain finalization rule
We additionally define the strongest chain as an i-strong chain with $i>\frac{f+1}{2}$. More than half of the correct validators participate in the strongest chain. This ensures that only one strongest chain can exist within $\Delta^{*}$, enabling correct validators to finalize every transaction, under this rule,
$C=\begin{cases}\infty,&\text{if }i>\frac{f+1}{2},\\ \frac{f}{f-i}D,&\text{if }\frac{f}{4}<i\leq\frac{f+1}{2},\\ D,&\text{otherwise.}\end{cases}$
5. CoBRA: Recovery Mechanism
In Section 4, we describe CoBRA’s finalization rule which implements an $(n,k,f^{*})$–resilient transformation for executions with $f^{*}\leq f$ Byzantine validators and $k\leq 2f-f^{*}$ rational validators. In this regime, although equivocations are possible with the participation of rational parties, they are not profitable for rational validators; thus, rational parties will not engage in forking attacks.
We now extend our protocol to executions with $f+1\leq f^{*}<5n/9$ Byzantine validators, and $k<2n/3-f^{*}$ rational validators. In this setting, equivocations cannot be prevented, since $f+1$ Byzantine validators alone suffice to create blocks with conflicting certificates within $\Delta^{*}$. Accordingly, we no longer aim to preclude safety violations; instead, we provide a recoverable economic-security regime that guarantees economic restitution with recovery parameter $\Delta_{R}=4\Delta^{*}$ (Definition 2.6).
Overview.
At a high level, CoBRA’s recovery protocol proceeds in the following steps:
  1. (1)
    Detecting conflicts: Upon observing conflicting certificates, correct validators halt further finalization to prevent additional financial damage to clients. Validators learn all client-acceptable finalized blocks by disseminating the relevant certificates containing finality votes.
  2. (2)
    Execution and reimbursement: Validators revert to a stable pre-fork state and deterministically re-execute all conflicting finalized ledgers in a common order, reimbursing provable client losses from the stake of provably misbehaving validators, without minting new coins.
  3. (3)
    New genesis: Validators issue a verifiable commitment to the recovered state, allowing the system to resume from a single consistent ledger.
5.1. Detecting Conflicts
We first bound the maximum economic damage that clients may incur and then describe how validators reliably observe all conflicting finalized blocks following an equivocation.
Economic damage
Clients accept only blocks with finality certificates. According to the finalization rules of CoBRA, the total amount of coins carrying finality votes during a fork is at most $fD$. Once the synchronous bound elapses and correct validators are aware of the equivocation they stop finalizing new blocks.
Observing conflicting certificates.
Ensuring that honest validators will witness all finalized blocks to reimburse affected clients is not straightforward. Recall that each vote in a finality certificate is included in a block with a valid SMR quorum certificate (namely an ss–phase certificate [^7] By the definition of $q$-commitable SMR protocols, a quorum certificate is formed by collecting $q$ votes across $s$ phases (e.g., $s=2$ in PBFT, $s=3$ in HotStuff).) which is broadcast to all validators. Nevetheless, malicious validators may withhold their votes for the ss–phase certificate from other validators while sending these votes directly to clients. As a result, a block may be accepted by clients even though honest validators have not observed its finality certificate.
To ensure that every block accepted by clients is eventually executed by validators, without requiring clients to disseminate certificates, validators will re-execute any block that carries $2f+1$ finality votes, provided that each vote is included in a block with at least an (s1)(s-1)–phase certificate. Whenever a block carries a valid ss–phase certificate whose finality votes form a finality certificate, at least one honest validator must have witnessed the corresponding (s1)(s\!-\!1)–phase certificates and finality votes.
Thus, upon detecting the first such violation, validators broadcast all (s1)(s\!-\!1)–phase certificates containing finality votes that form valid finality certificates but for which no corresponding ss–phase certificate has been observed (line 52). They then wait an additional $2\Delta^{*}$ (line 53) to ensure that all conflicting finalized ledgers have been revealed, as detailed in Section 4.2 (slashing mechanism). Thus, within $2\Delta^{*}$, all validators are guaranteed to have received the complete set $\mathcal{L}=\{L_{1},\ldots,L_{j}\}$, where $j\leq f+1$, which consists of all conflicting blocks whose $2f+1$ finality votes appear in at least (s1)(s\!-\!1)–phase certificates [^8] The withholding attack described above can happen only for the last block of any conflicting ledger in $\mathcal{L}$. For every intermediate block, the existence of an extending block implies that its ss–phase certificate was witnessed by an honest validator. Nevertheless, validators proactively broadcast all (s1)(s\!-\!1)–phase certificates upon equivocation to avoid waiting an additional $2\Delta^{*}$ until knowing the last block per ledger..
5.2. Execution and Reimbursement
We now describe how validators execute transactions, including those in the conflicting ledgers $\mathcal{L}$, to update the system state. Recovering a common state requires that all validators execute the same transactions in the same order, even in the presence of conflicting finalized ledgers. To this end, validators revert to the state up to a block for which no conflicting finalized block exists. To support this efficiently, validators locally maintain two distinct states.
  • Recent-finalized state: the state up to their most recent finalized block,
  • Old-finalized state: the state up to their most recent $2\Delta^{*}$–old finalized block (i.e., output $2\Delta^{*}$ ago).
Upon detecting an equivocation, validators revert to their old-finalized state (line 56), for which it is guaranteed that no fork exists. After waiting $2\Delta^{*}$, during which all conflicting finalized blocks in $\mathcal{L}=\{L_{1},\ldots,L_{j}\}$ are observed, validators update the system state.
To do so, validators first re-execute all transactions from their old-finalized state up to the first point of conflict, namely the first block in $\mathcal{L}=\{L_{1},\ldots,L_{j}\}$ (lines 56-59). For the rest blocks, they sort the conflicting ledgers in $\mathcal{L}$ lexicographically by the hash of their last finalized block, forming a new ordered set $\mathcal{L}^{*}$ (line 55). Finally, validators execute each ledger in $\mathcal{L}^{*}$ sequentially, processing all block within each ledger in order (lines 60-62).
During execution, transactions fall into two categories. First, transactions issued by honest clients who believed a branch to be canonical remain valid, since honest clients do not spend more coins that they hold. Second, transactions from malicious participants attempting to double-spend. For these, validators reimburse affected clients from the misbehaving validators’ slashed stake and burn any remaining stake to punish the malicious validators (line 63).
Algorithm 1 CoBRA protocol
[…]
9
10on SMR.deliver($QC(b),b,h$) do
11  if $quorumstore[h]=\bot$ then
12   $quorumstore[h]\leftarrow QC(b)$
13   $ledger\leftarrow ledger\cup(b,localtime)$
14   finalize-blocks()
15  else
16   $(block,\_\_\;)\leftarrow ledger[h]$
17   if $block\neq b$ then $violation\leftarrow true$
18
19on observing (“finality”, $u,b$) in a phase–(s1)(s-1) certificate do
20  finality_votes[b].add(u)finality\_votes[b].add(u)
21  if $|finality\_votes[b]|\geq 2f+1\land finalitystore[b.height]=\bot$ then
22   $finalized.apply(b)$ $\triangleright$ Execute $b$ to update the state
23   $(\_\_,time)\leftarrow ledger[b.height]$
24   if $localtime-time\geq 2\Delta^{*}\land b.height>oldFinalizedBlock.height$ then
25    $oldFinalized.apply(b)$, $oldFinalizedBlock\leftarrow b$
26  else if $|finality\_votes[b]|\geq 2f+1\land finalitystore[b.height]\neq\bot$ then
27   $(b^{\prime},\_\_)\leftarrow finalitystore[b.height]$
28   if $b^{\prime}\neq b$ then
29    $violation\leftarrow true$
30    $blacklist\leftarrow blacklist\cup(finalitystore[b.height]\cap finality\_votes[b])$
31    if this is the first time conflicting blocks are observed, Recover
32    finalitystore[b.height].add(finality_votes[b])finalitystore[b.height].add(finality\_votes[b])
33
34procedure finalize-blocks
35  $suffixvalue=0$
36  for $(block,blocktime)\in ledger$ from suffixindex to $|ledger|$ do
37   if $blocktime\leq localtime-\Delta^{*}\land\;(not\;violation)$ then
38    $block.finalize$
39    broadcast signed finality vote $v$ for block: ("finality", $v$, block)
40    $suffixindex++$
41   else if $suffixvalue+block.value\leq C\land\;(not\;violation)$ then
42    $block.finalize$
43    broadcast signed finality vote $v$ for block: ("finality", $v$, block)
44    $suffixvalue\leftarrow suffixvalue+block.value$
45   else
46    return
47
48procedure slash($(validators,state)$)
49  return $state\leftarrow\textsc{slashAll}(validators,state)$ $\triangleright$ penalize every validator in the set
50
51procedure recover
52  broadcast all (s1)(s\!-\!1)–phase certificates which form finality certificates
53  continue after $2\Delta^{*}$
54  $\mathcal{L}\leftarrow\{L_{1},...,L_{j}\}$ $\triangleright$ conflicting finalized ledgers stored in finalitystorefinalitystore
55  $\mathcal{L}^{*}\leftarrow$ $Sort(\mathcal{L})$ $\triangleright$ lexicographically sort ledgers by the hash of their last block
56  $recovered\_state\leftarrow oldFinalizedBlock$
57  for pospos from $oldFinalizedBlock.height+1$ to $|ledger|$ do
58   if $ledger[pos]\in L$ for some $L\in\mathcal{L}$ continue
59   recovered_state.apply(ledger[pos])recovered\_state.apply(ledger[pos])
60  for all ledger $L\in\mathcal{L}$ do
61   for all block $b\in L$ do
62    recovered_state.apply(brecovered\_state.apply(b)
63  $recovered\_state\leftarrow$ slash(blacklist,recovered_stateblacklist,\;recovered\_state)
64  broadcast signed state commitment to recovered_staterecovered\_state
5.3. New Genesis
After reaching to a common state, validators produce a verifiable commitment to the recovered state, allowing new participants to trust the correct state. In particular, validators sign a state commitment (e.g. Merkle commitment) for the recovered state. To ensure that only honest validators can produce this certificate, they also attach the finality votes of misbehaving validators on conflicting blocks as evidence (line 64). Once at least $f+1$ validators sign this state commitment, along with the proofs of misbehavior, the resulting certificate constitutes the new genesis block.
Since at least $f+1$ validators have provably misbehaved–and these must be Byzantine, as CoBRA ’s finalization and slashing rules make equivocation unprofitable for rational validators–the remaining population of $n^{\prime}\leq 2f$ contains an honest majority, and at least $2n^{\prime}/3$ validators in total are honest or rational. Consequently, the system can continue running CoBRA, ensuring that no further equivocations occur.
6. Security Guarantees
In this section, we provide sketch proofs and defer the full analysis to Appendix B.
Safety and client-safety
Rational validators may collude with Byzantine validators to create conflicting blocks within the maximum message delay $\Delta^{*}$. However, such attacks are unprofitable: to double-spend, misbehaving validators must present inclusion proofs for blocks with finality certificates, ensuring they are finalized by at least one correct validator. During a fork, at most $fD$ coins can be double-spent in total, and any single rational validator gains at most $D$ coins, for all finalization rules—stake-bounded, strong, or strongest chain—as shown in Lemmas B.2B.4, and B.6, respectively.
Next, we argue that any misbehaving validator will be detected and forfeit their stake. Every correct validator observes all equivocations for blocks with finality certificates within $2\Delta^{*}$ and blacklists misbehaving validators via our slashing mechanism. Hence, misbehaving validators cannot withdraw their stake (Lemma B.3).
Finally, we conclude that rational validators have no incentive to engage in forking attacks, as any potential gain does not exceed their staked amount, which is subject to slashing. This ensures the protocol’s safety (Theorem B.11). Moreover, since clients accept inclusion proofs only for blocks with at least $2f+1$ finality votes, client-safety is also guaranteed (Theorem B.12).
Liveness and client-liveness
Since transaction fees are equally distributed among block proposers, rational validators have no incentive to withhold proposals or votes for valid blocks. However, Byzantine validators may attempt to bribe them to silence correct proposers and censor transactions.
Byzantine validators claim rewards only within windows where at least $2f+1$ unique proposers exist, ensuring at least one is correct. Once their rewards are depleted, rational validators resume proposing and voting for valid blocks (Lemma B.10). Consequently, a correct validator will eventually propose a block, ensuring liveness (Theorem B.11). Additionally, to claim their rewards, rational validators commit finality votes, guaranteeing client-liveness (Theorem B.12).
Recovery and economic restitution.
After a fork, all correct validators observe the equivocation and, within $2\Delta^{*}$, witness all blocks carrying conflicting finality certificates sent to clients. By CoBRA ’s finalization rules, the total value of coins that can be double-spent across conflicting finalized blocks is bounded by $fD$. Validators provably identify at least $f+1$ misbehaving validators, whose combined stake $(f+1)D$ covers all double-spent transactions. They then deterministically execute the transactions from conflicting blocks in the same order to agree on a common state and reimburse affected clients using the slashed stake (Theorem B.14).
7. Case-Study
In this section, we present an empirical case study to evaluate the practical feasibility of CoBRA. To this end, we analyzed one million consecutive blocks from the Cosmos blockchain. Based on our analysis, over 99% of Cosmos blocks exceeded the $\frac{5}{6}$ participation threshold required by our strongest-chain finalization rule, allowing CoBRA to instantly finalize the vast majority of blocks.
The remaining blocks were produced during periods of lower participation, lasting up to 147 blocks. Given Cosmos’s $\sim$6-second block time, this corresponds to about 15 minutes. With a median daily transaction volume of \$17 million, this 15-minute window corresponds to an average transaction volume of approximately \$180,000. As such, given the minimum slashable stake of active validators of approximately \$470,000, CoBRA can compensate for low-participation periods by instantly finalizing blocks using our stake-based finalization rule.
These high participation rates are not unique to Cosmos. On Ethereum, the average daily participation has also consistently remained above 99% (Beaconscan, 2024). Thus, we conclude that CoBRA could be seamlessly integrated into production systems like Cosmos or Ethereum, enhancing system resilience without impacting block confirmation times.
8. Limitations and Extensions
Dynamic validator settings
We outline how to extend our protocol to the permissionless setting, where validators may dynamically join or leave the system. Similar to prior work on dynamic membership in BFT systems (Duan and Zhang, 2022), we assume that every membership change preserves the resilience assumptions required for safety and liveness. In our case, at any time at least one third of the active validators are honest, and between any two consecutive membership changes a quorum of at least two thirds—consisting of honest and rational validators—remains stable.
On-chain join requests. Similar to withdrawals, we treat join requests as on-chain state events. Once a join request is committed on-chain, validators update their local view of the validator set consistently. This guarantees that all existing validators agree on the current validator set, and the corresponding quorum size.
Challenges. Validator joining the system at a later time must obtain sufficient information to participate safely. In particular, late joiners must: (i) learn the current validator set, (ii) learn the current state of the system, (iii) detect past safety violations, ensuring that they neither enable malicious validators to withdraw their stake nor unknowingly contribute to the growth of a fork.
Membership and State Tracking. Late-joining validators can obtain the full chain and (i) execute the blocks to obtain the system state, (ii) follow the history of membership change events starting from the genesis block to reconstruct the current validator set. Optimizations such as introducing periodic checkpoints for the validator set and state commitments are left as future work.
Past equivocations. To ensure that late joiners do not mistakenly assist misbehaving validators, they must also learn whether any equivocation occurred in the past. We use the following mechanism:
  • Late joiner. A validator wishing to join in round $r$ broadcasts a join request and waits until round $r+2\Delta^{*}$ before participating fully. This waiting period ensures that it can receive any messages revealing past safety violations.
  • Existing validator. Any existing validator $p$ that receives the join request at some round $r^{\prime}\in\{r,r{+}\Delta^{*}\}$ responds to the joiner with all safety violations it is aware of up to round $r^{\prime}$. After that moment, $p$ treats the joining validator as active and interacts with it according to the protocol.
By round $r^{\prime}+\Delta^{*}\leq r+2\Delta^{*}$, the joining validator will have learned all equivocation evidence known to validator $p$ for rounds $\leq r^{\prime}$, and will learn about future inconsistencies during normal protocol participation.
Assessing transaction value
An important part of our protocol is assessing the value of transactions in a given block to weigh it against the locked stake. While this is simple in a payment system environment, it is not straightforward in a smart contract environment. While transactions transferring native on-chain currency can easily be evaluated, it is challenging to evaluate the value of ERC-20 tokens, NFTs, or even layer-2 rollup checkpoints.
To address this, clients can include an estimated transaction value in their submissions. If the receiver agrees with the estimate, they accept the stake-based finality rule; otherwise, they wait for the full finality threshold $2\Delta^{*}$ (or a strong chain). However, this mechanism introduces a vulnerability: a malicious client can delay the finality of other transactions by regularly issuing, or falsely declaring, high-value transactions. To mitigate this, we impose a maximum insured value per transaction, ensuring sufficient throughput.
Bribing mechanism
Trustless bribing mechanisms enable cooperation between Byzantine and rational nodes. A Byzantine node can bribe a rational validator $p$ to sign two conflicting blocks $b$ and $b^{\prime}$ (e.g., by sending pre-signed transactions). Similarly, censorship of a target client $p^{\prime}$ can be incentivized, effectively rewarding exclusion of $p^{\prime}$’s transactions. A malicious client can deploy a smart contract, funded every second block, that pays out in the next block if $p^{\prime}$’s wallet balance remains unchanged, allowing rational validators to claim the bribe only if $p^{\prime}$’s wallet is not updated.
Protocol Safety / Liveness Recovery Consistency Violations Recovery parameters Client Reimbursement
Classic BFT (Abraham et al., 2018; Buchman et al., 2018; Castro and Liskov, 2002) $n/3$ Byzantine
Accountable + Recoverable Partial Synchrony (Gong et al., 2025; Ranchal-Pedrosa and Gramoli, 2024) $n/3$ Byzantine $5n/9$ Byzantine 1 unbounded
$2n/3$ Byzantine 2
Accountable + Recoverable Synchrony (Lewis-Pye et al., 2025) $n/3$ Byzantine $5n/9$ Byzantine 1 $(f^{*}\Delta^{*},2\Delta^{*})$
$2n/3$ Byzantine 2
CoBRA (ours.)* $n/3$ Byzantine +n/3+\,n/3 rational $5n/9$ Byzantine +n/9+\,n/9 rational 1 $4\Delta^{*}$, no rollback
  • Columns: (i) resilience for safety/liveness, (ii) resilience for recovery, (iii–iv) number of consistency violations and recovery parameters ( time to achieve recovery, rollback of transactions finalized by clients), (v) client reimbursement. We denote by $\Delta^{*}$ the synchronous bound and by $f^{*}$ the actual number of Byzantine nodes.
    * Applying the recovery procedure of (Lewis-Pye and Roughgarden, 2025) on top of CoBRA ’s finalization rule (Sec. 4) guarantees client reimbursement against $2/3$ Byzantine adversary, at the cost of up to two safety violations, each requiring $O(f^{*}\cdot\Delta^{*})$ rounds.
Table 1. Comparison of classical BFT, accountable/recoverable consensus, and CoBRA.
Probabilistic network model.
In line with prior work, we assume an adversary that controls both malicious validators and network delays. We leave for future work models in which control over validators and the network is decoupled (Danezis et al., 2025) or network partitions are probabilistic (Guo et al., 2019). In such settings, attacks would succeed only with some probability, requiring the protocol to ensure that the expected utility of any attack for rational validators is strictly less than the stake they forfeit.
External incentives.
External incentives or attacks that are not observable within the blockchain ecosystem are incorporated by Byzantine behavior. By modeling a rational–Byzantine adversary, we capture a system that is resilient to both external incentives (Byzantine behavior), and internal incentives (rational behavior).
9. Related Work
Rational-Byzantine model in quorum-based protocol.
BAR (Aiyer et al., 2005) studies SMR with altruistic (honest), rational, and Byzantine parties. However, BAR is designed for general-purpose state machine replication and does not address the economic incentive structures or adversarial dynamics specific to payment systems. Moreover, BAR assumes that rational participants act independently and do not collude. Plenty of works considering rational and Byzantine parties also do not address collusions between participants (Amoussou-Guenou et al., 2019; McMenamin et al., 2021; Lev-Ari et al., 2019).
Abraham et al. (Abraham et al., 2006, 2008) introduce a solution concept which allows strategic collusion between rational and Byzantine players in the context of secret sharing and multi-party computation. Within this model, TRAP (Ranchal-Pedrosa and Gramoli, 2022) provides a protocol for one-shot consensus in the partially synchronous setting for $n>\max\left(3/2k+3f,2(k+f)\right)$. However, we show that this approach does not extend to multi-shot SMR in partial synchrony, as the total value that can be double-spent cannot be bounded. Assuming a synchronous network, we instead present an SMR protocol that tolerates up to $f+k<2/3$, improving upon TRAP’s resilience bound of $f+k<1/2$.
Accountability.
Several works focus on accountability in distributed protocols, including specific constructions (Civit et al., 2021; Buterin and Griffith, 2017; Shamis et al., 2022), analyzing forensic properties of well-known protocols (Sheng et al., 2021), and generic transformations that make protocols accountable (Civit et al., 2023). While these approaches ensure that misbehaving validators can eventually be detected, accountability alone does not guarantee rational resilience. If rational validators profit from double-spending is higher than their stake, they will misbehave even if eventually their stake is slashed. Moreover, this line of work does not address recovery after safety violations or client reimbursement. Our work closes these gaps by (i) identifying when accountability suffices to deter rational attacks and (ii) providing a recovery mechanism that ensures full reimbursement to clients when the Byzantine threshold is exceeded.
Economic safety
BLR24 (Budish et al., 2024b) examines economic safety in PoS blockchains, focusing on when slashing mechanisms can reliably punish equivocation after it occurs while ensuring honest validators retain their stake. They prove that slashing is impossible when more than $2f$ out of $3f+1$ validators misbehave, and propose a slashing mechanism when at most $2f$ do. Although our bounds align, as we also provide resilience up to $2f$ Byzantine and rational nodes, our focus is fundamentally different. BLR24 (Budish et al., 2024b) does not account for liquid coin transfers during SMR execution or explicitly model profit generation beyond slashing deterrents.
Recovery
Lewis-Pye et al. (Lewis-Pye and Roughgarden, 2025) study the impact of Byzantine adversaries controlling more than one third of the validators in quorum-based SMR. First, they show that adversaries controlling up to $5/9$ of validators can trigger at most one consistency violation, while adversaries controlling up to $2/3$ can trigger at most two. Moreover, they prove that in partial synchrony, recovery protocols cannot guarantee a bounded rollback window, implying that an unbounded number of finalized transactions may need to be reverted. As a consequence, such protocols (Ranchal-Pedrosa and Gramoli, 2024; Gong et al., 2025) cannot, in general, provide full client reimbursement and are therefore orthogonal to our approach, which targets client-level economic security.
Lastly, assuming a synchronous bound $\Delta^{*}$, they design a generic recovery protocol in which each recovery instance completes in $O(f^{*}\cdot\Delta^{*})$ rounds and rolls back transactions finalized within the previous $2\Delta^{*}$ rounds. In contrast, CoBRA provides a recovery procedure that completes within $4\Delta^{*}$, incurs at most one consistency violation, and ensures full reimbursement to all clients against an adversary controlling $5/9$ Byzantine plus $1/9$ rational validators. Notably, we can integrate the recovery procedure of (Lewis-Pye and Roughgarden, 2025) into CoBRA to tolerate up to $2/3$ Byzantine validators; full reimbursement would be still guaranteed, but recovery (up to two instances) would take $f^{*}\Delta^{*}$ rounds, essentially trading $1/9$ rational resilience for an extra $1/9$ Byzantine resilience at the cost of efficiency.
Sridhar et al. (Sridhar et al., 2023b) describe a gadget that can be layered onto synchronous SMR protocols to ensure that clients do not accept conflicting ledgers, but only under the assumption that clients actively communicate with one another and wait $3\Delta$ before finalizing transactions. In contrast, our clients remain silent: they simply accept blocks that carry finality votes. Any block meeting our finalization criteria can be finalized immediately. Moreover, (Sridhar et al., 2023b) does not specify recovery after safety violations and instead assumes that an honest majority is restored through an external process.
Algorithmic techniques
Decoupling the confirmation rules of clients and validators originates from the flexible BFT paradigm (Malkhi et al., 2019; Xiang et al., 2021; Neu et al., 2024) which separates the role of proposing and voting transactions (done by validators) with committing transactions (done by clients). We implement this separation using finality certificates for transactions and also leveraging certificates with high participation in our strong block finality rules. Finally, many protocols assume synchrony in quorum-based protocols for stronger guarantees, such as detecting equivocations (Sridhar et al., 2023a; Lewis-Pye and Roughgarden, 2025; Abraham et al., 2020), but in line with (Lewis-Pye and Roughgarden, 2025) our focus is to enable a transformation of partially synchronous protocols which can function in their assumed model without reliance on a synchrony bound.
Impossibilities
Finally, our impossibility results in partial synchrony align with (Tas et al., 2023; Budish et al., 2024b), which show that more than $1/3$ Byzantine validators can cause safety violations and avoid slashing once the security threshold is exceeded. We extend this by showing that the impossibility holds for any combination of Byzantine and rational participants exceeding $1/3$. In this setting, rational validators cannot be disincentivized from joining the attack.
Appendix A Rational attacks
In this section, we illustrate the challenges of extending (n,f)(n,f)–resilient SMR protocols towards (n,k,f)(n,k,f)–resiliency for $k\geq 1$.
Safety attacks
Conflicting ledgers and double-spending. Two blocks $b_{1}\neq b_{2}$ are conflicting if they both extend the same transaction ledger $T$. Moreover, we say that two ledgers $T_{1}$, $T_{2}$ are conflicting with each other, if there are two conflicting blocks $b_{1}\in T_{1}$, $b_{2}\in T_{2}$. Last, we say that there is a fork in the system if there are at least two conflicting ledgers $T_{1}$ and $T_{2}$.
The attack. Rational validators can collude with adversarial validators to create two conflicting blocks $b_{1}$ and $b_{2}$ with valid inclusion proofs, where they spend $\alpha$ coins to two distinct subsets of clients $C_{1}$ and $C_{2}$. For clients in $C_{1}$ (or $C_{2}$), the $\alpha$ coins have not been spent yet, and therefore, rational validators can spend them again. In that case, we say that rational validators double-spend $\alpha$ coins. We present this attack in Lemma A.1, showing that $(n,f)$-resilient blockchains are vulnerable to double-spending.
Lemma 0 (Double-Spending Attack).
Consider a system of $n=f+k+h$ validators, where $f$ are Byzantine, $k$ are rational, and $h$ are correct validators. Let $\Delta^{*}$ denote the maximum message delay (either known or not before execution). For any $q\leq f+k+\lfloor h/2\rfloor$ and $f+k>1$ and any q–commitable protocol $\Pi$ that is $(n,f)$-resilient
$\Pi$ is vulnerable to double-spending attacks.
Proof.
Assume the leader $L_{r}$ proposing the next block at some round $r$ (w.l.o.g.) is malicious. If conflicting blocks already exist, the claim holds. Otherwise, all correct validators initially agree on the same ledger $T$. A malicious leader partitions the correct validators into two subsets, $H_{1}$ and $H_{2}$, where $|H_{1}|=\lfloor h/2\rfloor$ and $|H_{2}|=\lceil h/2\rceil$. Before the bound $\Delta^{*}$ elapses, the leader sends a conflicting block $b_{1}$ to $H_{1}$ and $b_{2}$ to $H_{2}$. Rational validators sign both $b_{1}$ and $b_{2}$ because they can present inclusion proofs for $b_{1}$ and $b_{2}$ to disjoint client sets $C_{1}$ and $C_{2}$ in exchange for off-chain assets. Each conflicting block gathers $q$ signatures since $f+k+\lfloor h/2\rfloor\geq q$ and $f+k+\lceil h/2\rceil\geq q$. Therefore, there is a valid certificate for the conflicting blocks $b_{1}$ and $b_{2}$. ∎
Liveness Attacks
To guarantee liveness, q–commitable protocols typically assume at least $h\geq q$ correct validators. When this is not the case, rational validators must be incentivized to follow the protocol, introducing new challenges in aligning incentives with correct execution.
Censoring transactions. In traditional Proof-of-Stake blockchains, the majority of transaction fees are awarded to the validator that successfully proposes a block, incentivizing the inclusion of as many transactions as possible. However, if a proposer fails to produce a valid block, the pool of unprocessed transactions grows, potentially increasing the rewards available to subsequent proposers. This dynamic creates an incentive for rational validators to launch censorship attacks (Kogias et al., 2016) against competing proposers to maximize their own rewards.
Bribery attacks. Byzantine validators can use their share of the rewards to bribe rational validators and censor all correct validators. As an example, we consider a 2f+12f+1–commitable SMR protocol with $f$ Byzantine, $f$ rational, and $f+1$ correct validators. To successfully launch the attack, Byzantine validators only need to bribe a single rational validator $p$, by distributing all their rewards to $p$. In that scenario, all the rest block proposers gather at most $2f$ signatures for their proposals. Consequently, only the byzantine validators along with $p$ would successfully propose a block. As a result, the adversary would manage to censor any transaction held by correct validators.
Appendix B Analysis
We will prove that CoBRA is both a safety-preserving and liveness-preserving transformation. To this end, let $S_{f}\subseteq S_{\mathsf{bad}}$ denote the set of strategy profiles that cause a safety violation (possibly alongside a liveness violation), and define $S_{l}=S_{\mathsf{bad}}\setminus S_{f}$ as the set of strategy profiles that lead solely to liveness violations.
For each validator $p$, let $S_{f}^{p}$ (or $S_{l}^{p}$) be the set of individual strategies $s_{p}$ such that there is a strategy profile $s=(s_{p},s_{-p})\in S_{f}\textit{ (or $S_{l}$)}$ where $p$ actively contributes to the safety (or liveness) violation, and let $S_{f}^{-p}$ (or $S_{l}^{-p}$) denote the set of strategies $s_{-p}$ of all other validators in such profiles.
Safety-Preserving Transformation
First, we will show that regardless of the actions of the rest validators it is not beneficial for rational validators to engage in a forking attack. In that way, we demonstrate that for every validator $p$, and any strategy $s_{p}\in S^{p}_{f}$, there exists an alternative strategy $s_{p}^{\prime}\in S^{p}_{good}$ that weakly dominates $s_{p}$.
Liveness-preserving transformation
To establish that CoBRA is a liveness-preserving transformation, we show that at any round there will be future “good” leader windows during which at least one correct validator successfully proposes a block. Since this prevents liveness violations, we conclude that for every validator $p$ and any strategy $s_{p}\in S^{p}_{l}$, there exists an alternative strategy $s_{p}^{\prime}\in S^{p}_{\mathsf{good}}$ that weakly dominates $s_{p}$.
Combining both guarantees demonstrates that the composition of $\Pi$ and CoBRA forms a coalition compliant protocol. To formalize this, we define the utility function of the rational validators.
Utility of rational validators
To model dynamic behavior over intervals of rounds, we define the global history up to round $r$ as $H^{r}$, and the local view of player $i$ at that round as $H_{i}^{r}\subseteq H^{r}$. Let $S_{i}^{I}$ be the set of strategies available to player $i$ over an interval $I$, where each strategy is a function of their local view $H_{i}^{r}$ at each round $r\in I$. The overall strategy space for the interval is $S^{I}=\prod_{i\in[n]}S_{i}^{I}$, and the utility function of player $i$ over $I$ is given by $u_{i}^{I}:S^{I}\to\mathbb{R}$.
We now consider the set of strategies available to a rational validator $p$ over some interval $I$. First, $p$ may choose to participate in a coalition that attempts to fork the protocol. [^9] We assume a worst-case scenario: if $p$ chooses to join a coalition $\mathcal{F}$, all other validators in $\mathcal{F}$ are colluding. Otherwise, if $p$ does not engage in a forking attack, it may either follow actions that preserve liveness or abstain from doing so.
Forking. To perform this attack, a rational validator $p$ can join a coalition $\mathcal{F}\in\mathcal{F}_{R}$, where $\mathcal{F}_{R}$ consists of all the susbets of Byzantine and rational validators sufficienlty large to create a fork. We denote the partipication of $p$ in a coalition $\mathcal{F}$ during the interval $I$ with the indicator variable $b^{I}_{\mathcal{F},p}$. The function $\mathcal{M}$ gives an upper bound of the coins that can be double-spent during the attack:
$\mathcal{M}(\mathcal{F},I)=\frac{\gamma\cdot C(I)}{R(\mathcal{F})}$
, where $\gamma$ is the maximum number of forks that can be created by the coalition, $C(I)$ is an upper bound of coins finalized per fork during $I$ as specified by the protocol, and $R(\mathcal{F})$ returns the number of rational validators in the coalition [^10] To provide stronger guarantees, we assume that all the rewards from double-spending are distributed exclusively and equally to the rational validators..
Moreover, we denote the participation of $p$ in at least one coalition $\mathcal{F}\in\mathcal{F}_{R}$ by the indicator variable $b^{I}_{p}$. Let $\mathbf{b}^{I}_{p}$ be the vector whose elements are all indicator variables $b^{I}_{\mathcal{F},p}$ for every $\mathcal{F}\in\mathcal{F}_{R}$ and let $P$ be the slashing penalty imposed on $p$ (which is equal to the stake $D$).
Liveness-preserving actions. If validator $p$ does not join any coalition, it may instead follow the protocol’s liveness-preserving actions—such as proposing valid blocks when selected as leader, voting for valid blocks, participating in view-change procedures, and issuing finality votes for blocks committed to its local ledger.
We denote by $\sigma^{I}_{p}$ the strategy of validator $p$ over the interval $I$ with respect to these liveness-preserving actions. Let $R(\sigma^{I}_{p})$ denote the total participation rewards that $p$ earns by following $\sigma^{I}_{p}$, and let $B(\sigma^{I}_{p})$ denote any bribes received from an adversary to refrain from signing specific blocks.
Finally, the utility of $p$ in the interval $I$ while following the strategy $s_{p}=(\mathbf{b}^{I}_{p},\sigma^{I}_{p})$ is expressed as:
$u_{p}^{I}(s_{p},s_{-p})=\begin{cases}\sum_{\mathcal{F}\in\mathcal{F}_{R}}b^{I}_{\mathcal{F},p}\cdot\mathcal{M}(\mathcal{F,I})-P&\text{if $b^{I}_{p}=1$},\\ R(\sigma^{I}_{p})+B(\sigma^{I}_{p})&\text{otherwise}.\end{cases}$
B.1. Lemmas for proving safety
In Lemma B.1, we prove that once a correct validator finalizes a $2\Delta^{*}$-old block $b$, no conflicting block exists. Thus, conflicting block can exist only for $2\Delta^{*}$-recent blocks. We will use this lemma to establish the recovery guarantees of CoBRA.
Lemma 0.
Assume a correct validator finalizes a block $b_{1}$ that is $2\Delta^{*}$-old. Then, no correct validator will ever finalize a block $b_{2}$ conflicting to $b_{1}$.
Proof.
Assume, for the sake of contradiction, that the following scenario occurs. A correct validator $p_{i}$ includes block $b_{1}$ in its local ledger at time $localTime_{b_{1}}$ and triggers $b_{1}.\mathsf{finalize}$ at time $localTime$, where $localTime\geq localTime_{b_{1}}+2\Delta^{*}$. Furthermore, another correct validator $p_{j}$ triggers $b_{2}.\mathsf{finalize}$ for a block $b_{2}$ that conflicts with $b_{1}$.
By assumption, at $localTime_{b_{1}}$, $p_{i}$ broadcasts the valid certificate $QC(b_{1})$ for block $b_{1}$. Thus, $p_{j}$ receives $QC(b_{1})$ at some time $t^{*}\in\{localTime_{b_{1}},\dots,localTime_{b_{1}}+\Delta^{*}\}$. We distinguish two cases:
Case 1: $p_{j}$ triggers $b_{2}.\mathsf{finalize}$ at time $t<t^{*}$. Then $p_{j}$ must have received $QC(b_{2})$ at some time $t^{\prime}\leq t<t^{*}$. Since $p_{j}$ broadcasts $QC(b_{2})$, validator $p_{i}$ receives it by some time $localTime_{b_{2}}\leq t+\Delta^{*}<localTime_{b_{1}}+2\Delta^{*}$. At time $localTime_{b_{2}}$, $p_{i}$ detects the equivocation and slashes the misbehaving validators (line 63), resulting in a non-empty blacklist: $QC(b_{1})\cap QC(b_{2})\neq\emptyset$. Therefore, at time $localTime$, the finalization condition (line 37) is not satisfied, and $p_{i}$ will not finalize $b_{1}$ (line 38), contradicting the assumption.
Case 2: $p_{j}$ triggers $b_{2}.\mathsf{finalize}$ at time $t\geq t^{*}$. At time $t$ $p_{j}$ has received both $QC(b_{1})$ and $QC(b_{2})$. Let $t_{b_{2}}$ and $t_{b_{1}}$ be the times at which $p_{j}$ received $QC(b_{2})$ and $QC(b_{1})$, respectively, and assume without loss of generality that $t_{b_{1}}\geq t_{b_{2}}$. At time $t_{b_{1}}$, $p_{j}$ detects the equivocation and slashes the misbehaving validators (line 63), resulting in $QC(b_{1})\cap QC(b_{2})\neq\emptyset$. Therefore, at time $t$, the finalization condition (line 37) is not satisfied, and $p_{j}$ will not finalize $b_{2}$ (line 38), leading to a contradiction.
In Lemma B.2, we show that when validators employ the stake-bounded finalization rule, any rational validator engaging in a forking attack can gain at most $D$ coins.
Lemma 0.
Assume that validators employ the stake-bounded finalization rule and that at least one coalition of validators forks the system in some round.
  1. (1)
    Upper bound of double-spending: The coalition double spends up to $fD$ coins and any rational validator participating in the attack can double-spend at most an amount of coins equal to the stake $D$.
  2. (2)
    Tight example: There is a configuration where a rational validator in the coalition double-spends exactly $D$ coins.
Proof.
(1) Upper bound of double-spending. Misbehaving validators can double-spend only coins in conflicting blocks that are accepted by clients. Any inclusion proof $\pi_{tx}$ for a transaction $tx$ included in a block $b$ is accepted by some client $c$ only after witnessing a finality certificate $\mathcal{F}(b)$ including at least $2f+1$ finality votes for $b$. Therefore at least one correct validator $p_{i}$ has finalized block $b$. Since correct validators do not trigger finalizefinalize for conflicting blocks, and at least one correct validator is required to finalize a block, there can be at most $f+1$ conflicting ledgers.
Consider the time tt [^11] To refer at a specific time, we can fix the local clock of any validators, as only the time elapsed matters. where the first conflicting blocks with valid certificates are formed. Moreover, assume the worst case, where there are $f+1$ conflicting ledgers and let us fix any rational validator $p$ participating in f+1mf+1-m conflicting ledgers $L_{i\in[f+1-m]}$ with $m\in[f+1]$. The coins that misbehaving validators double-spend on these conflicting ledgers is upper bounded by the total coins spent in conflicting blocks finalized by at least one correct validator.
To bound this, we fix $L_{j}$ to be the longest ledger (i.e., the ledger with the most finalized blocks) at time $t$; if all ledgers have the same length, we fix $L_{j}$ for any $j\in[f+1-m]$. At time $t+\Delta^{*}$ every honest validator will have witnessed the equivocation and stop voting for new blocks. To upper bound the double spent coins during $t,t+\Delta^{*}]$ we fix some $L_{i}$ for any $i\in[f+1-m],i\neq j$ and upper bound the coins that finalized by at least one correct validator in $L_{i}$.
Let $Suffix^{[t,t+\Delta^{*}]}_{L_{i}}$ be the set of all blocks with valid certificates in $L_{i}$ during $t,t+\Delta^{*}]$. For every block $b\in Suffix^{[t,t+\Delta^{*}]}_{L_{i}}$, $p_{i}$ updates the variable suffixvaluesuffixvalue with the total output sum of transactions in the block $b$ and only finalizes $b$ if suffixvaluesuffixvalue is at most $C=D$ (line 41). Therefore, the amount of coins spent in $L_{i}$ is at most $D$.
Now, we will upper bound the total amount of coins validator $p$ can double-spend in all the f+1mf+1-m conflicting ledgers. First, at most $m+1$ correct validators participate in each ledger. We prove that by reaching a contradiction. Assume there is a ledger with at least $m+2$ correct validators. Then, in the remaining fmf-m ledgers, there must be at least one correct validator. Therefore, there are, in total, at least m+2+fm=f+2m+2+f-m=f+2 correct validators in the system. We reached a contradiction. Since there are at most $f$ byzantine and $m+1$ correct validators per ledger, at least fmf-m rational validators participate in each ledger. Therefore, at least fmf-m rational validators can double-spend at most $C=D$ coins in fmf-m ledgers, and since they distribute the benefit equally, each validator double-spends $max_{m\in[f+1]}\frac{f-m}{f-m}D=D$ coins.
(2) Tight example. Consider a run with $f$ byzantine, $f$ rational, $f+1$ correct validators. Moreover, consider as input a set of transactions $Tx=\cup_{i=1}^{f+1}Tx_{i}$ constructed as follows. There is a partition of clients in distinct sets $C=\cup_{i=1}^{f+1}C_{i}$ s.t. for each $i$, in $Tx_{i}$, each of the $f$ rational validators transfers $D$ coins to a subset of clients in $C_{i}$.
At some time, a malicious (Byzantine or rational) validator $p$ is the block proposer. Within the last $\Delta^{*}$, the adversary partitions the $f+1$ correct validators in the distinct sets $H_{i}$ for $i\in[f+1]$. Now, $p$ constructs the blocks $b_{i\in[f+1]}$, where each $b_{i}$ includes the transactions in $Tx_{i}$. Validator $p$ sends the block $b_{i}$, only to the correct node in $H_{i}$ and successfully generates a valid certificate for $b_{i}$ along with the respective finality votes which will present to the clients in $C_{i}$. Therefore, each rational validator spends the $D$ coins in the $f+1$ conflicting ledgers. In total, each rational validator double-spends $\frac{f*D}{f}=D$ coins.
Now, we argue that any misbehaving validator will be detected and forfeit their stake. We emphasize that we consider forks with conflicting finality certificates—not merely valid quorum certificates—as only these can be double-spent and yield a profit for rational participants.
Lemma 0.
Assume a coalition of validators $\mathcal{F}$ forks the system. No validator in $\mathcal{F}$ can withdraw its stake.
Proof.
Assume that the validators in $\mathcal{F}$ fork the system at time tt [^12] We refer to the first round when distinct correct validators observe conflicting blocks with valid certificates. As before, we use the local clock of any validator to refer to time elapsed between events. resulting in $j\in[f+1]$ conflicting ledgers $L_{i\in\{1,...,j\}}$ with finality certificates. Let us fix any validator $p\in\mathcal{F}$. We will show that $p$ cannot withdraw its stake in any of the conflicting ledgers.
Consider the first ledger $L_{i}$ in which validator $p$’s withdrawal request $R_{p}$ is included in a block $b$ at time $t^{\prime}$. We argue that $t\leq t^{\prime}+\Delta^{*}$. Suppose, for contradiction, that $t>t^{\prime}+\Delta^{*}$. Since block $b$ has a valid certificate, it must have been witnessed by at least one correct validator $p_{j}$, who then broadcasts $R_{p}$ at time $t^{\prime}$. Therefore all correct validators receive $R_{p}$ by time $t^{\prime}+\Delta^{*}$ and ignore the participation of $p$ at time $t$. This contradicts the assumption that $p$ participated in the forking attack at time $t$.
To complete the withdrawal, a second transaction $R_{p,s}$ must be committed to $L_{i}$ at a time $t^{\prime}+2\Delta^{*}$. Since $L_{i}$ is the first ledger to include $R_{p}$, the same condition applies to any other conflicting ledger, namely the withdrawal can only be finalized after $t^{\prime}+2\Delta^{*}$. However, since correct validators broadcast valid certificates, every correct validator will witness all conflicting ledgers at some time $t^{*}\leq t+\Delta^{*}\leq t^{\prime}+2\Delta^{*}$. As a result, they will detect equivocation and blacklist $p$, preventing the finalization of $R_{p,s}$. Thus, $p$’s withdrawal cannot succeed in any of the conflicting ledgers.
Lemma 0.
Assume that correct validators employ the Strong Chain finalization rule and at least one coalition of validators forks the system. The coalition double spends up to $fD$ coins and any rational validator participating in the attack can double spend at most $D$ coins.
Proof.
Similar to Lemma B.2, we will fix the longest ledger at the time of the equivocation and upper bound the total amount of coins in conflicting blocks that are finalized by at least one validator during $\Delta^{*}$.
Case 1: Every correct validator finalizes up to $D$. The proof is similar to Lemma B.2.
Case 2: At least one correct validator $p_{i}$ monitors an $i$-strong chain $L_{p_{i}}$ with $i>f/4$. Since $1+i$ correct validators monitor $L_{p_{i}}$, there can be at most fif-i additional conflicting ledgers. We assume the worst case, which is that fif-i conflicting ledgers exist. We fix any rational validator $p$ engaging in the attack, participating in fmif-m-i conflicting ledgers $L_{i\in[f-m-i]}$, with $m\in[f-i]$.
For each $i\in[f-m-i]$, at most $m+1$ correct validators and therefore at least fmf-m rational validators monitor ledger $L_{i}$. We prove that by contradiction. Assume there is a ledger $L_{i}$ with $m+2$ correct validators. At least $i+1$ correct validators monitor $L_{p_{i}}$, at least $m+2$ correct validators monitor $L_{i}$ and at least one correct validator exists in the rest fmi1f-m-i-1 conflicting ledgers $L_{j\neq i}$. In total, this accounts for $f+2$ correct validators, leading to a contradiction. Furthermore, since there are at most $m+1$ correct validators per ledger, the total amount of coins finalized by correct validators on any ledger is at most $\frac{f}{f-m}D$.
Finally, there are fmif-m-i conflicting ledgers with at most fmf-m rational validators spending up to $\frac{f}{f-m}D$ per ledger. So, the amount of double-spent coins is bounded by $max_{\{m\in[f-i]\}}\frac{(f-m-i)f}{(f-m)^{2}}D$ which is bounded by $D$ for $i\in\{\frac{f}{4},...,f\}$. We can show that, by taking the continuous extension of the function $F(m)=\frac{(f-m-i)f}{(f-m)^{2}}D$ for $m\in[0,f-i]$ and some constant $i\in\{\lfloor\frac{f}{4}+1\rfloor,...,f\}$, which takes its maximum value for m=f2im=f-2i. Since the function takes maximum value at an integer value, the maximum will be the same for the discrete function and thus $\frac{f}{4i}D\leq D$.
Lemma 0.
At any round, there can only exist a unique strongest chain.
Proof.
For the sake of contradiction, assume that at some round $r$, there exist two conflicting chains $T_{1}$, $T_{2}$ that both form a strongest chain. That means that $T_{1}$ is an $i_{1}$-strong chain for some $i_{1}>\frac{f-1}{2}$ and $T_{2}$ is an $i_{2}$-strong chain for some $i_{2}>\frac{f-1}{2}$. Therefore, there exist two distinct subsets $H_{1}$ and $H_{2}$ that consist of $1+i_{1}$ correct validators participating in $T_{1}$ and $1+i_{2}$ correct validators participating in $T_{2}$ respectively. However, $1+i_{1}+1+i_{2}>f+1$. Since there are $f+1$ only correct validators in the system, we reached a contradiction. ∎
Lemma 0.
Assume that correct validators employ the Strongest Chain finalization rule and at least one coalition of validators forks the system. The coalition double spends up to $fD$ coins and any rational validator participating in the attack can double-spend at most $D$ coins.
Proof.
We will prove that when validators employ the strongest chain finalization rule, rational validators participating in coalitions can double-spend up to $D$ coins. Similar to Lemma B.2, we fix the longest ledger at the time of the equivocation and provide an upper bound on the total amount of coins included in conflicting blocks that are finalized by at least one validator within the interval $\Delta^{*}$.
Case 1: There is no correct validator $p_{i}$ monitoring an $i_{1}$-strong chain with $i_{1}>(f+1)/2$. If every correct validator finalizes up to $D$, the proof follows from Lemma B.2. If at least one correct validator monitors an $i_{1}$-strong chain with $i_{1}>f/4$, the proof follows from Lemma B.4.
Case 2: At least one correct validator $p_{i}$ monitors an $i_{1}$-strong chain with $i_{1}>(f+1)/2$. According to Lemma B.5, only one strongest chain can exist. Thus we have the following cases.
Case 2a: Every correct validator in a conflicting ledger, finalizes up to $D$. A number of $1+i_{1}$ correct validators monitor $L_{p_{i}}$, and thus there are up to $f-i_{1}$ extra conflicting ledgers. Assume that there are $f-i_{1}$ extra conflicting ledgers and let us any rational validator participating in $f-m-i_{1}$ of the conflicting ledgers for $m\in[f-i]$. In each fork, there are at most $m+1$ correct validators and at least fmf-m rational validators (with the same analysis done in Lemma B.2). Therefore, rational misbehaving validators can double-spend up to $max_{\{m\in[f-i_{1}]\}}\frac{(f-m-i_{1})}{(f-m)}D\leq{m\in[f-i_{1}]\}}\frac{(f-m)}{(f-m)}D=D$.
Case 2b: At least one correct validator $p_{j}$ monitors an $i_{2}$-strong chain $L_{p_{j}}$ with $i_{2}>f/4$.. Let $i_{1}^{*}=i_{1}+1$ and $i_{2}^{*}=i_{2}+1$. In that scenario, at least $i_{1}^{*}>\lceil\frac{f+3}{2}\rceil$ correct validators monitor $L_{p_{i}}$ and at least $i_{2}^{*}$ correct validators, where $\lceil\frac{f+4}{4}\rceil<i_{2}^{*}<\lceil\frac{f-1}{2}\rceil$, monitor ledger $L_{p_{j}}$. Therefore, there are at $i_{3}=f+1-i_{1}^{*}-i_{2}^{*}<\frac{f-6}{4}<\frac{f}{4}$ remaining correct validators to participate in the rest conflicting ledgers. Thus, the there are up to $i_{3}$ conflicting ledgers, where correct validators finalize up to $D$ coins (since there is a participation of less than $f/4$ correct validators). Assume the worst case, where there are $i_{3}$ such conflicting ledgers $L_{i\in[i_{3}]}$.
Now, we fix any rational validator $p$ participating in $L_{p_{i}}$ and in $i_{3}-m$ of the conflicting ledgers $L_{i\in[i_{3}-m]}$ for $m\in[i_{3}]$. We will upper bound the coins that $p$ can spend in these conflicting blocks. First, in $L_{p_{j}}$, there are finalized up to $\frac{f}{(f-m)(f-i_{2})}D<A(m)=\frac{f}{(f-m)(f-\frac{f-1}{2})}D$. For the conflicting ledgers $L_{i\in[i_{3}-m]}$, we can show that there are at most $m+1$ correct validators and at least fmf-m rational validators per fork, with the same analysis done in B.2. Thus, each validator can double-spend up to $\frac{f+1-i_{1}^{*}-i_{2}^{*}-m}{f-m}D=(1+\frac{1-i_{1}^{*}-i_{2}^{*}}{f-m})D\leq B(m)=1-\frac{3(f+2)}{4(f-m)})D$. Now we need to upper bound $C(m)=A(m)+B(m)=(1-\frac{3f^{2}+f+6}{4(f-m)(f+1)})D<D$ since $m\in[f+1-i_{1}^{*}-i_{2}^{*}]$ and $f>0$.
Finally, in Lemmas B.7B.8, we conclude that rational validators have no incentive to engage in forking attacks, as any potential gain does not exceed their staked amount, which is subject to slashing.
Lemma 0.
When correct validators use the i) Stake-bounded finalization rule, or ii) the Strong Chain finalization rule, or iii) the Strongest Chain finalization rule, for any rational validator, following the protocol weakly dominates participating in a coalition to fork the system.
Proof.
Let us fix any rational validator $p$ and compute its utility to join coalitions forking the system. In Lemmas B.2B.4B.6, we showed that by performing the attack, $p$ can double-spend at most $D$ coins when validators employ i) finalization rule presented in 4, ii) strong ledger finalization rule, iii) strongest finalization rule respectively. According to Lemma B.3, $p_{j}$ cannot withdraw its stake $D$ before the conflicting ledgers are detected. Moreover, upon detecting equivocation the system halts and thus misbehaving validators cannot utilize further bribes. Therefore, for any strategy $s_{p}\in S^{p}_{\mathsf{f}}$ and $s_{-p}\in S^{-p}_{\mathsf{f}}$, the utility of $p$ is $u_{p}(s_{p},s_{-p})\leq 0$, while following the protocol yields non-negative utility.
To show that following the protocol weakly dominates the strategy of forking the system, we will construct a configuration where $p$ chooses a strategy $s_{p}\in S^{p}_{\mathsf{f}}$ and fix the rest strategies $s_{-p}\in S^{-p}_{\mathsf{f}}$ such that $u_{p}(s_{p},s_{-p})<0$ (this is sufficient since following the protocol has non-negative utility). To this end, suppose that $p$ participates in a coalition $\mathcal{F}$ consisting of $f$ Byzantine and $\lceil\frac{f+1}{2}\rceil$ rational validators. Moreover, consider a subset $H_{1}$ of $\lceil\frac{f+1}{2}\rceil$ correct validators, and a subset $H_{2}$ distinct to $H_{1}$ that consists of $\lfloor\frac{f+1}{2}\rfloor$ correct validators. Validators in $\mathcal{F}$ construct to conflicting ledgers $T_{1}$ and $T_{2}$ as described in Lemma A.1. The utility of $p$ in that case is $u_{p}(s_{p},s_{-p})\leq\frac{D}{\lceil\frac{f+1}{2}\rceil}-D<0$.
Lemma 0.
A fork of the system will never occur.
Proof.
For the sake of contradiction assume that two distinct correct validators $p_{1}$, $p_{2}$ finalize the conflicting blocks $b_{1}$ and $b_{2}$ respectively. Validator $p_{1}$ finalized $b_{1}$ after witnessing the respective valid certificate $QC(b_{1})$, and $p_{2}$ finalized $b_{2}$ after witnessing $QC(b_{2})$. By quorum intersection there exists a validator $p$ that is not Byzantine and has voted in both $QC(b_{1})$ and $QC(b_{2})$. Since correct validators do not equivocate, $p$ must be rational. However, according to Lemma B.7, for rational validators, following the protocol weakly dominates participating in a coalition that forks the system. So, $p$ will not sign for conflicting blocks and thus we reached a contradiction.
B.2. Lemmas for proving liveness
Since $\Pi$ is $(n,f)$-resilient and $f+1+k\geq n-f$, liveness is ensured as long as rational validators participate in liveness-preserving actions during some intervals. In Lemma B.9, we show that there exist windows during which it is a dominant strategy for rational validators to follow the protocol specifications. Building on this, Lemma B.10 concludes that some correct validator will successfully propose a block that gathers a valid certificate and finality votes.
Lemma 0.
For any round $r$, there is a leader window $W_{k}$ starting at some round $r^{\prime}\geq r$, such that following the protocol in every round of $W_{k}$ is a weakly dominant strategy for any rational validator.
Proof.
Assume for contradiction that in every window $W_{m}$ after round $r$, there exists at least one rational validator who prefers a strategy that deviates from the protocol. Since following the protocol yields a non-negative utility, any such deviating strategy must yield strictly positive utility during window $W_{m}$. Rational validators can deviate as follows:
Forking. From Lemma B.7, any strategy in the strategy profile $S^{p}_{f}$ (i.e., where $p$ actively contributes to a fork), is weakly dominated by following the protocol. Therefore, no rational validator will engage in forking attacks.
Only abstaining from liveness actions. A rational validator will only choose to abstain if the adversary offers a bribe using their participation rewards from previous rounds that outweighs the utility from honest participation.
Suppose the adversary is able to bribe at least one rational validator in every window $W_{m}$ following round $r$. In such windows, at most $2f$ validators can successfully propose valid blocks. Since participation rewards are distributed only when at least 2f+12f{+}1 distinct validators succeed, no rewards are paid out in these bribed windows. Because the adversary’s budget is finite, it is eventually depleted and the adversary cannot bribe any rational validator anymore. Therefore, there exists a future window $W_{k}$ (starting at some round $r^{\prime}\geq r$) such that, for any validator $p$, any strategy $s_{p}\in S_{p}$, and any $s_{-p}\in S_{-p}$ under which $p$ abstains from liveness actions during $W_{k}$, we have $u_{p}(s_{p},s_{-p})\leq 0$.
To show that following the protocol during $W_{k}$ is a weakly dominant strategy, fix any rational validator $p$, and consider two strategies: $s_{1}$, where $p$ follows the protocol during $W_{k}$, and $s_{2}$, where $p$ abstains from liveness actions during $W_{k}$ (we have already established the result for forking in B.7). Let $s_{-p}$ be a strategy profile in which at least $2f$ other validators follow the protocol during $W_{k}$. In the combined strategy profile $s_{a}=(s_{1},s_{-p})$, at least $2f+1$ validators follow the protocol, satisfying the reward condition, and $p$ receives the corresponding participation rewards. In contrast, under $s_{b}=(s_{2},s_{-p})$, only $2f$ validators follow the protocol, so no rewards are distributed. Hence, $u^{W_{k}}_{p}(s_{a})>u^{W_{k}}_{p}(s_{b})$.
Lemma 0.
For any round $r$, there exists a round $r^{*}\geq r$ at which a correct validator proposes a block $b$ such that:
  1. (1)
    $b$ is finalized by every correct validator, and
  2. (2)
    $2f+1$ finality votes for $b$ or a subsequent block are included in blocks with valid certificates.
Proof.
According to Lemma B.9, for any round $r$, there exists a round $r^{\prime}\geq r$ such that a time window $W_{k}$ begins, during which it is a weakly dominant strategy for rational validators to follow the protocol. Consequently, some block $b$ proposed by an honest validator during $W_{k}$ will obtain a valid quorum certificate $QC(b)$ and receive at least $2f+1$ finality votes (either for $b$ or for one of its descendants).
B.3. Final Theorems
Theorem B.11.
Validators running the composition of an $(3f+1,f)$-resilient, (2f+1)(2f+1)–commitable SMR protocol $\Pi$ and CoBRA agree on a total order in all executions with $f^{*}\leq f$ and $k\leq 2f-f^{*}$.
Proof.
Safety. In Lemma B.8, we show that validators will never finalize conflicting blocks.
Liveness. Consider any transaction $tx$ witnessed by all correct validators at some round $r$. According to Lemma B.10, there is a round $r^{*}\geq r$ where a correct validator $p$ will successfully propose a valid block $b$. If $tx$ has not already been included in a valid block proposed before round $r^{*}$, then, $p$ will include $tx$ at block $b$.
Theorem B.12.
The composition of an $(3f+1,f)$-resilient, (2f+1)(2f+1)–commitable SMR protocol $\Pi$ and CoBRA satisfies certifiability in all executions with $f^{*}\leq f$ and $k\leq 2f-f^{*}$.
Proof.
Client-safety. Consider any client $c$. We will show that any transaction $tx$ for which $c$ has accepted an inclusion proof is included in the transaction ledger of at least one correct validator.
To this end, consider the inclusion proof $\pi_{tx}$ for the transaction $tx$. Client $c$ has accepted $\pi_{tx}$ only after receiving a proof showing the inclusion of $tx$ in a block $b$, where $b$ has $2f+1$ finality votes. At least one correct validator $p$ has committed a finality vote for $b$. Validator $p$ commits finality votes only for blocks included in its transaction ledger.
Client-liveness. For every transaction $tx$ committed to a block $b$ accepted by all correct validators, any correct validator $p_{i}$ can construct a proof $\pi$ as follows. Validator $p_{i}$ sends an inclusion proof $\pi_{tx}$ of $tx$ in $b$ to a client $c$, along with a proof $\pi_{f}(b)$ consisting of $2f+1$ finality votes for $b$ or subsequent block, which is feasible according to Lemma B.10. Any client $c$ will accept the proof $\pi$ by definition of the client confirmation rule.
Theorem B.13.
The composition of CoBRA and a (2f+1)(2f\!+\!1)–committable protocol $\Pi$ is an $(n,k,f^{*})$-resilient SMR protocol in all executions with $f^{*}\leq f$ and $k\leq 2f-f^{*}$.
Proof.
Validators running the composition of CoBRA and $\Pi$ agree in a total order according to Theorem B.11. Moreover, in Theorem B.12 we show that the composition of CoBRA and $\Pi$ satisfies certifiability.
Theorem B.14.
The composition of an $(3f+1,f)$-resilient, (2f+1)(2f+1)–commitable SMR protocol $\Pi$ and CoBRA regains economic restitution with recovery parameter $\Delta_{R}=4\Delta^{*}$ in all executions with $f^{*}<5n/9$ and, $k<2n/3-f^{*}$.
Proof.
Assume an equivocation occurs, i.e., conflicting blocks containing finality votes are formed, and let time tt [^13] We fix the local clock fo any participant, as only elapsed time matters. be the first time at which finality certificates for conflicting blocks are formed.
Clients accept a block $b$ only if it carries a valid finality certificate $\mathcal{F}(b)$. Finality votes forming $\mathcal{F}(b)$ are included only in a subset of blocks $\mathcal{B}$ where each block in $\mathcal{B}$ carries a valid quorum certificate. Since $\mathcal{F}(b)$ contains $2f+1$ finality votes and honest validators issue finality votes only for blocks whose certificates satisfy the finalization rule, at least one honest validator must have witnessed a valid certificate for each conflicting block and broadcast the corresponding certificate. Therefore, at time $t+\Delta^{*}$ every honest validator will have witnessed the first conflicting certificates votes and stop finalizing new blocks (line 17 or 29). Thus, the last conflicting blocks must have been created before $t+\Delta^{*}$ and will be witnessed by all validatots by time $t+2\Delta^{*}$.
We now argue that all honest validators will witness the same set of conflicting finalized ledgers $\mathcal{L}$. Every block whose finality votes are included in valid quorum certificates observable by at least one honest validator will eventually be observed by all honest validators, since this honest validator broadcasts the certificate. We now consider the case where quorum certificates are initially withheld by Byzantine validators. Since the $2f+1$ finality votes forming a finality certificate are included in a block with an ss–phase quorum certificate, and honest validators vote for an ss–phase certificate only after witnessing a complete (s1)(s\!-\!1)–phase certificate for the same block, at least one honest validator must have observed the corresponding (s1)(s\!-\!1)–phase certificate for every finalized block and broadcast it (line 52). Consequently, even if some ss–phase certificates were initially withheld, within $2\Delta^{*}$ all honest validators receive the complete set of finalized blocks accepted by clients and construct the same set of conflicting ledgers $\mathcal{L}$.
Each validator returns to the state of its last $2\Delta^{*}$-finalized block; by Lemma B.1, no block conflicting with this block exists. After witnessing an equivocation, validators then wait at least $2\Delta^{*}$ before executing transactions. Consequently, each validator update their state only at some time $t^{\prime}\in[t+2\Delta^{*},t+3\Delta^{*}]$, by which time they have observed all conflicting blocks with finality certificates in $\mathcal{L}$. Using the same deterministic rule (e.g., ordering by the hash of the last finalized block), all validators execute transactions from the conflicting ledgers in the same order.
Any transaction issued by an honest sender is executed successfully, since the sender holds the corresponding coins in the respective ledger. Transactions whose recipients lack sufficient balance correspond to double-spent coins. By Lemmas B.2B.4B.6, within the interval $t,\,t+\Delta^{*}]$, at most $fD$ coins are double-spent in total. Because, all honest validators identify (the same) at least $f+1$ misbehaving validators, these coins are covered by the stake of misbehaving validators which is in total $(f+1)D>fD$; any remaining stake of misbehaving validators is burned (line 63).
Thus, all honest validators compute the same resulting state and sign the same state commitment by time $t+3\Delta^{*}$. In particular, at least $f+1$ honest validators sign the state commitment which they broadcast to every validator (line 64). Thus every honest validator will have received the verifiable state commitment by time $4\Delta^{*}$. In the resulting state, all affected clients are reimbursed and all misbehaving validators are penalized.
Appendix C Impossibility results
C.1. Proof Sketches
Proof sketch of Theorem 3. The adversary can partition correct validators in two distinct subsets $H_{1}$, $H_{2}$ and perform the double-spending attack described in Lemma A.1.This attack can be beneficial for rational validators since for any value of stake $D$ they lock before the run, there is an input of conflicting transactions $T_{1}$ and $T_{2}$ where rational validators transfer all their liquid coins to distinct subset of clients $C_{1}$ and $C_{2}$ respectively. Notably, misbehaving validators can withdraw their stake without even being slashed.
Let us fix the stake $D$ and any withdrawal period $\Delta_{w}$. Consider as input a set of conflicting transactions $T_{1}$ and $T_{2}$ constructed as explained before. To guarantee liveness, correct validators must finalize any input of valid transactions. Therefore, validators in $H_{1}$ must finalize transactions in $T_{1}$, and validators in $H_{2}$ must finalize transactions in $T_{2}$.
However, since $GST$ occurs at an unknown time, it is impossible for correct validators in different subsets to detect the conflicting blocks including transactions in $T_{1},T_{2}$ before finalizing the transactions. Similarly, it impossible for the correct validators to slash the misbehaving ones before they withdraw their stake. ∎
Proof sketch of Theorem 3. We modify the attack described in Theorem 3 as follows. For any stake value $D$ locked by rational validators before the execution, there exists an input consisting of conflicting transactions $T_{1}$ and $T_{2}$ such that rational validators transfer a total of $qD$ coins to clients in $C_{1}$ and $C_{2}$, respectively, and hence double-spend $qD$ coins. Even if the attack is eventually detected and the misbehaving validators are slashed and forfeit their stake, or even if the stake of misbehaving validators is instead offered as a reward for reporting misbehavior, the attack remains strictly beneficial and therefore preferable for rational validators.
Such a configuration can be constructed whenever at least one of the following resources is unbounded: (i) the initial liquid coins held by rational validators, (ii) the coins available as bribes from malicious clients, or (iii) the participation rewards. In particular, when participation rewards are unbounded, rational validators can repeatedly reclaim their initial funds through participation rewards, spend them again in conflicting transactions, and repeat this process sufficiently many times before GST, thereby accumulating a total double-spent value exceeding $qD$. ∎
Proof sketch of Theorem 3. When $f+k\geq q$, Byzantine and rational validators can successfully construct valid certificates for an arbitrary number of conflicting blocks and provide them to clients. For any stake $D$ that rational validators have locked before the run, there is a configuration in which each rational validator can double-spend more than $D$ coins. That is because the number of clients is arbitrary and unknown before the run. Even if the conflicting blocks are detected and the misbehaving validators get slashed, the attack is still beneficial for rational validators. ∎
Proof sketch of Theorem 3. The adversary can perform the double-spend attacks described in Theorems 33, before the synchronous bound $\Delta^{*}$ elapses. That is because both cases assume the same resilience against Byzantine adversaries and operate under similar network conditions during a network partition in $\Delta^{*}$. ∎
C.2. Full proofs
Proof of Theorem 3
Proof.
For the sake of contradiction assume that there is a q–commitable SMR protocol $\Pi$ that is $(n,k,f)$-resilient for $f\geq\lceil h/2\rceil$, where h=nkfh=n-k-f is the number of correct validators. Assume that each validator has deposited $D\in\mathbf{N}$ stake to participate in $\Pi$ and assume any period $\Delta_{W}$ where a validator can withdraw their stake.
We will construct a run where colluding with Byzantine validators to create valid certificates for conflicting blocks yields greater utility for rational validators.
Double spending attack. Since $f\geq\lceil h/2\rceil$, it holds that $f+k+\lfloor h/2\rfloor\geq q$. In Lemma A.1, we showed that a malicious validator can partition correct validators in two distinct subsets $H_{1}$ and $H_{2}$ monitoring two conflicting ledgers $T_{c}\cup T_{1}$ and $T_{c}\cup T_{2}$ respectively, where $T_{c}$ is the common part of the ledgers. Assume the attack takes place during an interval  [^14] We can use a reference point the local clock of any correct validator. $t,t^{\prime}]$ where $t^{\prime}>t+\Delta_{W}$ and let $\mathcal{F}$ be the set of the $f+k$ misbehaving validators.
We will construct $T_{1}$ and $T_{2}$ as follows. There is a partition of the clients in two sets $C_{1},C_{2}$, and a set of transactions $Tx=Tx_{1}\cup Tx_{2}$ input to the SMR protocol $\Pi$, where a set $\mathcal{R}$ of at least $q-f-min(|H_{1}|,|H_{2}|)$  [^15] If $q-f-\min(|H_{1}|,|H_{2}|)=0$, meaning there are enough Byzantine validators to carry out the attack, the analysis follows trivially. rational validators spend all their liquid coins $\epsilon>0$ in $Tx_{1}$ to a subset of clients in $C_{1}$ and the same $\epsilon>0$ coins in $Tx_{2}$ to a subset of clients $C_{2}$. Moreover, in $Tx_{1}$ and $Tx_{2}$, there are included all necessary transactions so that validators can withdraw their stake.
We will show that such an attack is feasible. That is, if rational nodes in $\mathcal{R}$ perform the attack, it leads to a safety violation. Consider a correct validator $p_{1}\in H_{1}$ and any correct validator $p_{2}\in H_{2}$. Since $\Pi$ solves SMR, it guarantees liveness, and therefore, every transaction in $Tx$ is committed to the ledger of $p_{1}$ and $p_{2}$  [^16] We assume that correct validators gossip valid transactions, so the transaction will eventually be witnessed by every correct validators. . We prove that $p_{1}$ and $p_{2}$ will finalize the blocks in the conflicting ledgers $T_{1}$ and $T_{2}$ respectively, using an indistinguishability argument.
World 1. (GST = 0.) Validators in $H_{1}$ and $F$ are correct. Validators in $H_{2}$ are Byzantine and do not participate in the protocol. The input to $\Pi$ are the transactions in $Tx_{1}$. Validator $p_{1}\in H_{1}$ witnesses blocks in $T_{1}$  [^17] We mean that $p_{1}$ sequentially finalizes blocks in $T_{1}$. with quorum certificates of nfn-f votes.
World 2. (GST = 0) This is symmetrical to World 1. Now, validators in $H_{2}$ and $F$ are correct. Validators in $H_{1}$ are Byzantine and do not participate in the protocol. The input to $\Pi$ are the transactions in $Tx_{2}$. Validator $p_{2}\in H_{2}$ witnesses blocks in $T_{2}$ along with the quorum certificates of nfn-f votes.
World 3. ($GST=t^{\prime}$). Validators in $F$ perform the Double Spending Attack, as explained before, until $GST$. First, the adversary partitions the network between validators in $H_{1}$ and $H_{2}$, such that messages sent by validators in $H_{1}$ are received from validators in $H_{2}$ only after $GST$ happens, and vice versa. The input to $p_{1}\in H_{1}$ are the transactions in $Tx_{1}$ and the input to $p_{2}\in H_{2}$ are the transactions in $Tx_{2}$. Then, every validator in $F$ simulates two correct validators. The first one simulates the execution of World 1 to any validator $p_{1}\in H_{1}$ and the second one simulates the execution of World 2 to any validator $p_{2}\in H_{2}$.
In the view of $p_{1}$ World 1 and World 3 are indistinguishable. Thus, $p_{1}$ finalizes blocks in $T_{1}$ after $T_{c}$. In the view of $p_{2}$ World 2 and World 3 are indistinguishable. Therefore, $p_{2}$ finalizes blocks in $T_{2}$ after $T_{c}$ leading to a safety violation.
Then, to argue that the attack is profitable for rational nodes we will show that clients will accept the inclusion proofs for conflicting blocks. Since $\Pi$ satisfies client-liveness, there must be a rule so that clients in $C_{1}$ (or $C_{2}$) can receive inclusion proofs for blocks in $T_{c}$. We will show that any rule $\rho$ that satisfies client-liveness, leads to clients accepting the inclusion proofs for the conflicting blocks in the double spending attack.
In q–commitable protocols, the rule $\rho$ can either depend on some time bound $\Delta$ or on a certificate of votes. However, we show that rule $\rho$ cannot depend on a time bound in the partial synchronous model using an indistinguishability argument. Consider the time bound $\Delta$ after GST. Moreover, consider any rule $\rho_{1}$ according which clients wait for the message delay $k\Delta$ for any $k\in\mathbb{Z}$ to receive a proof of equivocation and, w.l.o.g., let us fix some client $c_{1}\in C_{1}$.
World A. (GST=0). The next blocks with valid certificates extending $T_{c}$ are the blocks in $T_{1}$. Validator, $p_{1}$ commits blocks in $T_{1}$ in its local ledger. Client $c_{1}$ receives a valid certificate for any block $b\in T_{1}$ from $p_{i}$ and waits for $k\Delta$. Since there are not conflicting blocks in the system, client $c_{1}$ does not receive a proof of equivocation.
World B. ($GST>t^{\prime}+k\Delta)$. Validators in $\mathcal{F}$ perform the double spending attack at time $t$. Validator $p_{1}\in H_{1}$ finalizes blocks in $T_{1}$ to extend $T_{c}$ and validator $p_{2}\in H_{2}$ finalizes blocks in $T_{2}$ to extend $T_{c}$. Validator $p_{1}$ sends the inclusion proof for blocks in $T_{1}$ to $c_{1}$. Client $c_{1}$ waits for $k\Delta$. At $t^{\prime}+k\Delta<GST$, $c_{1}$ does not receive any message from validators in $H_{2}$ and does not witness conflicting blocks.
Client $c_{1}$ cannot distinguish between World A and World B and therefore, the rule $\rho_{1}$ is not sufficient. Next, consides certificates of stages of at least $t$ votes. since the $f$ Byzantine validators might never participate, $t\leq n-f$. Now, consider any rules $\rho_{a}$ and $\rho_{b}$, which are receiving stages of at least $t_{1}=n-f$ votes and $t_{2}$ votes for some $t_{2}\in[1,n-f)$ respectively. Any certificate validating $\rho_{1}$ also validates $\rho_{2}$. Therefore, an inclusion proof consisting of stages of $t_{1}=n-f\geq q$ votes is the strongest certificate that clients can receive, and will therefore accept the inclusion proofs for the blocks in the conflicting ledgers $T_{1},T_{2}$.  [^18] A different approach is to notice that $c_{1}$ cannot distinguish between World 1 and World 2 of the Double Spending Attack.
Proof of Theorem 3
Proof.
For the sake of contradiction, assume that there exists a qq–commitable SMR protocol $\Pi$ that is $(n,k,f)$-resilient for some $q$ s.t. $q\leq f+k+\lfloor h/2\rfloor$, where h=nkfh=n-k-f is the number of correct validators. Also, assume that each validator has deposited $D\in\mathbf{N}$ stake to participate in $\Pi$.
We construct an attack in which rational validators collude with Byzantine validators to produce valid certificates for conflicting blocks, yielding strictly greater utility for rational validators and leading to a safety violation. Similar to the proof of Theorem 3, a malicious validator can partition correct validators in two distinct subsets $H_{1}$ and $H_{2}$ which monitor the conflicting ledgers $T_{c}\cup T_{1}$ and $T_{c}\cup T_{2}$ respectively. Additionally, there is a partition of the clients in two sets $C_{1},C_{2}$, and the input to the blockchain protocol is $Tx=Tx_{1}\cup Tx_{2}$, where a set $\mathcal{R}$ of at least $p-f-min(|H_{1}|,|H_{2}|)$ rational validators spend $qD$ coins in $Tx_{1}$ to a subset of clients in $C_{1}$ and $qD$ coins in $Tx_{2}$ to a subset of clients $C_{2}$ where $\epsilon_{1},\epsilon_{2}>0$. Therefore, even if the attack is eventually detected and rational validators forfeit their stake, or if the stake of misbehaving validators is instead offered as a reward for reporting misbehavior, the total value double-spent makes the attack strictly beneficial for rational validators.
We will now construct a run where, for any $D$, every rational validator spend $qD$ coins in $T_{1}$ (same for $T_{2}$). To this end, cosnider the following notation. $G^{p}_{0}$ denotes the coins of validator $p$ in $T_{c}$ plus the maximum amount of bribes from malicious validators to fork the system. Then, we denote by $S^{p}_{i}$ the total coins paid to clients in $C_{1}$ by validator $p$ and by $R^{p}_{i}$ the participation rewards $p$ has received until the $i$-th block of $T_{1}$ (excluding $T_{c}$).
We need to show that for every rational validator $p$ there is a $k$ s.t. $S^{p}_{k}>qD$, or $G^{p}_{0}+R^{p}_{k}>qD$. If the initial coins held by $p$, or the coins provided by malicious clients as bribes, are unbounded, this condition trivially holds. We now show that the condition also holds if the participation rewards $R^{p}_{k}$ are unbounded.
Toward this direction, consider the first $i\in\mathcal{N}$ such that $R^{p}_{i}=\mathcal{C}^{p}_{0}$ as follows. Every validator in $\mathcal{R}$ spends all their coins, and therefore all coins belong to the hands of all the other users. When enough blocks has been processed in the system, any rational validator $p$ actively participating (by proposing or voting blocks depending on the reward scheme) will receive at least $\mathcal{C}^{p}_{0}$ in participation rewards. Specifically, w.l.o.g., if the participation rewards per block are $r$,$n$ blocks have been processed in the system and $p$ has participated in a percentage of $\alpha$ of them, then there is a run where $\alpha\cdot n\cdot r\geq\mathcal{C}^{p}_{0}$, which holds since it can be $n\to\infty$. Then, rational validators will spend all of their money to pay clients in $C_{1}$ and the above scenario is repeated for $j$ times such that there is $i\in\mathcal{N}$, $R^{p}_{i}=j\mathcal{C}^{p}_{0}$ and $(j+1)\mathcal{C}^{p}_{0}>qD$. Since it can be that $j\to\infty$, this is a feasible scenario. ∎
Proof of Theorem 3
Proof.
For the sake of contradiction assume that there is a q–commitable SMR protocol $\Pi$ that is $(n,k,f)$-resilient for $f\geq h$, where h=nkfh=n-k-f is the number of correct validators. Assume that each validator has deposited $D\in\mathbf{N}$ stake to participate in $\Pi$.
We will construct a run where it is a dominant strategy for rational validators to collude with Byzantine validators and create valid certificates for blocks that are not witnessed, and therefore not finalized, by any correct validator. Clients will accept the certificates, leading to a client-safety violation. Towards this direction, consider a run of $\Pi$ and let us denote by $C_{t}$ the number of clients in the system at time $t$. Assume that all correct validators agree in the transaction ledger $T_{c}$ at time $t-\epsilon$ for some $\epsilon>0$.
Double spending attack. The input of $\Pi$ at time $t$ is a set of transactions $Tx=\cup_{i=1}^{C_{t}}Tx_{i}$ constructed as follows. There is a set $\mathcal{R}$ of at least r=max(qf,0)r=max(q-f,0) rational validators such that every validator in $\mathcal{R}$ transfers $c>0$ coins to the client $c_{i\in\{1,\dots,C_{t}\}}$ in $Tx_{i\in\{1,\dots,C_{t}\}}$. Moreover, we construct the run such that $(C_{t}-1)C>D$. Since the deposit $D$ is decided before the run of the protocol and $C_{t}$ is arbitrary, it is feasible to construct such a run.
Validators in $\mathcal{R}$, along with the $f$ Byzantine validators, create the blocks $b_{i\in\{1,\dots,C_{t}\}}$, where for each $i,b_{i}$ includes the transactions $Tx_{i}$. Since $r+f\geq q$, they can construct $s$ phases (for any $s$ specified by $\Pi$) of $q$ votes for every blocks $b_{i\in\{1,\dots,C_{t}\}}$. Note that none of the correct validators has witnessed the valid certificate for any of those blocks.
We will show that clients $c_{i\in\{1,\dots,C_{t}\}}$ will accept the inclusion proofs for blocks $b_{i\in\{1,\dots,C_{t}\}}$. In that scenario, any rational validator $p\in\mathcal{R}$ double spends $(C_{t}-1)C>D$ coins. Even if validator $p$ is slashed in $\Pi$, the attack is still beneficial. This will lead to violation of client-safety.
Clients must accept the inclusion proofs. Since $\Pi$ ensures client-liveness, there is a rule $\rho$ that allows clients to accept inclusion proofs for every block committed to the ledger $T_{c}$. We will show that for any such rule $\rho$, all clients $c_{i\in\{1,\dots,C_{t}\}}$ will accept inclusion proofs for blocks $b_{i\in\{1,\dots,C_{t}\}}$.
First, consider a rule $\rho_{1}$ where clients wait for a fixed time bound $k\Delta$, where $\Delta$ is the synchrony bound and $k>0$. We fix a client $c_{i\in\{1,\dots,C_{t}\}}$. We prove that this rule is insufficient using an indistinguishability argument.
World 1. The next block extending $T_{c}$ with a valid certificate is block $b_{i}$. At least one correct validator $p_{i}$, commits $b_{i}$ in its local ledger. There are not conflicting blocks in the system. Client $c_{i}$ receives a valid certificate for $b_{i}$ from $p_{i}$ and waits for $k\Delta$. After the period passes $c_{i}$ detects no conflict.
World 2. Rational nodes collude with Byzantine nodes to perform the double spending attack described above. A rational node $p\in\mathcal{R}$ sends the inclusion proof of $b_{i}$ to $c_{i}$. Client $c_{i}$ waits for $k\Delta$. Since no correct validator has witnessed the certificates for any conflicting block $b_{j\in\{1,\dots,C_{t}\},j\neq i}$, client $c_{i}$ detects no conflict.
Client $c_{i}$ cannot distinguish between World 1 and World 2 and therefore, the rule $\rho_{1}$ is not sufficient. Now, consider any rule $\rho_{1}$ which requires stages of at least $t$ votes. Similar to the proof of Theorem 3, an inclusion proof of stages of t=nft=n-f votes is the strongest certificate that clients can receive. Therefore, each client $c_{i\in\{1,\dots,C_{t}\}}$ clients will accept the inclusion proofs for the respective block $b_{i}$.
Proof of Theorem 3
Proof.
Assume, for contradiction, that there exists a qq–commitable SMR protocol $\Pi$ that is $(n,k,f)$-resilient for i) $f\geq\lceil h/2\rceil$, ii) or , where h=nkfh=n-k-f represents the number of correct validators, and $\Pi$ does not incur a finality delay of at least $\Delta^{*}$. Moreover, suppose each validator has staked $D\in\mathbf{N}$ to take part in $\Pi$.
Theorems 3 and 3 share the same Byzantine resilience and assume similar network conditions during a partition lasting up to $\Delta^{*}$, as in Theorem 3, with $f\geq\lceil h/2\rceil$ and $q\leq f+k+\lfloor h/2\rfloor$, respectively. Thus, the proofs of Theorems 3 and 3 extend to show that: (1) safety is violated, with correct validators $p_{1}\in H_{1}$ and $p_{2}\in H_{2}$ finalizing conflicting blocks in ledgers $T_{1}$ and $T_{2}$, and (2) clients accept inclusion proofs backed by valid certificates. ∎